### v4

# New systems of log-canonical coordinates on $SL(2,{\mathbb{C}})$ character varieties of compact Riemann surfaces

Abstract. We construct new sets of log-canonical coordinates on $SL(2,\mathbb{C})$ character varieties of compact Riemann surfaces. These coordinates are obtained by combining shear type coordinates with length-twist type coordinates. On the real component corresponding to the moduli space of compact Riemann surfaces $\mathcal{M}_{g}$, the generating function corresponding to the symplectomorphism between these new coordinates and Fenchel-Nielsen coordinates is explicitly computed.
1. Introduction
The $SL(2,{\mathbb{C}})$ character variety $\mathcal{V}_{g}$ of compact Riemann surfaces of genus $g$ is equipped with the natural complex symplectic form, inverse to the Goldman Poisson bracket on $\mathcal{V}_{g}$ [13]. There exists a set of Darboux coordinates on $\mathcal{V}_{g}$ which is obtained by analytic continuation to the complex domain of the Fenchel-Nielsen coordinates on the Teichmüller space $\mathcal{T}_{g}$ [20].
Once one introduces $n$ boundary components on the Riemann surface and considers the corresponding character variety, there exists a set of coordinates which are extensions of Thurston’s shear coordinates and are log-canonical for the Goldman form on the symplectic leaves (the coordinates are called log-canonical if the coefficients of the 2-form written in terms of those coordinates are constant, the actual Darboux coordinates can then be obtained as linear combinations of those coordinates).
However, so far it remains unclear how to use the (complexified) shear coordinates efficiently on $\mathcal{V}_{g}$, although important steps in this direction have been made in [5]. In the case of $\mathcal{V}_{2}$, a set of log-canonical coordinates alternative to Fenchel-Nielsen were recently constructed in [6]; however, generalizing this construction to higher genus was still an open question.
In this paper, we construct new systems of log-canonical coordinates on $\mathcal{V}_{g}$ for any $g$ by combining the ideas behind the (complexified) shear coordinates and the twist-length coordinates. The technical framework is in [3], where for each graph on a Riemann surface and an assignment of $SL(2,{\mathbb{C}})$ matrices on edges, one associates a canonical two-form.
The main result is that we associate a new system of (local) log-canonical coordinates on $\mathcal{V}_{g}$ to any system of simple non-intersecting closed contours $\{\gamma_{1},\dots,\gamma_{m}\}$, for any $m=1,\dots,3g-3$. In the case of maximal m=3g3m=3g-3, the set of contours define a “trinion decomposition” of the Riemann surface into three-holed spheres (the trinions or pairs of pants). In this case, our coordinates are closely related (but not exactly coinciding) with the complexified Fenchel-Nielsen length-twist coordinates; their relationship is explained in Section 5.
We start by presenting this construction for the simplest case $m=1$, which corresponds to a single contour $\gamma$ that separates $\mathcal{C}$} into two Riemann surfaces, $\widehat{{\mathcal{C}}}$ and $\widetilde{{\mathcal{C}}}$ of genera $\widetilde{g}$ and $\widehat{g}$, respectively. The associated system of log-canonical coordinates on $\mathcal{V}_{g}$ consists of a pair of coordinates of length-twist type along $\gamma$, and also complex shear type coordinates on $SL(2,{\mathbb{C}})$ character varieties of Riemann surfaces $\widehat{{\mathcal{C}}}$ and $\widetilde{{\mathcal{C}}}$ with the boundaries collapsed to a point.
More precisely, given a representation $\rho$ in $\mathcal{V}_{g}$, let $\lambda_{\gamma}$ and $\lambda_{\gamma}^{-1}$ be the eigenvalues of the monodromy matrix $\rho(\gamma)=M_{\gamma}$, and denote $\ell_{\gamma}=\ln\lambda_{\gamma}$, with some choice of branch for the complex logarithm. Let us choose a point $\widetilde{v}\in\widetilde{{\mathcal{C}}}$ in a neighborhood of $\gamma$ on the $\widetilde{{\mathcal{C}}}$ side, and similarly another point $\widehat{v}\in\widehat{{\mathcal{C}}}$ that is also close to $\gamma$ but lying on the $\widehat{{\mathcal{C}}}$ side. Consider a fatgraph $\widetilde{\Sigma}$ on $\widetilde{{\mathcal{C}}}$ with only one vertex at $\widetilde{v}$, and a fatgraph $\widehat{\Sigma}$ on $\widehat{\mathcal{C}}$ with a single vertex at $\widehat{v}$. These graphs are chosen to triangulate $\widetilde{{\mathcal{C}}}$ and $\widehat{{\mathcal{C}}}$ if their boundaries were collapsed to a point.
The number of edges of $\widetilde{\Sigma}$ and $\widehat{\Sigma}$ equals $\widetilde{n}=6\widetilde{g}-3$ and $\widehat{n}=6\widehat{g}-3$, respectively. The valences of the vertices $\widetilde{v}$ and $\widehat{v}$ are equal to $2\widetilde{n}$ and $2\widehat{n}$, respectively, since every edge of $\widetilde{\Sigma}$ (resp. $\widehat{\Sigma}$) comes to $\widetilde{v}$ (resp. $\widehat{v}$) twice.
To each edge, $\widetilde{e}_{i}$, of the graph $\widetilde{\Sigma}$ we assign a coordinate $\widetilde{\zeta}_{e_{i}}\in{\mathbb{C}}$ and $\widetilde{z}_{e_{i}}={\rm e}^{\widetilde{\zeta}_{e_{i}}}$, which are complex coordinates of shear type parametrizing the character variety $\mathcal{V}_{\widetilde{g},1}$ of $\widetilde{{\mathcal{C}}}$. The complex “length” $\ell_{\gamma}$ of the separating loop $\gamma$ is related to these coordinates by
$2\sum_{i=1}^{\widetilde{n}}\widetilde{\zeta}_{e_{i}}=\ell_{\gamma}\;.$ (1)
We shall continue calling $\widetilde{\zeta}_{e_{i}}$ shear or shear type coordinates in this paper although they really coincide with logarithms of Thurston’s shear coordinates only when the length of the boundary curve is zero. Similar coordinates $\widehat{\zeta}_{e_{j}}$’s and $\widehat{z}_{e_{j}}$’s are assigned to the edges $\widehat{e}_{j}$ of $\widehat{\Sigma}$ and they satisfy the similar relation
$2\sum_{j=1}^{\widehat{n}}\widehat{\zeta}_{e_{j}}=\ell_{\gamma}\;.$ (2)
Using (1), (2) we can eliminate $\widehat{\zeta}_{1}$ and $\widetilde{\zeta}_{1}$ so that the full set of independent complex local coordinates on $\mathcal{V}_{g}$ looks as follows:
$\left\{\{\widetilde{\zeta}_{e_{i}}\}_{i=2}^{\widetilde{n}},\,\{\widehat{\zeta}_{e_{j}}\}_{j=2}^{\widehat{n}},\ell_{\gamma},\beta_{\gamma}\right\}\;.$ (3)
The total number of variables in the set (3) is given by $\widetilde{n}+\widehat{n}=6g-6$ which coincides with the dimension of $\mathcal{V}_{g}$. The coordinate $\beta_{\gamma}$ is the complex toric variable along contour $\gamma$ (similar but not coinciding with the complexification of the Fenchel-Nielsen twist).
Then, by Theorem 7.1, the form $\Omega$ on $\mathcal{V}_{g}$ can be written as follows in terms of coordinates (3):
$\Omega=\Omega_{0}+\widetilde{\Omega}_{1}+\widehat{\Omega}_{1}+\widetilde{\Omega}_{2}+\widehat{\Omega}_{2}\;,$ (4)
where
$\Omega_{0}=\mathrm{d}\beta_{\gamma}\wedge\mathrm{d}\ell_{\gamma}\;,$
$\widetilde{\Omega}_{1}=\sum_{\begin{subarray}{c}i,j=2\\ i<j\end{subarray}}^{6\widetilde{g}-3}\left(\widetilde{c}_{ij}-\widetilde{c}_{1j}+\widetilde{c}_{1i}\right)\mathrm{d}\widetilde{\zeta}_{i}\wedge\mathrm{d}\widetilde{\zeta}_{j}\;,\qquad\widehat{\Omega}_{1}=\sum_{\begin{subarray}{c}k,l=2\\ k<l\end{subarray}}^{6\widehat{g}-3}\left(\widehat{c}_{kl}-\widehat{c}_{1l}+\widehat{c}_{1k}\right)\mathrm{d}\widehat{\zeta}_{k}\wedge\mathrm{d}\widehat{\zeta}_{l}\;,$
$\widetilde{\Omega}_{2}=\frac{1}{2}\sum_{m=2}^{6\widetilde{g}-3}\widetilde{c}_{1m}\mathrm{d}\ell_{\gamma}\wedge\mathrm{d}\widetilde{\zeta}_{m}\;,\qquad\widehat{\Omega}_{2}=\frac{1}{2}\sum_{n=2}^{6\widehat{g}-3}\widehat{c}_{1n}\mathrm{d}\ell_{\gamma}\wedge\mathrm{d}\widehat{\zeta}_{n}\;.$
The integer coefficients $\widetilde{c}_{ij}$, $\widehat{c}_{kl}$ lie in the interval between 4-4 and $4$. They are determined by the order of the edges of the one-valent graphs $\widetilde{\Sigma}$ and $\widehat{\Sigma}$ at their vertex.
This construction extends to an arbitrary system of simple non-intersecting contours $\gamma_{1},\dots,\gamma_{m}$}, $1\leq m\leq 3g-3$, which split $\mathcal{C}$} into $n$ subsurfaces ${\mathcal{C}}^{(i)}$ with $k^{(i)}$ boundary components. Then we have $m$ pairs of log-canonical coordinates $\{\beta_{\gamma_{j}},\ell_{\gamma_{j}}\}_{j=1}^{m}$. Each subsurface ${\mathcal{C}}^{(i)}$ carries a triangulation graph $\Sigma^{(i)}$ with $6g^{(i)}-6+3k^{(i)}$ edges, to which we assign shear coordinates $\zeta_{e}$. In these coordinates, by Theorem 8.1, one can write the Goldman form on $\mathcal{V}_{g}$ as
$\Omega=\sum_{i=1}^{n}\Omega^{(i)}+\sum_{j=1}^{m}{\mathrm{d}}\beta_{\gamma_{j}}\wedge{\mathrm{d}}\ell_{\gamma_{j}}\;,$ (5)
where
$\Omega^{(i)}=\sum_{v\in V(\Sigma^{(i)})}\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v\\ \tilde{e}<\widetilde{e}\end{subarray}}{\mathrm{d}}\zeta_{e}\wedge{\mathrm{d}}\zeta_{\widetilde{e}}\;$.
When m=3g3m=3g-3, the surface $\mathcal{C}$} decomposes into 2g22g-2 trinions. Since each ${\mathcal{C}}^{(i)}$ is a trinion, all shear type coordinates can be expressed in terms of the boundary lengths of ${\mathcal{C}}^{(i)}$, due to the constraints (1). This yields 3g33g-3 log-canonical coordinates $\{\beta_{\gamma_{j}},\ell_{\gamma_{j}}\}_{j=1}^{3g-3}$ for the Goldman form $\Omega$ on $\mathcal{V}_{g}$.
Using the trinion graph $\Gamma_{trin}$, whose vertices correspond to trinions and edges to the gluings of their boundary components, one may rewrite $\Omega$ as (cf. Theorem 9.2)
$\Omega=\frac{1}{2}\sum_{v\in V(\Gamma_{trin})}\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v\\ e<\widetilde{e}\end{subarray}}\mathrm{d}\ell_{e}\wedge\mathrm{d}\ell_{\widetilde{e}}+\sum_{j=1}^{3g-3}{\mathrm{d}}\beta_{\gamma_{j}}\wedge{\mathrm{d}}\ell_{\gamma_{j}}\;.$ (6)
For each vertex, the inner sum runs over pairs of incident edges, ordered according to the inherited orientation of the ciliated graph. This expression is independent of the choice of cilium.
When restricted to the real Fuchsian component which projects to the Teichmuller space $\mathcal{T}_{g}$, the form $\Omega$ coincides with the Weil-Petersson form $\omega_{WP}$. In terms of Fenchel-Nielsen coordinates $\{\tau_{\gamma_{j}},\ell_{\gamma_{j}}\}_{j=1}^{3g-3}$, the form $\omega_{WP}$ can be written [20]
$\omega_{WP}=\sum_{j=1}^{3g-3}{\mathrm{d}}\tau_{\gamma_{j}}\wedge{\mathrm{d}}\ell_{\gamma_{j}}\;.$ (7)
The length coordinates $\ell_{\gamma_{j}}$ are the same in (7) and (6) while each Fenchel-Nielsen twist coordinate $\tau_{\gamma_{j}}$ differs from $\beta_{\gamma_{j}}$ by adding a function that depends only on the lengths. The generating function of the symplectomorphism between log-canonical coordinates $\{\beta_{\gamma_{j}},\ell_{\gamma_{j}}\}_{j=1}^{3g-3}$ and Fenchel-Nielsen coordinates $\{\tau_{\gamma_{j}},\ell_{\gamma_{j}}\}_{j=1}^{3g-3}$ can be expressed in terms of Euler dilogarithms.
This construction admits a straightforward generalization to the case of an arbitrary $SL(N,\mathbb{C})$ character variety of $\pi_{1}({\mathcal{C}})$. Here, one needs to use the Fock-Goncharov coordinates associated to each fatgraph $\Gamma^{(i)}$ together with $N$ eigenvalues and toric variables associated to each $M_{\gamma_{j}}$.
Some explicit computations of the Goldman symplectic form $\Omega$ for low genera are given in an appendix.
2. Main spaces
2.1. Character variety of a compact Riemann surface
Let $\mathcal{C}$} be an oriented compact Riemann surface of genus $g$. Then the standard generators $\{\alpha_{k},\beta_{k}\}_{k=1}^{g}$ of the fundamental group $\pi_{1}({\mathcal{C}})$ can be chosen such that
$\prod_{i=1}^{g}\left[\alpha_{i},\beta_{i}\right]=id\;,$ (8)
where
$\left[\alpha_{i},\beta_{i}\right]=\alpha_{i}\beta_{i}^{-1}\alpha_{i}^{-1}\beta_{i}\;.$
A point of the $SL(2,{\mathbb{C}})$ character variety $\mathcal{V}_{g}$ is represented by the equivalence class under conjugation of the set of monodromy matrices $\{M_{\alpha_{1}},M_{\beta_{1}}...,M_{\alpha_{g}},M_{\beta_{g}}\}\subset SL(2,{\mathbb{C}})$ satisfying
$\prod_{i=1}^{g}\left[M_{\alpha_{i}},M_{\beta_{i}}\right]=\mathbb{I}_{2}\;.$ (9)
Traces of monodromy matrices can be used as coordinates on $\mathcal{V}_{g}$.
The Goldman Poisson bracket on $\mathcal{V}_{g}$ (the complexification of the Weil-Petersson bracket on $\mathcal{M}_{g}$) can be written as follows in terms of trace coordinates
[…]
where $\gamma,\tilde{\gamma}\in\pi_{1}({\mathcal{C}})$ and $\nu(p)$ is the contribution of the point $p$ to the intersection index of $\gamma$ and $\tilde{\gamma}$.
The Goldman bracket on $\mathcal{V}_{g}$ is non-degenerate. The symplectic form inverting the Goldman’s bracket on an arbitrary $SL(N,\mathbb{C})$ character variety of a compact Riemann surface was found in [2], but the corresponding log-canonical coordinates for them are unknown.
Below, we show how to construct systems of Darboux coordinates on $\mathcal{V}_{g}$ which can be naturally extended to higher $SL(N,\mathbb{C})$ character varieties using the machinery of Fock-Goncharov coordinates.
Let now $\mathcal{C}$} be a closed and oriented Riemann surface of genus $g$ with $n\geq 1$ marked points $t_{1},\dots,t_{n}$. Then the standard generators $\{\gamma_{j}\}_{j=1}^{n}$, $\{\alpha_{k},\beta_{k}\}_{k=1}^{g}$ of the fundamental group $\pi_{1}({\mathcal{C}}\setminus\{t_{j}\})$ can be chosen such that:
$\quad\gamma_{1}...\gamma_{n}\prod_{i=1}^{g}\left[\alpha_{i},\beta_{i}\right]=id\;.$
Consider now a representation of $\pi_{1}({\mathcal{C}}\setminus\{t_{j}\})$ in $SL(2,{\mathbb{C}})$. Then monodromy matrices corresponding to the standard generators of the fundamental group satisfy the same relation:
$M_{\gamma_{1}}...M_{\gamma_{n}}\prod_{i=1}^{g}\left[M_{\alpha_{i}},M_{\beta_{i}}\right]=\mathbb{I}_{2}\;.$ (11)
We denote the eigenvalues of $M_{\gamma_{j}}$ by $\lambda_{\gamma_{j}}$ and $1/\lambda_{\gamma_{j}}$.
Define the character variety $\mathcal{V}_{g,n}$ as the space of monodromy representations modulo simultaneous conjugation by $SL(2,\mathbb{C})$. The coordinates on $\mathcal{V}_{g,n}$ can be again chosen to be the trace coordinates, i.e. ${\rm tr}M_{\gamma}$, for sufficiently many elements $\gamma$ of the fundamental group.
The Goldman bracket on $\mathcal{V}_{g,n}$ is defined by the same formula (10). However, on $\mathcal{V}_{g,n}$ it is degenerate, with the Casimir elements being the eigenvalues of $M_{\gamma_{j}}$.
2.2. Extended character varieties
The character varieties $\mathcal{V}_{g,n}$ admit natural extensions ${{\mathcal{H}}}_{g,n}$ that possess nondegenerate Poisson structures. The idea of such an extension is to introduce canonical partners to the Casimirs $\ell_{\gamma_{1}},...,\ell_{\gamma_{n}}$ of the Goldman bracket on $\mathcal{V}_{g,n}$. Suppose that all monodromies $M_{\gamma_{k}}$ are diagonalizable with
$M_{\gamma_{k}}=C_{k}\Lambda_{k}C_{k}^{-1}\;.$
Here we do not fix the order of the eigenvalues and eigenvectors of $M_{\gamma_{k}}$. The Casimirs $\ell_{\gamma_{k}}$ are defined by $\ell_{\gamma_{k}}=\pm\ln{\lambda_{\gamma_{k}}}$, where $\ln$ denotes a branch of the complex logarithm (on the real slice we will assume $\lambda_{\gamma_{k}}>1$ and analytic continuation can be obtained by choosing the principal value).
Define the extended space $\mathcal{H}_{g,n}$ as follows
${{\mathcal{H}}}_{g,n}=\left\{\{M_{\alpha_{j}},M_{\beta_{j}}\}_{j=1}^{g},\{C_{k},\Lambda_{k}\}_{k=1}^{n}:\qquad C_{1}\Lambda_{1}C_{1}^{-1}\dots C_{N}\Lambda_{N}C_{N}^{-1}\prod_{i=1}^{g}\left[M_{\alpha_{i}},M_{\beta_{i}}\right]=\mathbb{I}_{2}\right\}/\sim\;.$ (12)
The sets related by transformation $M_{\gamma_{k}}\to GM_{\gamma_{k}}G^{-1}$, $C_{k}\to GC_{k}$ for a fixed $G\in SL(2,{\mathbb{C}})$ (while $\Lambda_{k}$ remain invariant) are assumed to be equivalent.
The space ${{\mathcal{H}}}_{g,n}$ for an arbitrary matrix dimension was studied in [10], [4] and [3]. It possesses a natural symplectic form [4, 3]. The canonical partners of the Casimirs are the coordinates $\beta_{\gamma_{k}}$, which are the logarithms of the so-called toric variables $b_{\gamma_{k}}$. The toric variables are the normalizing factors of the eigenvectors of $M_{\gamma_{k}}$ encoded in the matrices $C_{k}$. In [3] this symplectic form was represented in a log-canonical form for an arbitrary $SL(N,\mathbb{C})$ group by combining Fock-Goncharov coordinates with the toric variables.
3. Graphs on surfaces and canonical symplectic form
Let $\Gamma$ be a finite ciliated oriented graph embedded into a Riemann surface $\mathcal{C}$}, with $V(\Gamma)$ and $E(\Gamma)$ denoting the sets of its vertices and edges respectively. The chosen cilium at each vertex induces the ordering of edges attached to this vertex. One writes $e<e^{\prime}$ when the edge $e$ precedes the edge $e^{\prime}$ in a counterclockwise order from the cilium.
We will use the following notation: $e=[v_{1},v_{2}]$ denotes the oriented edge connecting $v_{1}$ to $v_{2}$ and $e=[v_{2},v_{1}]$ denotes the oriented edge connecting $v_{2}$ to $v_{1}$. If an edge $e$ is incident to a vertex $v$ we write $e\perp v$. We define the valency $n_{v}$ of a given vertex $v$ to be the total number of edges incident to $v$ counted with multiplicity.
Definition 3.1.
A pair $(\Gamma,J)$ consisting of an oriented graph on $\mathcal{C}$} up to isotopy and a map $J:E(\Gamma)\rightarrow SL(2,{\mathbb{C}})$ is canonical if it satisfied the following conditions:
  1. (1)
    The map $J:E(\Gamma)\rightarrow SL(2,{\mathbb{C}})$ satisfies $J(-e)=J(e)^{-1}$. The matrix $J(e)$ is called the jump matrix associated to the edge $e$.
  2. (2)
    For each vertex $v\in V(\Gamma)$ of valence $n_{v}\geq 2$ with incident incoming edges $\{e_{1},e_{2},...,e_{n_{v}}\}$ enumerated counterclockwise starting from an arbitrary chosen edge the local monodromy around $v$ is trivial, i.e.
    $J(e_{1})J(e_{2})...J(e_{n_{v}})=\mathbb{I}_{2}\;.$ (13)
To each pair $(\Gamma,J)$ we associate the canonical two-form $\Omega(\Gamma)$ as in [3]:
$\Omega({\Gamma})=\frac{1}{2}\,{\rm tr}\sum_{v\in V(\Gamma)}\sum_{l=1}^{n_{v}-1}\left(J_{[1...l]}^{(v)}\right)^{-1}{\mathrm{d}}J_{[1...l]}^{(v)}\wedge\left(J_{l}^{(v)}\right)^{-1}{\mathrm{d}}J_{l}^{(v)}\;,$ (14)
where $J_{l}^{(v)}=J(e_{l})$, $e_{l}\perp v$, are the matrices defined by the map in Definition 3.1 corresponding to the edges $e_{1},e_{2},...,e_{n_{v}}$ incident to $v$, and
$J_{[1...l]}^{(v)}=J_{1}^{(v)}J_{2}^{(v)}...J_{l}^{(v)}\;.$
The form $\Omega({\Gamma})$ is invariant under the set of the following admissible moves [3]:
  1. (1)
    merging of two vertices $v_{1}$ and $v_{2}$ of valency $n_{v_{1}}+1$ and $n_{v_{2}}+1$ respectively, connected by a common edge $e_{0}$, to get a single vertex $v$ of valency $n_{v_{1}}+n_{v_{2}}$,
  2. (2)
    zipping of two edges $e_{1}$ and $e_{2}$ with jump matrices $J_{1}$ and $J_{2}$ respectively, connecting the same two vertices $v_{1}$ and $v_{2}$, to get a single edge $e$ with jump matrix $J_{1}J_{2}$.
4. Standard graph and symplectic form on $\mathcal{V}_{g}$
Let $\mathcal{C}$} be a compact unpunctured Riemann surface of genus $g$. Choose a base point $v_{0}$ and a set of standard generators $\{\alpha_{j},\beta_{j}\}_{j=1}^{g}$ of $\pi_{1}({\mathcal{C}},v_{0})$ satisfying (8). Denote by $M_{\alpha_{i}}$, $M_{\beta_{i}}$ the corresponding monodromies which satisfy (11) for $n=0$:
$\prod_{j=1}^{g}[M_{\alpha_{j}},M_{b_{j}}]=\mathbb{I}_{2}\;.$ (15)
Let us choose some oriented closed contours on $\mathcal{C}$} starting at $v_{0}$ representing the elements $\alpha_{j}$, $\beta_{j}$; with a slight abuse of notations we denote these loops by the same letters. Consider the one-valent ciliated graph $\Gamma_{0}$ (which we call the standard graph, see Figure 1 (a)) with vertex at $v_{0}$ and $2g$ oriented edges $\{\alpha_{j},\beta_{j}\}_{j=1}^{g}$; the valency of $v_{0}$ equals $4g$. Define the jump matrices on the edges of $\Gamma_{0}$ as follows:
$J(\alpha_{j})=M_{\beta_{j}}\;,\qquad J(\beta_{j})=M_{\alpha_{j}}\;.$ (16)
The cilium at $v_{0}$ is chosen between $\alpha_{g}$ and $\beta_{1}$ (see Fig 1 (a)). Notice that the relation (13) in the general definition of jump matrices on a given graph is written under assumption that all the edges $e_{j}$ are incoming to the vertex. In our case, since every edge meets the vertex twice, while defining the canonical form (14) we need to inverse the orientation of $2g$ out of $4g$ half-edges incident to $v_{0}$ as shown in Fig 1 (b).
Figure 1. Orientation of the edges of the standard graph $\Gamma_{0}$ for $g=2$.
Then we get:
$J(e_{1})=M_{\alpha_{1}}\;,\qquad J(e_{2})=M_{\beta_{1}}^{-1}\;,\qquad J(e_{3})=M_{\alpha_{1}}^{-1}\;,\qquad J(e_{4})=M_{\beta_{1}}\;,\dots\;.$
Under these conventions we get
$J(e_{1})J(e_{2})...J(e_{4g})=\prod_{i=1}^{g}M_{\alpha_{i}}M_{\beta_{i}}^{-1}M_{\alpha_{i}}^{-1}M_{\beta_{i}}=\mathbb{I}_{2}\;.$
Therefore, the standard graph $\Gamma_{0}$ with our choice of jump matrices satisfies the condition (13).
Applying the formula (14) to the pair $(\Gamma_{0},J)$ we get the symplectic form $\Omega(\Gamma_{0})$ which can be written as follows:
$\Omega(\Gamma_{0})=\frac{1}{2}\sum_{j=1}^{2g-1}\mathrm{tr}\left(K_{[1...j]}^{-1}dK_{[1...j]}\wedge K_{j}^{-1}dK_{j}\right)\;,$ (17)
where $K_{j}$ is the $j$th matrix in the product (15); $K_{[1...j]}$ denotes the ordered product of matrices entering in the same product (15) up to $j$; for instance
$K_{[1]}=M_{\alpha_{1}}\;,\qquad K_{[12]}=M_{\alpha_{1}}M_{\beta_{1}}^{-1}\;,\qquad K_{[123]}=M_{\alpha_{1}}M_{\beta_{1}}^{-1}M_{\alpha_{1}}^{-1}\;,\dots\;.$
Proposition 4.1.
The form $\Omega(\Gamma_{0})$ given by (17) coincides with the Goldman symplectic form $\Omega$ on $\mathcal{V}_{g}$.
Proof.
The symplectic form $\Omega$ inverting the Goldman bracket on the $SL(N,\mathbb{C})$ character variety of $\pi_{1}({\mathcal{C}})$ was computed by Alekseev and Malkin [2]. The form of [2] can be represented via (14) for a special graph $G$ and the jump matrices [3]. In turns out that the Alekseev-Malkin graph can be transformed to the standard graph $\Gamma_{0}$ by a series of admissible moves preserving the symplectic form. Therefore, the form $\Omega(\Gamma_{0})$ coincides with the Goldman symplectic form (for an arbitrary $SL(N,\mathbb{C})$ case).
5. Parametrization of $\mathcal{V}_{g}$ by plumbing two lower genus Riemann surfaces
Here, we are going to construct another oriented ciliated graph on $\mathcal{C}$} (with jump matrices on its edges) which can be transformed to $\Gamma_{0}$ by a series of admissible moves. This allows us to get an alternative representation for the form $\Omega(\Gamma_{0})$ which will be used in the next section to construct new systems of log-canonical coordinates on $\mathcal{V}_{g}$.
5.1. Oriented graph on $\mathcal{C}$} by amalgamation of standard graphs on two Riemann surfaces of lower genera
Consider two topological Riemann surfaces, $\widetilde{{\mathcal{C}}}$ and $\widehat{{\mathcal{C}}}$, of genera $\widetilde{g}$ and $\widehat{g}$, respectively (such that $g=\widetilde{g}+\widehat{g}$) with one boundary component each. We denote by $\widetilde{\gamma}$ (resp. $\widehat{\gamma}$) a contour slightly shifted from the boundary to the interior of $\widetilde{{\mathcal{C}}}$ (resp. $\widehat{{\mathcal{C}}}$).
[…]
$M_{\widetilde{\gamma}}^{-1}=\prod_{i=1}^{\widetilde{g}}\left[M_{\widetilde{\alpha}_{i}},M_{\widetilde{\beta}_{i}}\right]\;.$ (18)
We assume that $M_{\widetilde{\gamma}}$ is diagonalizable, $M_{\widetilde{\gamma}}=\widetilde{C}\widetilde{\Lambda}\widetilde{C}^{-1}$ where the matrix of eigenvectors $\widetilde{C}$ has a freedom of multiplication with $SL(2,{\mathbb{C}})$ diagonal matrix from the right.
[…]
$M_{\widehat{\gamma}}^{-1}=\prod_{i=1}^{\widehat{g}}\left[M_{\widehat{\alpha}_{i}},M_{\widehat{\beta}_{i}}\right]\;,$ (19)
which is also assumed to be diagonalizable, $M_{\widehat{\gamma}}=\widehat{C}\widehat{\Lambda}\widehat{C}^{-1}$.
Define the following moduli space:
$\widetilde{{\mathcal{H}}}=\left\{\left\{M_{\widetilde{\alpha}_{j}},M_{\widetilde{\beta}_{j}}\right\}_{j=1}^{\widetilde{g}},\;\;\{\widetilde{C},\widetilde{\Lambda}\}:\qquad\widetilde{C}\widetilde{\Lambda}\widetilde{C}^{-1}\prod_{i=1}^{\widetilde{g}}\left[M_{\widetilde{\alpha}_{i}},M_{\widetilde{\beta}_{i}}\right]=\mathbb{I}_{2}\right\}/\sim\;,$ (20)
where $\widetilde{C},\widetilde{\Lambda}\in SL(2,{\mathbb{C}})$ and $\widetilde{\Lambda}=\mathrm{diag}(\lambda_{\widetilde{\gamma}},\lambda_{\widetilde{\gamma}}^{-1})$.
The equivalence relation $\sim$ is the same as in (12). The monodromies $\{M_{\widetilde{\alpha}_{j}},M_{\widetilde{\beta}_{j}}\}$ can be chosen arbitrarily; they determine the matrix $\widetilde{\Lambda}$ uniquely. The matrix $\widetilde{C}$ is then defined up to the right multiplication with a diagonal matrix.
Similarly, consider the moduli space:
$\widehat{{\mathcal{H}}}=\left\{\left\{M_{\widehat{\alpha}_{j}},M_{\widehat{\beta}_{j}}\right\}_{j=1}^{\widehat{g}},\{\widehat{C},\widehat{\Lambda}\}:\qquad\widehat{C}\widehat{\Lambda}\widehat{C}^{-1}\prod_{i=1}^{\widehat{g}}\left[M_{\widehat{\alpha}_{i}},M_{\widehat{\beta}_{i}}\right]=\mathbb{I}_{2}\right\}/\sim\;,$ (21)
with $\widehat{C},\widehat{\Lambda}\in SL(2,\mathbb{C})$ and $\widehat{\Lambda}=\mathrm{diag}(\lambda_{\widehat{\gamma}},\lambda_{\widehat{\gamma}}^{-1})$.
The idea of the following proposition is to construct an element of $\mathcal{V}_{g}$ starting from a set of matrices representing a point of $\widetilde{{\mathcal{H}}}$ and a set of matrices representing a point of $\widehat{{\mathcal{H}}}$ that satisfy the additional constraint
$\widehat{\Lambda}=\widetilde{\Lambda}\;$. (22)
We observe that for a diagonal matrix $\Lambda$, the inverse matrix is in the same adjoint orbit, namely;
$\Lambda^{-1}=\mathfrak{b}^{-1}\Lambda\mathfrak{b}\;,$ (23)
where
$\mathfrak{b}=i\sigma_{2}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\;$.
Assuming (22) we can glue (topologically) the surfaces $\widetilde{{\mathcal{C}}}$ and $\widehat{{\mathcal{C}}}$ along the boundary to get a topological surface $\mathcal{C}$} of genus $g=\widetilde{g}+\widehat{g}$. We denote by $\gamma$ contour on $\mathcal{C}$} placed between $\widetilde{\gamma}$ with $\widehat{\gamma}$. The diagonal monodromy matrix representing $\gamma$ is given by
$\Lambda=\mathrm{diag}(\lambda_{\gamma},\lambda_{\gamma}^{-1})\;,$
such that $\Lambda=\widehat{\Lambda}=\widetilde{\Lambda}$.
[…]
$F:\widetilde{{\mathcal{H}}}\times\widehat{{\mathcal{H}}}\Big|_{\widehat{\Lambda}=\widetilde{\Lambda}}\to\mathcal{V}_{g}\;.$ (24)
Proposition 5.1.
Consider a set of matrices representing a point of $\widetilde{{\mathcal{H}}}$ and a set of matrices representing a point of $\widehat{{\mathcal{H}}}$ such that $\widehat{\Lambda}=\widetilde{\Lambda}$.
Then the following set of $SL(2,{\mathbb{C}})$ matrices
$M_{\alpha_{i}}=\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}M_{\widetilde{\alpha_{i}}}\widetilde{C}\mathfrak{b}\widehat{C}^{-1}\;,\qquad M_{\beta_{i}}=\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}M_{\widetilde{\beta_{i}}}\widetilde{C}\mathfrak{b}\widehat{C}^{-1},\quad i=1,...,\widetilde{g}\;$,
$M_{\alpha_{i+\widetilde{g}}}=M_{\widehat{\alpha_{i}}}\;,\qquad M_{\beta_{i+\widetilde{g}}}=M_{\widehat{\beta_{i}}}\;,\quad i=1,...,\widehat{g}\;$. (25)
defines a point of $\mathcal{V}_{g}$.
Proof.
From the definitions (20) of $\widetilde{{\mathcal{H}}}$ and $\widehat{{\mathcal{H}}}$ one has
[…]
Imposing $\widetilde{\Lambda}=\widehat{\Lambda}$ we get
$\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}\left(\prod_{i=1}^{\widetilde{g}}\left[M_{\widetilde{\alpha}_{i}},M_{\widetilde{\beta_{i}}}\right]\right)\widetilde{C}\mathfrak{b}\widehat{C}^{-1}\left(\prod_{i=1}^{\widehat{g}}\left[M_{\widehat{\alpha}_{i}},M_{\widehat{\beta}_{i}}\right]\right)=\mathbb{I}_{2}\;.$ (27)
Defining $M_{\alpha_{i}}=\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}M_{\widetilde{\alpha}_{i}}\widetilde{C}\mathfrak{b}\widehat{C}^{-1}$ and $M_{\beta_{i}}=\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}M_{\widetilde{\beta_{i}}}\widetilde{C}\mathfrak{b}\widehat{C}^{-1}$ for $i=1,...,\widetilde{g}$, equation (27) can be rewritten as:
$\prod_{i=1}^{\widetilde{g}}\left[M_{\alpha_{i}},M_{\beta_{i}}\right]\prod_{i=1}^{\widehat{g}}\left[M_{\widehat{\alpha}_{i}},M_{\widehat{\beta}_{i}}\right]=\mathbb{I}_{2}\;.$
Now, defining $M_{\alpha_{i+\widetilde{g}}}=M_{\widehat{\alpha}_{i}}$ and $M_{\beta_{i+\widetilde{g}}}=M_{\widehat{\beta_{i}}}$ for $i=1,...,\widehat{g}$, this last identity can be rewritten:
$\prod_{i=1}^{\widetilde{g}}\left[M_{\alpha_{i}},M_{\beta_{i}}\right]\prod_{i=\widetilde{g}+1}^{\widetilde{g}+\widehat{g}}\left[M_{\alpha_{i}},M_{\beta_{i}}\right]=\mathbb{I}_{2}\;,$
which coincides with (9). ∎
Let us now use this plumbing formalism to construct the natural graph $\Gamma_{1}$ on $\mathcal{C}$} equipped with jump matrices on its edges such that the canonical form $\Omega(\Gamma_{1})$ coincides with the Goldman form $\Omega$ on $\mathcal{V}_{g}$.
The graph $\Gamma_{1}$ is constructed by amalgamation of three graphs:
  • the graph $\widetilde{\Gamma}_{1}$ (which can be considered to be the standard graph on $\widetilde{{\mathcal{C}}}$ if the boundary component $\widetilde{\gamma}$ is shrunk to a regular point),
  • the graph $\widehat{\Gamma}_{1}$ (which can be considered to be the standard graph on $\widehat{{\mathcal{C}}}$ if the boundary component $\widehat{\gamma}$ is shrunk to a regular point),
  • and the five-vertex “plumbing” graph $\Gamma_{pl}$.
These graphs can be seen in Fig. 3. The five-valent graph $\Gamma_{pl}$ has one-valent vertices at $\widetilde{v}$, $\widehat{v}$, and three four-valent vertices which we denote by $\widetilde{q}$, $q$ and $\widehat{q}$; these are the points where the segment connecting $\widetilde{v}$ and $\widehat{v}$ crosses $\widetilde{\gamma}$, $\gamma$ and $\widehat{\gamma}$, respectively.
The jump matrices on the edges of $\widetilde{\Gamma}$ are given by matrices $M_{\widetilde{\alpha}_{j}},M_{\widetilde{\beta}_{j}}$; the jump matrices on the edges of $\widehat{\Gamma}$ are given by matrices $M_{\widehat{\alpha}_{j}},M_{\widehat{\beta}_{j}}$ according to the rule (16).
Finally, the jump matrices on the edges of $\Gamma_{pl}$ are given by
$J([\widetilde{v},\widetilde{q}])=J(\widetilde{e}_{1})=M_{\widetilde{\gamma}}\;,\qquad\quad J(\widetilde{e}_{2})=J(\widehat{e}_{2})=\Lambda\;,\qquad J([\widehat{v},\widetilde{q}])=J(\widehat{e}_{1})=M_{\widehat{\gamma}}\;$,
$J(\widetilde{\gamma})=\widetilde{C}\;,\qquad J(\widehat{\gamma})=\widehat{C}\;,\qquad J(\gamma)=\mathfrak{b}\;,$
as shown in Fig. 4 (recall that $M_{\widetilde{\gamma}}$ and $M_{\widehat{\gamma}}$ are defined by (18) and (19) and that the orientation of $\gamma$ is irrelevant because $\mathfrak{b}=\mathfrak{b}^{-1}\equiv\mathfrak{b}$ are the same in $PSL(2,{\mathbb{C}})$).
Figure 3. The graphs $\widetilde{\Gamma}_{1}$, $\Gamma_{pl}$ and $\widehat{\Gamma}_{1}$ for $g=2$.
Figure 4. The amalgamated graph $\Gamma_{1}$ drawn on $\mathcal{C}$} for $g=2$.
One can verify that the local monodromies around all five vertices are trivial. Consider, for example, $q$, $\widetilde{q}$ and $\widetilde{v}$ (the computation at $\widehat{v}$ and $\widehat{q}$ is identical).
  • For $q$: Let us choose the cilium position to be between the edge $q,\widehat{q}$] and the half edge corresponding to $\gamma$ with the same orientation as $q,\widehat{q}$]. The four edges incident to $q$ in counterclockwise order have jump matrices: $J(e_{1})=\mathfrak{b}^{-1}$, $J(e_{2})=\Lambda$, $J(e_{3})=\mathfrak{b}$ and $J(e_{4})=\Lambda$. Then
    $J(e_{1})J(e_{2})J(e_{3})J(e_{4})=\mathfrak{b}^{-1}\Lambda\mathfrak{b}\Lambda=\Lambda^{-1}\Lambda=\mathbb{I}_{2}\;.$
  • For $\widetilde{q}$: Let us choose the cilium position to be between the edge $\tilde{e}=[\widetilde{v},\widetilde{q}$] and the half edge corresponding to $\widetilde{\gamma}$ with the same orientation as $\tilde{e}$. Then we have $J(e_{1})=M_{\widetilde{\gamma}}$, $J(e_{2})=\widetilde{C}$, $J(e_{3})=\Lambda^{-1}$ and $J(e_{4})=\widetilde{C}^{-1}$. Then
    $J(e_{1})J(e_{2})J(e_{3})J(e_{4})=M_{\widetilde{\gamma}}\widetilde{C}\Lambda^{-1}\widetilde{C}^{-1}=M_{\widetilde{\gamma}}M_{\widetilde{\gamma}}^{-1}=\mathbb{I}_{2}\;.$
  • For $\widetilde{v}$: Let us choose the cilium position between the edge $\tilde{e}$ and the half edge corresponding to $\tilde{\beta}_{1}$ with the same orientation as $\tilde{e}$. Then
    $J(e_{1})=M_{\widetilde{\alpha}_{1}}\;,\qquad J(e_{2})=M_{\widetilde{\beta}_{1}}^{-1}\;,\qquad J(e_{3})=M_{\widetilde{\alpha}_{1}}^{-1}\;,\qquad J(e_{4})=M_{\widetilde{\beta}_{1}}\;,$
    $J(e_{5})=M_{\widetilde{\alpha}_{2}},\qquad\dots,\qquad J(e_{4\widetilde{g}+1})=J(\tilde{e})=M_{\widetilde{\gamma}}^{-1}\;.$
    Therefore,
    $\prod_{i=1}^{4\widetilde{g}+1}J(e_{i})=\prod_{i=1}^{\widetilde{g}}\left[M_{\widetilde{\alpha}_{i}},M_{\widetilde{\beta}_{i}}\right]M_{\widetilde{\gamma}}^{-1}=\mathbb{I}_{2}\;.$
The graph $\Gamma_{1}$ can be transformed into the standard graph $\Gamma_{0}$ by a sequence of admissible moves in the sense of [3]:
  1. (1)
    Merge the vertices $\widetilde{v}$ with $\widetilde{q}$ and $\widehat{v}$ with $\widehat{q}$.
  2. (2)
    Merge the vertices $q$ , $\widetilde{q}$ and $\widehat{q}$.
  3. (3)
    Zip together the three edges $\gamma$, $\widetilde{\gamma}$ and $\widehat{\gamma}$ whose jump matrices are respectively $\mathfrak{b}$, $\widetilde{C}^{-1}$ and $\widehat{C}$ to get a single edge $\gamma_{0}$ with jump matrix $J(\gamma_{0})=\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}$.
  4. (4)
    Zip the edge $\gamma_{0}$ with the edges $\{\widetilde{\alpha}_{i},\widetilde{\beta}_{i}\}_{i=1}^{\widetilde{g}}$ to get edges $\{\alpha_{i},\beta_{i}\}_{i=1}^{\widetilde{g}}$ with corresponding jump matrices:
    $J(\beta_{i})=M_{\alpha_{i}}=\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}M_{\widetilde{\alpha}_{i}}\widetilde{C}\mathfrak{b}\widehat{C}^{-1}\quad J(\alpha_{i})=M_{\beta_{i}}=\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}M_{\widetilde{\beta}_{i}}\widetilde{C}\mathfrak{b}\widehat{C}^{-1},\quad i=1,...,\widetilde{g}$.
Therefore, the canonical symplectic form $\Omega(\Gamma_{1})$ coincides with the canonical symplectic form $\Omega(\Gamma_{0})$. In other words, one can use the graph $\Gamma_{1}$ to compute the Goldman symplectic form $\Omega$ on $\mathcal{V}_{g}$.
In the next section we construct another graph $\Gamma_{2}$ which can also be transformed to $\Gamma_{1}$ by a sequence of standard moves; the jump matrices on edges of $\Gamma_{2}$ will be expressed in terms of Fock-Goncharov coordinates, and, therefore, can be used to write $\Omega$ in the log-canonical form.
6. Triangulation of $\widetilde{{\mathcal{C}}}$ and $\widehat{{\mathcal{C}}}$ and log-canonical coordinates
6.1. Graph $\Gamma_{2}$ from triangulations of $\widetilde{{\mathcal{C}}}$ and $\widehat{{\mathcal{C}}}$
For explicit computation of the Weil-Petersson form in terms of log-canonical coordinates we are going to construct one more graph on $\mathcal{C}$}, which we denote by $\Gamma_{2}$, with jump matrices on its edges which make it equivalent to $\Gamma_{1}$ and $\Gamma_{0}$.
The graph $\Gamma_{2}$ and jump matrices on its edges are constructed via amalgamation of the following three graphs
  1. (1)
    Graph $\widetilde{\Gamma}_{2}$ constructed as follows. Fix an oriented triangulation (i.e. such that an orientation is assigned to each edge) $\widetilde{\Sigma}$ of $\widetilde{{\mathcal{C}}}$ with a single vertex $\widetilde{v}$ such that the boundary $\gamma$ and contour $\widetilde{\gamma}$ fall into one of the triangles. The number of edges of $\widetilde{\Sigma}$ equals $6\widetilde{g}-3$. The jump matrices on the edges $\widetilde{e}_{j}$ of $\widetilde{\Sigma}$ are chosen as follows
    $J(\widetilde{e}_{j})=\widetilde{{S}}_{e_{j}}=\begin{pmatrix}0&1/\widetilde{z}_{e_{j}}\\ -\widetilde{z}_{e_{j}}&0\\ \end{pmatrix},\quad\widetilde{z}_{e_{j}}\in{\mathbb{C}}\;.$
    Note that $\widetilde{S}_{e_{j}}^{-1}=-\widetilde{S}_{e_{j}}=\widetilde{S}_{-e_{j}}$. We also denote
    $\widetilde{\zeta}_{e_{j}}=\ln{\widetilde{z}_{e_{j}}}\;,$
    for $j=1,...,6\widetilde{g}-3$.
    Choose on each face $\widetilde{F}_{k}$ of $\widetilde{\Sigma}$ an extra internal vertex $\widetilde{w}_{k}$ connected to $\widetilde{v}$ by three internal edges oriented toward $\widetilde{w}_{k}$. To each internal edge (shown in red in Fig. 5 and Fig. 6) assign the constant jump matrix:
    $A=\begin{pmatrix}0&-1\\ 1&-1\\ \end{pmatrix}\;$.
  2. (2)
    Graph $\widehat{\Gamma}_{2}$ constructed as follows. Fix an oriented triangulation $\widehat{\Sigma}$ of $\widehat{{\mathcal{C}}}$ with a single vertex $\widehat{v}$ such that the boundary $\gamma$ and contour $\widehat{\gamma}$ fall into one of the triangles. The number of edges of $\widehat{\Sigma}$ equals $6\widehat{g}-3$. The jump matrices on edges $\widehat{e}_{i}$ of $\widehat{\Sigma}$ are chosen as follows
    $J(\widehat{e}_{i})=\widehat{S}_{e_{i}}=\begin{pmatrix}0&1/\widehat{z}_{e_{i}}\\ -\widehat{z}_{e_{i}}&0\\ \end{pmatrix},\quad\ \widehat{z}_{e_{i}}\in{\mathbb{C}}\;,$
    and denote $\widehat{\zeta}_{e_{i}}=\ln{\widehat{z}_{e_{i}}}$, for $i=1,...,6\widehat{g}-3$. Introduce on each face of $\widehat{\Sigma}$ a point $\widehat{w}_{k}$ together with three internal edges linking it to $\widehat{v}$. Assign the jump matrix $A$ to each of these internal edges.
    The variables $\widehat{\zeta}$ and $\widetilde{\zeta}$ are assumed to satisfy the constraint
    $\sum_{j=1}^{6\widetilde{g}-3}\widetilde{\zeta}_{e_{j}}=\sum_{i=1}^{6\widehat{g}-3}\widehat{\zeta}_{e_{i}}$ (28)
    or, equivalently,
    $\prod_{j=1}^{6\widetilde{g}-3}\widetilde{z}_{e_{j}}=\prod_{i=1}^{6\widehat{g}-3}\widehat{z}_{e_{i}}\;.$ (29)
  3. (3)
    The “plumbing” 5-vertex graph part $\Gamma_{pl}$ remains the same while the jump matrices on its edges are defined from the variables $\widetilde{\zeta}_{e_{j}}$ and $\widehat{\zeta}_{e_{j}}$ as follows.
[…]
    $M_{\widetilde{\gamma}}=\left(\prod_{l=1}^{n_{\widetilde{v}}}A\widetilde{S}_{e_{l}}\right)^{-1}=\begin{pmatrix}\lambda_{\widetilde{\gamma}}&0\\ *&\lambda_{\widetilde{\gamma}}^{-1}\end{pmatrix}\;$.
    where
    $\lambda_{\widetilde{\gamma}}=\prod_{j=1}^{6\widetilde{g}-3}\widetilde{z}^{2}_{e_{j}}\;.$ (30)
    Choose the special matrix $\widetilde{C}$ of eigenvectors of $M_{\widetilde{\gamma}}$ having the form
    $\widetilde{C}=\begin{pmatrix}1&0\\ *&1\end{pmatrix}\;$.
    Then the general $SL(2,{\mathbb{C}})$ matrix $\widetilde{C^{\prime}}$ diagonalizing $M_{\widetilde{\gamma}}$ (such that $M_{\widetilde{\gamma}}=\widetilde{C^{\prime}}\widetilde{\Lambda}\widetilde{C^{\prime}}^{-1}$) has the form
    $\widetilde{C^{\prime}}=\widetilde{C}\begin{pmatrix}b_{\widetilde{\gamma}}^{-1}&0\\ 0&b_{\widetilde{\gamma}}\end{pmatrix}\;$, (31)
    where the complex number $b_{\widetilde{\gamma}}$ is called the toric variable. The corresponding logarithmic coordinate is denoted $\beta_{\widetilde{\gamma}}=\ln{b_{\widetilde{\gamma}}}$.
    Similarly, we define the matrix
    $M_{\widehat{\gamma}}=\left(\prod_{l=1}^{n_{\widehat{v}}}A\widehat{S}_{e_{l}}\right)^{-1}=\begin{pmatrix}\lambda_{\widehat{\gamma}}&0\\ *&\lambda_{\widehat{\gamma}}^{-1}\end{pmatrix}\;$,
    with
    $\lambda_{\widehat{\gamma}}=\prod_{j=1}^{6\widehat{g}-3}\widehat{z}^{2}_{e_{j}}\;,$ (32)
    the matrix $\widehat{C}$ of the form
    $\widehat{C}=\begin{pmatrix}1&0\\ *&1\end{pmatrix}\;$,
    and the general matrix diagonalizing $M_{\widehat{\gamma}}$ of the form
    $\widehat{C^{\prime}}=\widehat{C}\begin{pmatrix}b_{\widehat{\gamma}}^{-1}&0\\ 0&b_{\widehat{\gamma}}\end{pmatrix}\;$, (33)
    where $b_{\widehat{\gamma}}$ is another toric variable; we denote $\beta_{\widehat{\gamma}}=\ln b_{\widehat{\gamma}}$.
Note that the necessary condition
$\lambda_{\widehat{\gamma}}=\lambda_{\widetilde{\gamma}}$
is the corollary of the constraint (28).
In the sequel we shall drop the decorations and denote
$\lambda_{\widetilde{\gamma}}=\lambda_{\gamma}=\lambda_{\widehat{\gamma}}\;$.
Define also
$\ell_{\gamma}=\ln\lambda_{\gamma}\;.$
An example of graph $\Gamma_{2}$ for genus 2 surface $\mathcal{C}$} is shown in Fig. 5.
Figure 5. The graph $\Gamma_{2}$ drawn on $\mathcal{C}$} for $g=2$.
Another representation of $\Gamma_{2}$ with explicitly shown triangulation is given schematically in Fig. 6. To get from Fig. 5 to Fig. 6 we cut the genus two surface $\mathcal{C}$} along the contour $\gamma$ (shown in blue) positioned ”in the middle” between $\widetilde{\gamma}$ and $\widehat{\gamma}$. Then the edge $e$ connecting $\widetilde{q}$ and $\widehat{q}$ is cut in half, leaving the half-edges inside the small circles in both genus one ”halfs” shown in Fig. 6.
Figure 6. Triangulations of genus one components $\widetilde{{\mathcal{C}}}$ and $\widehat{{\mathcal{C}}}$ for $g=2$. The segments inside of the circles represent the two edges $\widetilde{q},q]$ and $q,\widehat{q}$] in Fig. 5.
The graph $\Gamma_{2}$ can be transformed into the graph $\Gamma_{1}$ by the following sequence of standard moves:
  1. (1)
    First, merge the internal vertices $\widetilde{w}_{k}$ with $\widetilde{v}$ and $\widehat{w}_{k}$ with $\widehat{v}$.
  2. (2)
    Second, zip together the edges with jump matrices $A$ and the edges with jump matrices $\widetilde{S}_{e_{i}}$ (resp. $\widehat{S}_{e_{i}}$) to get edges with jump matrices $A\widetilde{S}_{e_{i}}$ or $A^{-1}\widetilde{S}_{e_{i}}$ (resp. $A\widehat{S}_{e_{i}}$ or $A^{-1}\widehat{S}_{e_{i}}$).
  3. (3)
    Third, zip together the edges of the triangulations $\widetilde{\Sigma}$ and $\widehat{\Sigma}$ to get the set of edges $\{\widetilde{\alpha}_{i},\widetilde{\beta}_{i}\}_{i=1}^{\widetilde{g}}$ and $\{\widehat{\alpha}_{i},\widehat{\beta}_{i}\}_{i=1}^{\widehat{g}}$. The jump matrices
    $\{M_{\widetilde{\alpha}_{j}}\;,M_{\widetilde{\beta}_{j}}\;,M_{\widehat{\alpha}_{j}}\;,M_{\widehat{\beta}_{j}}\}$ (34)
    on these new edges become then some products of $\widetilde{S}_{e_{i}}$ $\widehat{S}_{e_{i}}$, $A$ and $A^{-1}$.
The Fuchsian monodromy matrices $M_{\alpha_{j}}$, $M_{\beta_{j}}$ can be obtained from matrices (34) using formulas (25) with $\widehat{C}$ and $\widetilde{C}$ replaced by the general diagonalizing matrices $\widetilde{C^{\prime}}$ and $\widehat{C^{\prime}}$ (31) and (33).
The matrix
$\widetilde{C^{\prime}}^{-1}\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}\widehat{C^{\prime}}=\widetilde{C}^{-1}\begin{bmatrix}0&-b_{\widetilde{\gamma}}\,b_{\widehat{\gamma}}\\ \left(b_{\widetilde{\gamma}}\,b_{\widehat{\gamma}}\right)^{-1}&0\end{bmatrix}\widehat{C}$ (35)
depends only on the product of the toric variables; thus the Fuchsian monodromies depend only on the sum of their logarithms and we can define the new variable
[…]
Note that the variables $\widetilde{\zeta}_{e_{j}}$ and $\widehat{\zeta}_{e_{j}}$ are not independent but satisfy the constraint (28). The independent set of variables can be chosen as follows:
$\left\{\{\widetilde{\zeta}_{e_{i}}\}_{i=2}^{6\widetilde{g}-3},\,\{\widehat{\zeta}_{e_{i}}\}_{i=2}^{6\widehat{g}-3},\ell_{\gamma},\beta_{\gamma}\right\}\;.$ (37)
Then the remaining coordinates $\widetilde{\zeta}_{1}$ and $\widehat{\zeta}_{1}$ can be obtained from the logarithmic version of (32):
$\widetilde{\zeta}_{e_{1}}=\frac{1}{2}\ell_{\gamma}-\sum_{j=2}^{6\widetilde{g}-3}\widetilde{\zeta}_{e_{j}}\;,\qquad\widehat{\zeta}_{e_{1}}=\frac{1}{2}\ell_{\gamma}-\sum_{j=2}^{6\widehat{g}-3}\widehat{\zeta}_{e_{j}}\;.$ (38)
7. Computation of Goldman symplectic form as $\Omega(\Gamma_{2})$
We are in a position to state our main theorem.
Theorem 7.1.
The Goldman form $\Omega$ on $\mathcal{V}_{g}$ has constant coefficients in terms of variables (37) i.e. these variables are log-canonical for $\Omega$. Namely, we have
$\Omega=\Omega_{0}+\widetilde{\Omega}_{1}+\widehat{\Omega}_{1}+\widetilde{\Omega}_{2}+\widehat{\Omega}_{2}\;,$ (39)
where
$\Omega_{0}=\mathrm{d}\beta_{\gamma}\wedge\mathrm{d}\ell_{\gamma}\;,$
$\widetilde{\Omega}_{1}=\sum_{\begin{subarray}{c}i,j=2\\ i<j\end{subarray}}^{6\widetilde{g}-3}\left(\widetilde{c}_{ij}-\widetilde{c}_{1j}+\widetilde{c}_{1i}\right)\mathrm{d}\widetilde{\zeta}_{e_{i}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{j}}\;,\qquad\widehat{\Omega}_{1}=\sum_{\begin{subarray}{c}k,l=2\\ k<l\end{subarray}}^{6\widehat{g}-3}\left(\widehat{c}_{kl}-\widehat{c}_{1l}+\widehat{c}_{1k}\right)\mathrm{d}\widehat{\zeta}_{e_{k}}\wedge\mathrm{d}\widehat{\zeta}_{e_{l}}\;,$
$\widetilde{\Omega}_{2}=\frac{1}{2}\sum_{m=2}^{6\widetilde{g}-3}\widetilde{c}_{1m}\mathrm{d}\ell_{\gamma}\wedge\mathrm{d}\widetilde{\zeta}_{e_{m}}\;,\qquad\widehat{\Omega}_{2}=\frac{1}{2}\sum_{n=2}^{6\widehat{g}-3}\widehat{c}_{1n}\mathrm{d}\ell_{\gamma}\wedge\mathrm{d}\widehat{\zeta}_{e_{n}}\;,$
The integer coefficients $\widetilde{c}_{ij}$, $\widehat{c}_{kl}$ lie in the interval between 4-4 and $4$. They are determined by the order of the edges of the uni-valent graphs $\widetilde{\Sigma}$ and $\widehat{\Sigma}$ at their vertices.
Proof.
The symplectic form $\Omega({\Gamma_{2}})$ can be computed using the general formula (14) by summing the contributions of all four vertices $\widetilde{v}$, $\widehat{v}$, $\widetilde{q}$ and $\widehat{q}$ of the graph $\Gamma_{2}$.
The contributions coming from the vertices $\widetilde{q}$, $\widehat{q}$ and $q$ are given by the following lemma:
Lemma 7.2.
The contribution of the vertex $\widetilde{q}$ to $\Omega({\Gamma_{2}})$ equals to $\Omega_{\widetilde{q}}=2\mathrm{d}\beta_{\widetilde{\gamma}}\wedge\mathrm{d}\ell_{\gamma}$, the contribution of the vertex $\widehat{q}$ equals to $\Omega_{\widehat{q}}=2\mathrm{d}\beta_{\widehat{\gamma}}\wedge\mathrm{d}\ell_{\gamma}$ and the contribution of the vertex $q$ equals 0.
Choose the cilium position to be between the edge $\widetilde{v},\widetilde{q}$] and the half edge corresponding to $\widetilde{\gamma}$ with the same orientation as $\widetilde{v},\widetilde{q}$]. The four edges incident to $\widetilde{q}$ in counterclockwise order have jump matrices: $J(e_{1})=J_{1}=M_{\widetilde{\gamma}}$, $J(e_{2})=J_{2}=\widetilde{C^{\prime}}$, $J(e_{3})=J_{3}=\Lambda^{-1}$ and $J(e_{4})=J_{4}=\widetilde{C^{\prime}}^{-1}$.
According to equation (14) the contribution of $\widetilde{q}$ to $\Omega({\Gamma_{2}})$ reads:
$\Omega_{\widetilde{q}}$ $=\frac{1}{2}\sum_{l=1}^{3}\mathrm{tr}\left(J_{[1...l]}^{-1}{\mathrm{d}}J_{[1...l]}\wedge J_{l}^{-1}{\mathrm{d}}J_{l}\right)$
$=\frac{1}{2}\mathrm{tr}\left((J_{1}J_{2})^{-1}{\mathrm{d}}(J_{1}J_{2})\wedge J_{2}^{-1}{\mathrm{d}}J_{2}\right)+\frac{1}{2}\mathrm{tr}\left((J_{1}J_{2}J_{3})^{-1}{\mathrm{d}}(J_{1}J_{2}J_{3})\wedge J_{3}^{-1}{\mathrm{d}}J_{3}\right)$
$=\frac{1}{2}\mathrm{tr}\left((M_{\widetilde{\gamma}}\widetilde{C^{\prime}})^{-1}{\mathrm{d}}(M_{\widetilde{\gamma}}\widetilde{C^{\prime}})\wedge\widetilde{C^{\prime}}^{-1}{\mathrm{d}}\widetilde{C^{\prime}}\right)+\frac{1}{2}\mathrm{tr}\left(\widetilde{C^{\prime}}^{-1}{\mathrm{d}}\widetilde{C^{\prime}}\wedge\Lambda{\mathrm{d}}(\Lambda^{-1})\right)$
[…]
Using the identity $M_{\widetilde{\gamma}}^{-1}{\mathrm{d}}M_{\widetilde{\gamma}}=\widetilde{C^{\prime}}\Lambda^{-1}\widetilde{C^{\prime}}^{-1}\left({\mathrm{d}}\widetilde{C^{\prime}}+{\mathrm{d}}\Lambda+{\mathrm{d}}(\widetilde{C^{\prime}}^{-1})\right)$ we see that the first term is equal to the second one, so the contribution $\Omega_{\widetilde{q}}$ simplifies to:
[…]
Using the definition (31) of $\widetilde{C^{\prime}}$ we get:
$\Omega_{\widetilde{q}}=\mathrm{tr}\left(\Lambda^{-1}{\mathrm{d}}\Lambda\wedge\widetilde{C^{\prime}}^{-1}{\mathrm{d}}\widetilde{C^{\prime}}\right)=-2\frac{\mathrm{d}\lambda_{\gamma}}{\lambda_{\gamma}}\wedge\frac{\mathrm{d}b_{\widetilde{\gamma}}}{b_{\widetilde{\gamma}}}=2\mathrm{d}\beta_{\widetilde{\gamma}}\wedge\mathrm{d}\ell_{\gamma}\;.$
Similarly,
$\Omega_{\widehat{q}}=-2\frac{\mathrm{d}\lambda_{\gamma}}{\lambda_{\gamma}}\wedge\frac{\mathrm{d}b_{\widehat{\gamma}}}{b_{\widehat{\gamma}}}=2\mathrm{d}\beta_{\widehat{\gamma}}\wedge\mathrm{d}\ell_{\gamma}\;.$
The four edges incident to $q$ have jump matrices $\mathfrak{b}^{-1}$, $\Lambda$, $\mathfrak{b}$ and $\Lambda$. Since the matrix $\mathfrak{b}$ is constant by immediate consequence of equation (14) the contribution of $q$ to $\Omega$ is zero. ∎
To compute the contributions of the vertex $\widetilde{v}$ we can use the result of [3] which states that
$\Omega_{\widetilde{v}}=\sum_{\begin{subarray}{c}i,j=1\\ e_{i}<e_{j}\end{subarray}}^{2(6\widetilde{g}-3)}\mathrm{d}\widetilde{\zeta}_{e_{i}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{j}}\;,$
where each edge comes to $\widetilde{v}$ twice. Eliminating repetitions of similar terms we get
$\Omega_{\widetilde{v}}=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{6\widetilde{g}-3}\widetilde{c}_{ij}\mathrm{d}\widetilde{\zeta}_{e_{i}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{j}}\;,$ (40)
where $\widetilde{c}_{ij}$ are some (depending on the order of the edges at the vertex) integers from the range $\{-4,...,4\}$.
Similarly,
$\Omega_{\widehat{v}}=\sum_{\begin{subarray}{c}k,l=1\\ k<l\end{subarray}}^{6\widehat{g}-3}\widehat{c}_{kl}\mathrm{d}\widehat{\zeta}_{e_{k}}\wedge\mathrm{d}\widehat{\zeta}_{e_{l}}\;,$ (41)
where $\widehat{c}_{kl}$ are also some integers from the same set $\{-4,...,4\}$.
If one eliminates $\widetilde{\zeta}_{1}$ and $\widehat{\zeta}_{1}$ using the constraints (38) we get also
$\begin{split}\Omega_{\widetilde{v}}&=\sum_{\begin{subarray}{c}i,j=2\\ i<j\end{subarray}}^{6\widetilde{g}-3}\left(\widetilde{c}_{ij}-\widetilde{c}_{1j}+\widetilde{c}_{1i}\right)\mathrm{d}\widetilde{\zeta}_{e_{i}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{j}}+\frac{1}{2}\sum_{m=2}^{6\widetilde{g}-3}\widetilde{c}_{1m}\mathrm{d}\ell_{\gamma}\wedge\mathrm{d}\widetilde{\zeta}_{e_{m}}\\ \Omega_{\widehat{v}}&=\sum_{\begin{subarray}{c}k,l=2\\ k<l\end{subarray}}^{6\widehat{g}-3}\left(\widehat{c}_{kl}-\widehat{c}_{1l}+\widehat{c}_{1k}\right)\mathrm{d}\widehat{\zeta}_{e_{k}}\wedge\mathrm{d}\widehat{\zeta}_{e_{l}}+\frac{1}{2}\sum_{n=2}^{6\widehat{g}-3}\widehat{c}_{1n}\mathrm{d}\ell_{\gamma}\wedge\mathrm{d}\widehat{\zeta}_{e_{n}}\end{split}$
Summing up the contributions of all four vertices to $\Omega(\Gamma_{2})$ we get (39).
8. Generalization to multiple cutting contours
A system of log-canonical coordinates generalizing (3) can be naturally associated to any system of $m$ closed nonintersecting geodesics $\gamma_{1},\dots\gamma_{m}$ for any $m=1,\dots,3g-3$. These contours split $\mathcal{C}$} into one or more Riemann surfaces with boundaries which we denote by ${\mathcal{C}}^{(1)},\dots\mathcal{C}^{(n)}$. Denote the number of boundary components of ${\mathcal{C}}^{(i)}$ by $k^{(i)}$ and the genus by $g^{(i)}$; we have $\sum_{i=1}^{n}k^{(i)}=2m$. The boundary components of each ${\mathcal{C}}^{(i)}$ are denoted $\gamma_{j}^{(i)}$, $j=1,...,k^{(i)}$, see Figure 7.
Figure 7. An example of multiple contours for $m=4$ and $g=5$.
Then we can choose two points $v_{k}^{\pm}$ near the opposite sides of the curve $\gamma_{k}$. Consider now a fat-graph $\Sigma^{(i)}$ with $k^{(i)}$ vertices $v_{k}^{\pm}\in{\mathcal{C}}^{(i)}$ that provides triangulation of each ${\mathcal{C}}^{(i)}$ if the boundary components shrink to a point. The number of edges of $\Sigma^{(i)}$ equals $6g^{(i)}-6+3k^{(i)}$, which coincides with the dimension of the moduli space of Riemann surfaces of genus $g^{(i)}$ with $k^{(i)}$ boundary components of variable length.
We associate complex shear coordinates $\zeta_{e_{j}}^{(i)}$, $j=1,\dots,6g^{(i)}-6+3k^{(i)}$ to the edges of $\Sigma^{(i)}$; for each $e\in\Sigma^{(i)}$ we simply denote the coordinate by $\zeta_{e}$. Adding internal vertices and edges we get a graph $\Gamma^{(i)}$ on each ${\mathcal{C}}^{(i)}$. The graph in $\mathcal{C}$} obtained by amalgamating all graphs $\Gamma^{(i)}$, in a manner similar to that of Section 3.6, is denoted by $\Gamma$.
There are $m$ constraints that relate the oriented lengths $\ell_{\gamma_{k}}$ of $\gamma_{k}$ to $\zeta_{e}$’s
$\sum_{e\in E(\Sigma^{(i)}),\;e\perp v_{k}}\zeta_{e}=\ell_{\gamma_{k}}\;,\qquad k=1,\dots m\;,$ (42)
where some terms on the left-hand side come with coefficient 1, and some other terms with coefficient 2 (depending on whether the edge “comes back” or not).
The system of constraints (42) is linear in terms of $\zeta_{e}$. The system of log-canonical coordinates contains “complex lengths” $\ell_{\gamma_{k}}$ for $k=1,\dots,m$ and the shear-type variables $\zeta_{e}$ (there are linear constraints (42) relating $\ell_{\gamma_{k}}$ to $\zeta_{e}$’s).
Theorem 8.1.
The Goldman form $\Omega$ can be represented as
$\Omega=\sum_{i=1}^{n}\Omega^{(i)}+\sum_{j=1}^{m}{\mathrm{d}}\beta_{\gamma_{j}}\wedge{\mathrm{d}}\ell_{\gamma_{j}}\;,$ (43)
where
$\Omega^{(i)}=\sum_{v\in V(\Sigma^{(i)})}\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v\\ e<\widetilde{e}\end{subarray}}{\mathrm{d}}\zeta_{e}\wedge{\mathrm{d}}\zeta_{\widetilde{e}}\;$. (44)
Proof.
The proof is by direct computation of $\Omega$ as $\Omega(\Gamma)$, which is a straightforward generalization of the one given in the proof of Theorem 7.1. ∎
Remark 8.2.
In (44) some terms repeat and some coordinates $\zeta_{e}$ should also be expressed by a linear combination of the remaining ones and $\ell_{\gamma_{k}}$ due to constraints (42).
9. Complete trinion decomposition
Let us discuss in detail the case of complete trinion decomposition. This decomposition is obtained by splitting $\mathcal{C}$} along a system $S$ of m=3g3m=3g-3 closed non-intersecting geodesics $\gamma_{1},...,\gamma_{3g-3}$ to obtain 2g22g-2 trinions -that is, spheres with three boundary components- denoted by $\mathcal{T}^{(j)}$.
Consider a single trinion $\mathcal{T}$. A triangulation on $\mathcal{T}$ is shown in Figure 8.
Figure 8. Triangulation and monodromies associated to a single trinion $\mathcal{T}$.
The triangulation in Figure 8 has edges ordered as follows:
$\begin{split}&e_{1},e_{2}\perp v_{1},\quad e_{2}<e_{1}\;,\\ &e_{1},e_{3}\perp v_{2},\quad e_{1}<e_{3}\;,\\ &e_{2},e_{3}\perp v_{3},\quad e_{3}<e_{2}\;.\end{split}$ (45)
The boundary components near the vertices $v_{1}$, $v_{2}$, $v_{3}$ are denoted by $\gamma_{1}$, $\gamma_{2}$, $\gamma_{3}$, respectively. The induced cyclic ordering of the boundary components, based on the edge ordering (45), is:
$\left(\gamma_{3},\gamma_{2},\gamma_{1}\right)\;.$ (46)
Taking the base point $p$ inside the black triangle (Figure 8), the local $SL(2,\mathbb{C})$ monodromy matrices around $v_{j}$ are denoted by $M_{j}$, with complex eigenvalues $\lambda_{j}$ and $1/\lambda_{j}$, $j=1,2,3$. The shear-type coordinates $z_{e_{1}}$, $z_{e_{2}}$ and $z_{e_{3}}$ can be expressed in terms of $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ as:
$z_{e_{1}}=\sqrt{\frac{\lambda_{3}}{\lambda_{2}\lambda_{1}}}\;,\qquad z_{e_{2}}=\sqrt{\frac{\lambda_{2}}{\lambda_{1}\lambda_{3}}}\;,\qquad z_{e_{3}}=\sqrt{\frac{\lambda_{1}}{\lambda_{3}\lambda_{2}}}\;.$ (47)
Then a representation $\rho:\pi_{1}(\mathcal{T},p)\rightarrow SL(2,\mathbb{C})$ is explicitly given by the following three monodromies:
$M_{1}=\left[\begin{array}[]{cc}-\lambda_{1}&0\\ \frac{\lambda_{1}\lambda_{2}+\lambda_{3}}{\lambda_{1}\lambda_{3}}&-\frac{1}{\lambda_{1}}\end{array}\right]\;,\hskip 19.91684ptM_{2}=\left[\begin{array}[]{cc}\frac{\lambda_{3}}{\lambda_{1}}&\frac{-\lambda_{2}\lambda_{3}-\lambda_{1}}{\lambda_{1}\lambda_{2}}\\ \frac{\lambda_{2}\lambda_{1}+\lambda_{3}}{\lambda_{1}}&-\frac{\lambda_{1}\lambda_{2}^{2}+\lambda_{3}\lambda_{2}+\lambda_{1}}{\lambda_{1}\lambda_{2}}\end{array}\right]\;,\hskip 19.91684ptM_{3}=\left[\begin{array}[]{cc}-\frac{1}{\lambda_{3}}&\frac{-\lambda_{2}\lambda_{3}-\lambda_{1}}{\lambda_{2}}\\ 0&-\lambda_{3}\end{array}\right]\;.$ (48)
These monodromies satisfy the fundamental relation in homotopy:
$M_{1}M_{2}M_{3}=\mathbb{I}_{2}\;,\hskip 19.91684pt\mathrm{Tr}(M_{j})=-\lambda_{j}-\frac{1}{\lambda_{j}}\;,\hskip 19.91684ptj=1,2,3\;.$
For an embedded trinion $\mathcal{T}^{(j)}$, the complex lengths of its boundary components must match the lengths of three geodesics in the system $S$. This matching imposes additional constraints.
The coordinates $\{\ell_{\gamma_{j}},\beta_{\gamma_{j}}\}_{j=1}^{3g-3}$ form a set of log-canonical coordinates on $\mathcal{V}_{g}$. The exact expression of the Goldman form $\Omega$ in those coordinates depends on the chosen trinion decomposition. Nevertheless, a general expression can be written using the “trinion graph” corresponding to a complete trinion decomposition.
Definition 9.1.
The trinion graph $\Gamma_{trin}$ of a complete trinion decomposition is the graph whose vertices are individual trinions. The edges of $\Gamma_{trin}$ represent the boundary components where two trinions are glued (or a closed handle for a loop edge).
Figure 9. Gluing of a trinion $\mathcal{T}^{(j)}$ with two other trinions $\mathcal{T}^{(j-1)}$ and $\mathcal{T}^{(j+1)}$ -with graph $\Gamma^{(j)}$ drawn on $\mathcal{T}^{(j)}$- (left) and associated edges and vertices of the trinion graph (right).
Figure 9, right, shows a vertex $v^{(j)}$ of the trinion graph. It corresponds to a trinion $\mathcal{T}^{(j)}$, which is glued along its boundaries to two other trinions $\mathcal{T}^{(j-1)}$ and $\mathcal{T}^{(j+1)}$.
Let the boundary components of $\mathcal{T}^{(j)}$ be $\gamma_{1}^{(j)}$, $\gamma_{2}^{(j)}$ and $\gamma_{3}^{(j)}$, ordered as (46). Figure 9, left, shows that $\gamma_{1}^{(j)}$ lies in a collar neighborhood of $\gamma_{m}$, $\gamma_{2}^{(j)}$ lies near $\gamma_{r}$ and $\gamma_{3}^{(j)}$ lies near $\gamma_{n}$, where $\gamma_{n},\gamma_{r},\gamma_{m}\in S$ and $n,r,m\in\{1,2,...,3g-3\}$. This induces the following ordering of the geodesics:
$(\gamma_{n},\gamma_{r},\gamma_{m})\;.$ (49)
This is equivalent to the following ordering of the edges $e_{n}$, $e_{r}$ and $e_{m}$ incident to the vertex $v^{(j)}$ of $\Gamma_{trin}$:
$(e_{n},e_{r},e_{m})\;.$ (50)
With the edge ordering (50) fixed, we formulate the following theorem.
Theorem 9.2.
Let $\Gamma_{trin}$ be the trinion graph of a complete trinion decomposition. Then the Goldman form on $\mathcal{V}_{g}$ can be expressed as:
$\begin{split}\Omega=\frac{1}{2}\sum_{v\in V(\Gamma_{trin})}\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v\\ e<\widetilde{e}\end{subarray}}\mathrm{d}\ell_{e}\wedge\mathrm{d}\ell_{\widetilde{e}}+\sum_{e\in E(\Gamma_{trin})}\mathrm{d}\beta_{e}\wedge\mathrm{d}\ell_{e}\;,\end{split}$ (51)
where the 3g33g-3 variables $\ell_{e}$ are the “complex oriented lengths” of the edges of the trinion graph, and $\beta_{e}$’s are the conjugated variables.
Proof.
Let $\{\lambda_{i}^{(j)},b_{i}^{(j)}\}_{i=1,2,3}$ be the eigenvalues and toric variables associated with the jthj-th trinion $\mathcal{T}^{(j)}$. Their logarithmic counterparts are $\{\ell_{i}^{(j)},\beta_{i}^{(j)}\}_{i=1,2,3}$. These variables are pairwise duplicated due to trinion gluing. Applying (17) to the monodromies (48) yields the contribution of $\mathcal{T}^{(j)}$ to the symplectic form $\Omega$:
$\Omega_{\mathcal{T}^{(j)}}=\sum_{i=1}^{3}\left(\mathrm{d}\ell_{i+1}^{(j)}\wedge\mathrm{d}\ell_{i}^{(j)}+2\mathrm{d}\beta_{i}^{(j)}\wedge\mathrm{d}\ell_{i}^{(j)}\right),$ (52)
where the summation index $i$ taken modulo 3. Importantly, the expression (52) is independent of the trinion configuration and is compatible with the ordering (46).
Let $v^{(j)}$ be a vertex of $\Gamma_{trin}$ corresponding to $\mathcal{T}^{(j)}$. The three gluing constraints (for the configuration shown on Figure 9) imply:
$\ell_{3}^{(j)}=\ell_{3}^{(j+1)}=\ell_{e_{n}}\;,\qquad\ell_{2}^{(j)}=\ell_{2}^{(j-1)}=\ell_{e_{r}}\;,\qquad\quad\ell_{1}^{(j)}=\ell_{1}^{(j-1)}=\ell_{e_{m}}\;.$ (53)
The corresponding toric variables are:
$\beta_{e_{n}}=2\left(\beta_{3}^{(j)}+\beta_{3}^{(j+1)}\right)\;,\qquad\quad\beta_{e_{r}}=2\left(\beta_{2}^{(j)}+\beta_{2}^{(j-1)}\right)\;,\qquad\quad\beta_{e_{m}}=2\left(\beta_{1}^{(j)}+\beta_{1}^{(j-1)}\right)\;.$
These constraints correspond specifically to $\mathcal{T}^{(j)}$ glued to $\mathcal{T}^{(j-1)}$ and $\mathcal{T}^{(j+1)}$.
For a loop edge $e_{p}$ incident to a vertex $v^{(q)}$ (corresponding to a “closed handle”), the constraints become:
$l_{1}^{(q)}=l_{2}^{(q)}=l_{e_{p}}\;,\qquad\beta_{e_{p}}=2\left(\beta_{1}^{(q)}+\beta_{2}^{(q)}\right)\;.$
For a trinion $P^{(s)}$ glued to three distinct trinions $\mathcal{T}^{(s-2)}$, $\mathcal{T}^{(s-1)}$ and $\mathcal{T}^{(s+1)}$, one possible choice of constraints is:
$\ell_{3}^{(s)}=\ell_{1}^{(s+1)}=\ell_{e_{t}}\;,\qquad\ell_{2}^{(s)}=\ell_{3}^{(s-1)}=\ell_{e_{v}}\;,\qquad\ell_{1}^{(s)}=\ell_{3}^{(s-2)}=\ell_{e_{w}}\;.$
The corresponding toric variables are:
$\beta_{e_{t}}=2\left(\beta_{3}^{(s)}+\beta_{1}^{(s+1)}\right)\;,\qquad\beta_{e_{v}}=2\left(\beta_{2}^{(s)}+\beta_{3}^{(s-1)}\right)\;,\qquad\beta_{e_{w}}=2\left(\beta_{1}^{(s)}+\beta_{3}^{(s-2)}\right)\;.$
In general, gluing of boundary components always falls into one of the three above cases. The particular choice of constraints does not affect the final form of $\Omega$. Summing over all vertices $v$ of the trinion graph and using the ordering (45), we obtain
$\Omega=\frac{1}{2}\sum_{v\in V(\Gamma_{trin})}\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v\\ e<\widetilde{e}\end{subarray}}\mathrm{d}\ell_{e}\wedge\mathrm{d}\ell_{\widetilde{e}}+\sum_{e\in E(\Gamma_{trin})}\mathrm{d}\beta_{e}\wedge\mathrm{d}\ell_{e}\;.$
The prefactor $1/2$ in the first term corrects the duplication of variables due to gluing.
10. Application to $\mathcal{T}_{g}$ and $\mathcal{M}_{g}$
The $SL(2,\mathbb{R})$ character variety $\mathcal{V}_{g}^{\mathbb{R}}$ has several connected components classified by their Euler class (Euler numbers of the corresponding representations) [14]. In this section, we prove that the representations in $\mathcal{V}_{g}^{\mathbb{R}}$ constructed by plumbing of two or more lower genus Riemann surfaces with boundaries lie in the Fuchsian component of $\mathcal{V}_{g}^{\mathbb{R}}$. Projection to the Fuchsian component of the $PSL(2,\mathbb{R})$ character variety enables us to parametrize the Teichmüller space $\mathcal{T}_{g}$ using log-canonical coordinates. We explain in detail the argument for two Riemann surfaces, the generalization to multiple plumbing being straightforward. These results are already known in the existing literature; we only adapt them to our setting and give proofs for the sake of self-completeness.
Given the setting of Section 5, consider two Fuchsian representations $\widetilde{\rho}:\pi_{1}(\widetilde{{\mathcal{C}}},\widetilde{v})\rightarrow SL(2,\mathbb{R})$ and $\widehat{\rho}:\pi_{1}(\widehat{{\mathcal{C}}},\widehat{v})\rightarrow SL(2,\mathbb{R})$. Such Fuchsian representations are given by hyperbolic matrices
$\left\{\left\{M_{\widetilde{\alpha}_{j}},M_{\widetilde{\beta}_{j}}\right\}_{j=1}^{\widetilde{g}},M_{\widetilde{\gamma}}\right\},\hskip 19.91684pt\left\{\left\{M_{\widehat{\alpha}_{j}},M_{\widehat{\beta}_{j}}\right\}_{j=1}^{\widehat{g}},M_{\widehat{\gamma}}\right\}$,
still satisfying relations (18) and (19), respectively. The representations $\widetilde{\rho}$ and $\widehat{\rho}$ act on $\mathbb{RP}^{1}$ by Möbius transformations, since $\mathbb{RP}^{1}\cong S^{1}=\mathbb{R}/\mathbb{Z}$ (by stereographic projection), one can lift these actions to $\mathbb{R}$. We denote by $R_{\widetilde{\rho}}$$\mathbb{R}\rightarrow\mathbb{R}$ and by $R_{\widehat{\rho}}$$\mathbb{R}\rightarrow\mathbb{R}$ the lifted actions of $\widetilde{\rho}$ and $\widehat{\rho}$, respectively.
Taking an element $M\in SL(2,\mathbb{R})$, its action on $\mathbb{RP}^{1}\cong S^{1}$ can be lifted to an action $R_{M}:\mathbb{R}\to\mathbb{R}$. The Poincaré rotation number is then
$\mathrm{rot}\left(R_{M}\right)=\lim_{n\to\infty}\frac{R_{M}^{n}(\theta_{0})-\theta_{0}}{n}\;,$
for $\theta_{0}\in\mathbb{R}$ not a fixed point. For hyperbolic elements this limit is an integer, independent of $\theta_{0}$, and invariant under conjugation.
[…]
$e(\widetilde{\rho})=\mathrm{rot}\left(R_{M_{\widetilde{\gamma}}}\right)+\sum_{i=1}^{\widetilde{g}}\mathrm{rot}\left(\left[R_{M_{\widetilde{\alpha}_{i}}},R_{M_{\widetilde{\beta}_{i}}}\right]\right),\hskip 19.91684pte(\widehat{\rho})=\mathrm{rot}\left(R_{M_{\widehat{\gamma}}}\right)+\sum_{i=1}^{\widehat{g}}\mathrm{rot}\left(\left[R_{M_{\widehat{\alpha}_{i}}},R_{M_{\widehat{\beta}_{i}}}\right]\right).$
Lemma 10.1.
For the two Fuchsian representations $\widetilde{\rho}:\pi_{1}(\widetilde{{\mathcal{C}}},\widetilde{v})\rightarrow SL(2,\mathbb{R})$ and $\widehat{\rho}:\pi_{1}(\widehat{{\mathcal{C}}},\widehat{v})\rightarrow SL(2,\mathbb{R})$ we have:
$e(\widetilde{\rho})=2\widetilde{g}-1\;,\hskip 19.91684pte(\widehat{\rho})=2\widehat{g}-1\;.$
Proof.
[…]
$\mathrm{rot}\left(R_{M_{\widetilde{\gamma}}}\right)=\mathrm{rot}\left(R_{M_{\widehat{\gamma}}}\right)=\pm 1\;.$
We adopt the +1+1 convention, which corresponds to the orientation-preserving choice. If we close the surface $\widetilde{{\mathcal{C}}}$ by capping off the boundary, the resulting closed surface of genus $\widetilde{g}$ carries a Fuchsian representation with Euler number $2\widetilde{g}-2$. Hence,
$\sum_{i=1}^{\widetilde{g}}\mathrm{rot}\!\left([R_{M_{\widetilde{\alpha}_{i}}},R_{M_{\widetilde{\beta}_{i}}}]\right)=2\widetilde{g}-2\;.$
The same argument applies for $\widehat{{\mathcal{C}}}$. Putting everything together it follows that
$e(\widetilde{\rho})=2\widetilde{g}-1\;,\hskip 19.91684pte(\widehat{\rho})=2\widehat{g}-1\;.$
Proposition 10.2.
Consider the representation $\rho:\pi_{1}({\mathcal{C}},\widehat{v})\rightarrow SL(2,\mathbb{R})$ given by the set of matrices constructed in Proposition 5.1. We have $e(\rho)=2g-2$.
Proof.
For the representation $\rho$ constructed in Proposition 5.1 we have
$e(\rho)$ $=\sum_{i=1}^{\widetilde{g}}\mathrm{rot}\left(R_{\widehat{C}\mathfrak{b}^{-1}\widetilde{C}^{-1}}\left[R_{M_{\widetilde{\alpha}_{i}}},R_{M_{\widetilde{\beta}_{i}}}\right]R_{\widetilde{C}\mathfrak{b}\widehat{C}^{-1}}\right)+\sum_{i=1}^{\widehat{g}}\mathrm{rot}\left(\left[R_{M_{\widehat{\alpha}_{i}}},R_{M_{\widehat{\beta}_{i}}}\right]\right)$
$=\sum_{i=1}^{\widetilde{g}}\mathrm{rot}\left(\left[R_{M_{\widetilde{\alpha}_{i}}},R_{M_{\widetilde{\beta}_{i}}}\right]\right)+\sum_{i=1}^{\widehat{g}}\mathrm{rot}\left(\left[R_{M_{\widehat{\alpha}_{i}}},R_{M_{\widehat{\beta}_{i}}}\right]\right)$
$=e(\widetilde{\rho})+e(\widehat{\rho})-\left(\mathrm{rot}\left(\Lambda\right)+\mathrm{rot}\left(\Lambda^{-1}\right)\right)$
$=2\widetilde{g}-1+2\widehat{g}-1-0$
$=2g-2\;.$
Therefore, the representation constructed in Proposition 5.1 lies in the Fuchsian component of $\mathcal{V}_{g}^{\mathbb{R}}$. Projecting to $PSL(2,\mathbb{R})$, we have the following corollary by Goldman’s classification of connected components of $PSL(2,\mathbb{R})$ character varieties.
Corollary 10.3.
Let $\rho:\pi_{1}(\mathcal{C})\to SL(2,\mathbb{R})$ be the representation constructed by Proposition 5.1, and let $\bar{\rho}:\pi_{1}(\mathcal{C})\to PSL(2,\mathbb{R})$ be its projection. Then the conjugacy class of $\bar{\rho}$ defines a point of the Teichmüller space $\mathcal{T}_{g}$.
Proof.
By Proposition 10.2 the Euler number of the representation $\rho$ is
$e(\rho)=2g-2\;.$
The central subgroup $\{\pm\mathbb{I}_{2}\}\subset SL(2,\mathbb{R})$ acts trivially on $\mathbb{RP}^{1}$, therefore the Euler number of the induced $PSL(2,\mathbb{R})$ representation $\bar{\rho}$ coincides with $e(\rho)$. In particular $\bar{\rho}$ achieves the maximum possible value of the Euler number:
$|e(\bar{\rho})|=2g-2\;.$
By Goldman’s classification of connected components of $PSL(2,\mathbb{R})$ character varieties over compact Riemann surfaces [14, Corollary C], representations whose Euler number is maximal (in absolute value) are exactly the discrete, faithful representations. Therefore, $\bar{\rho}$ is discrete and faithful and its conjugacy classes determine a point of the Teichmüller space $\mathcal{T}_{g}$.
The following lemma gives the other direction ; any point in $\mathcal{T}_{g}$ can be recovered from the data of a Fuchsian $SL(2,\mathbb{R})$ representation $\rho$.
Lemma 10.4.
Consider the following subsets of $\mathcal{V}_{g}^{\mathbb{R}}$:
$\mathcal{V}_{g}^{+}=\{\,\rho^{+}:\pi_{1}(\mathcal{C})\to SL(2,\mathbb{R})\mid\;\prod_{i=1}^{g}[M_{\alpha_{i}},M_{\beta_{i}}]=+\mathbb{I}_{2}\,\}/\!\!/SL(2,\mathbb{R})\;,$
where the representations $\rho^{+}\in\mathcal{V}^{+}$ are assumed to be Fuchsian. Then the natural projection
$\Pi^{+}:\;$ $\mathcal{V}_{g}^{+}\rightarrow\mathcal{T}_{g}$
$\rho^{+}\mapsto\bar{\rho}$
is surjective.
Proof.
Every point of the Teichmüller space $\mathcal{T}_{g}$ corresponds to a Fuchsian representation
$\bar{\rho}:\pi_{1}(\mathcal{C})\to PSL(2,\mathbb{R})\;,$
which can be realized as the side-pairing group of a fundamental $4g$-gon in the hyperbolic plane $\mathbb{H}^{2}$. Each side-pairing is an isometry of $\mathbb{H}^{2}$ and each such isometry has two preimages in $SL(2,\mathbb{R})$. W. Abikoff, in [1], proved that one can choose the lifts consistently by imposing a canonical sign convention (for instance, requiring the lower-left entry of each matrix to be positive). With this convention, the relation
$\prod_{i=1}^{g}[M_{\alpha_{i}},M_{\beta_{i}}]=+\mathbb{I}_{2}$
holds in $SL(2,{\mathbb{R}})$. Thus every Fuchsian representation $\bar{\rho}:\pi_{1}({\mathcal{C}})\to PSL(2,{\mathbb{R}})$ has a canonical lift $\rho:\pi_{1}({\mathcal{C}})\to SL(2,{\mathbb{R}})$ lying in $\mathcal{V}_{g}^{+}$. Since every point of $\mathcal{T}_{g}$ arises from a Fuchsian representation, the projection
$\Pi^{+}:\mathcal{V}_{g}^{+}\longrightarrow\mathcal{T}_{g}$
is surjective. ∎
In other words, Corollary 10.3 and Lemma 10.4 prove that conjugacy classes of representations constructed by gluing of Fuchsian representations presented in Proposition 5.1 and points of the Teichmuller space $\mathcal{T}_{g}$ are in one to one correspondence.
We can deduce from the above results that expression (39) for the Goldman symplectic form $\Omega$ in Theorem 7.1 holds on $\mathcal{T}_{g}$. The form must therefore coincide with the Weil-Petersson symplectic form $\omega_{WP}$. By the results of S. Wolpert, $\omega_{WP}$ is invariant under the action of the mapping class group and hence descends to the moduli space of Riemann surfaces $\mathcal{M}_{g}$.
The above reasoning can be generalized to an arbitrary number $1\leq m\leq 3g-3$ of Fuchsian representations $\rho^{(i)}:\pi_{1}({\mathcal{C}}^{(i)},v^{(i)})\rightarrow SL(2,\mathbb{R})$ (using the notation of Section 8). So, expression (43) for the Goldman form $\Omega$ also lifts to $\mathcal{T}_{g}$ (and $\mathcal{M}_{g}$). Essentially, Theorem 8.1 assigns a different set of log-canonical coordinates to any component of the Deligne-Mumford boundary of $\mathcal{M}_{g}$. The real slice for m=3g3m=3g-3 (complete trinion decomposition) is discussed in detail in the next section.
Remark 10.5.
Let $\mathcal{M}_{\widetilde{g},1}[\ell_{\widetilde{\gamma}}$] (resp. $\mathcal{M}_{\widehat{g},1}[\ell_{\widehat{\gamma}}$]) be the moduli space of Riemann surfces of genus $\widetilde{g}$ (resp. $\widehat{g}$) with one geodesic boundary of length $\ell_{\widetilde{\gamma}}$ (resp. $\ell_{\widehat{\gamma}}$). Starting directly at the level of the moduli space $\mathcal{M}_{g}$ another idea of proof for Theorem 7.1 would be to parametrize each space $\mathcal{M}_{\widetilde{g},1}[\ell_{\widetilde{\gamma}}$] and $\mathcal{M}_{\widehat{g},1}[\ell_{\widehat{\gamma}}$] independently and then quotient $\mathcal{M}_{\widetilde{g},1}[\ell_{\widetilde{\gamma}}]\times\mathcal{M}_{\widehat{g},1}[\ell_{\widehat{\gamma}}$] by the $\mathbb{R}$ action
$\beta_{\widetilde{\gamma}}\mapsto\beta_{\widetilde{\gamma}}+t\;,\qquad\beta_{\widehat{\gamma}}\mapsto\beta_{\widehat{\gamma}}+t\;,\qquad t\in\mathbb{R}\;.$
The moment map $\mu:\mathcal{M}_{\widetilde{g},1}[\ell_{\widetilde{\gamma}}]\times\mathcal{M}_{\widehat{g},1}[\ell_{\widehat{\gamma}}$$\rightarrow\mathbb{R}$ of this action is given by the Hamiltonian function
$\mu(\ell_{\widetilde{\gamma}},\ell_{\widehat{\gamma}})=\ell_{\widetilde{\gamma}}-\ell_{\widehat{\gamma}}\;$.
The moduli space $\mathcal{M}_{g}$ is symplectomorphic to $\mu^{-1}(0)\sslash\mathbb{R}$, i.e:
$\mathcal{M}_{g}\cong\mu^{-1}(0)\sslash\mathbb{R}\;,$
and the Goldman symplectic form $\Omega$ can be written as in Theorem 7.1 on the reduced space $\mu^{-1}(0)\sslash\mathbb{R}$.
11. Relationship to Fenchel-Nielsen coordinates
We continue to restrict our attention to the situation in which the log-canonical coordinates are defined on the Fuchsian component of the $SL(2,\mathbb{R})$ character variety $\mathcal{V}_{g}^{{\mathbb{R}}}$. Our goal is to establish a precise relationship to the real Fenchel-Nielsen coordinates in terms of algebraic expressions that can be simply extended to the complex Fenchel-Nielsen coordinates on the $SL(2,\mathbb{C})$ character variety $\mathcal{V}_{g}$.
11.1. Dehn twist action on the variable $\beta_{\gamma}$
To simplify the discussion, we adopt for the rest of Chapter 3 a slightly different but equivalent convention for the toric actions on the spaces $\widetilde{{\mathcal{H}}}$ and $\widehat{{\mathcal{H}}}$. Instead of defining it on jump matrix $\widetilde{C}$ as $\widetilde{C}\mapsto\widetilde{C}b_{\widetilde{\gamma}}^{\boldsymbol{\sigma_{3}}}$ and $\widehat{C}\mapsto\widehat{C}b_{\widehat{\gamma}}^{\boldsymbol{\sigma_{3}}}$, we define a reduced toric action directly on the matrix $\mathfrak{b}$:
$\mathfrak{b}\mapsto\mathfrak{T}=b_{\widetilde{\gamma}}^{-\boldsymbol{\sigma_{3}}}\mathfrak{b}b_{\widehat{\gamma}}^{\boldsymbol{\sigma_{3}}}=\begin{bmatrix}0&-\sqrt{b_{\gamma}}\\ \frac{1}{\sqrt{b_{\gamma}}}&0\end{bmatrix}$,
where $b_{\gamma}=(b_{\widetilde{\gamma}}\,b_{\widehat{\gamma}})^{2}$ is the toric variable on $\mathcal{V}_{g}^{{\mathbb{R}}}$. Thus the jump matrix on the curve $\gamma$ is taken to be $\mathfrak{T}$. Figure 10 illustrates how a Dehn twist along $\gamma$ deforms the graph $\Gamma_{2}$. Under such a twist, the log–canonical variable $\beta_{\gamma}$ transforms as follows:
Figure 10. Action of a Dehn twist on the plumbing region.
Proposition 11.1.
A Dehn twist on the plumbing region acts on the log–canonical variable $\beta_{\gamma}$ by:
$\beta_{\gamma}\mapsto\beta_{\gamma}+\ell_{\gamma}\;.$
Proof.
To see how the Dehn twist acts on the monodromies, consider the red loop corresponding to the trivial monodromy $M=\mathbb{I}_{2}$. Before the Dehn twist this monodromy (Figure 10 left) is given by
$M=\widetilde{C}M_{\widetilde{\gamma}}^{-1}\widetilde{C}^{-1}\mathfrak{T}^{-1}\widehat{C}M_{\widehat{\gamma}}\widehat{C}^{-1}\mathfrak{T}=\mathbb{I}_{2}\;.$
After the Dehn twist (Figure 10 right) it becomes
$M_{d}=\widetilde{C}M_{\widetilde{\gamma}}^{-1}\widetilde{C}^{-1}\Lambda^{-1}\mathfrak{T}^{-1}\widehat{C}M_{\widehat{\gamma}}\widehat{C}^{\prime-1}\Lambda\mathfrak{T}=\mathbb{I}_{2}\;.$
Therefore, this twist only acts on $\mathfrak{T}$ by
$\mathfrak{T}\mapsto\Lambda\mathfrak{T}\;.$
Consequently, $b_{\gamma}$ transforms as $b_{\gamma}\mapsto b_{\gamma}\lambda_{\gamma}$, which in logarithmic form gives $\beta_{\gamma}\mapsto\beta_{\gamma}+\ell_{\gamma}$. ∎
For a closed geodesic $\gamma$, one can take the ratio of the real Fenchel-Nielsen twist parameter $\tau_{\gamma}$ over the length $\ell_{\gamma}$. Then, the quantity (with some suitable choice of origin and orientation for $\tau_{\gamma}$)
$\theta_{\gamma}=2\pi\,\frac{\tau_{\gamma}}{\ell_{\gamma}}\;,$
can be interpreted as an angle. In the present setting, the corresponding quantity is
$2\pi\,\frac{\beta_{\gamma}}{\ell_{\gamma}}\;,$
where $\beta_{\gamma}$ is the log-canonical coordinate conjugated to $\ell_{\gamma}$. To describe the relation between $\beta_{\gamma}$ and $\tau_{\gamma}$, we consider next the real length-twist coordinates arising from the Fuchsian uniformization of a trinion.
11.2. Fuchsian uniformization of a trinion
Consider again a single trinion $\mathcal{T}$ with the triangulation shown in Figure 8. Take the Fuchsian representation $\rho:\pi_{1}(\mathcal{T},p)\to SL(2,\mathbb{R})$ determined by the three hyperbolic monodromies whose explicit form is given by (48). The eigenvalues of $M_{j}$ are labeled so that $\lambda_{j}>1$.
The edge variables $z_{e_{j}}$, $j=1,2,3$, are now real and positive. They can still be expressed in terms of the eigenvalues of $M_{j}$ via (47). The next lemma describes the fixed points of the Möbius transformations corresponding to the monodromies $M_{j}$ in the upper half-plane $\mathbb{H}$.
Lemma 11.2.
For each monodromy $M_{j}$, let $\mathrm{Fix}(M_{j})=\{f_{j}^{(+)},f_{j}^{(-)}\}$ denote the pair of attracting and repelling fixed points of the associated Möbius transformation. Then
$\begin{split}{\rm Fix}(M_{1})&=\left\{\frac{(1-\lambda_{1}^{2})\lambda_{3}}{\lambda_{1}\lambda_{2}+\lambda_{3}},0\right\}=\{f_{1}^{(+)},0\}\;,\\ {\rm Fix}(M_{2})&=\left\{\frac{\lambda_{2}\lambda_{3}+\lambda_{1}}{\lambda_{2}\left(\lambda_{1}\lambda_{2}+\lambda_{3}\right)},1\right\}=\{f_{2}^{(+)},1\}\;,\\ {\rm Fix}(M_{3})&=\left\{\frac{\lambda_{3}(\lambda_{2}\lambda_{3}+\lambda_{1})}{\lambda_{2}\left(\lambda_{3}^{2}-1\right)},\infty\right\}=\{f_{3}^{(+)},\infty\}\;.\end{split}$ (54)
In addition, the fixed points satisfy
$f_{1}^{(+)}<0<f_{2}^{(+)}<1<f_{3}^{(+)}\;.$
Proof.
The proof is obtained by direct inspection of the matrices (48) and recalling the running assumption that $\lambda_{j}>1$. ∎
11.3. Gluing of two trinions and real Fenchel-Nielsen twist
Glue a trinion $\widetilde{\mathcal{T}}$ with monodromies $\widetilde{M}_{1},\widetilde{M}_{2},\widetilde{M}_{3}$ (as in (48)) to another trinion $\widehat{\mathcal{T}}$ with monodromies $\widehat{M}_{1},\widehat{M}_{2},\widehat{M}_{3}$, defined in the same manner, along the axis of $\widetilde{M}_{3}$ (Figure 11). Let $\widetilde{C}_{j}$ (resp. $\widehat{C}_{j}$) denote the matrices diagonalizing $\widetilde{M}_{j}$ (resp. $\widehat{M}_{j}$).
Lemma 11.3.
Let $\widetilde{M}_{1},\widetilde{M}_{2},\widetilde{M}_{3}$} be normalized monodromies as in (48) for a trinion $\widetilde{\mathcal{T}}$, and $\widehat{M}_{1},\widehat{M}_{2},\widehat{M}_{3}$} be similarly normalized monodromies for another trinion $\widehat{\mathcal{T}}$, with $\widetilde{\Lambda}_{3}=\widehat{\Lambda}_{3}=\Lambda$.
Then the following $SL(2,\mathbb{R})$ matrices
$M_{1}=\widetilde{M}_{1}\;,\hskip 19.91684ptM_{2}=\widetilde{M}_{2}\;,\hskip 19.91684ptM_{3}=\widetilde{C}_{3}\mathfrak{T}^{-1}\widehat{C}_{3}^{-1}\widehat{M}_{1}\widehat{C}_{3}\mathfrak{T}\widetilde{C}_{3}^{-1}\;,\hskip 19.91684ptM_{4}=\widetilde{C}_{3}\mathfrak{T}^{-1}\widehat{C}_{3}^{-1}\widehat{M}_{2}\widehat{C}_{3}\mathfrak{T}\widetilde{C}_{3}^{-1}\;,$
form a Fuchsian representation of $\pi_{1}(\mathcal{S}^{2}[\boldsymbol{\ell}],p)$, where $\mathcal{S}^{2}[\boldsymbol{\ell}]$ is a four-holed sphere with boundary lengths $\boldsymbol{\ell}=(\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widehat{\ell}_{1},\widehat{\ell}_{2})$.
Proof.
Diagonalize $\widetilde{M}_{3}$ (resp. $\widehat{M}_{3}$) by an upper-triangular matrix $\widetilde{C}_{3}$ (resp. $\widehat{C}_{3}$) so that $\widetilde{C}_{3}\widetilde{\Lambda}_{3}\widetilde{C}_{3}^{-1}=\widetilde{M}_{3}$ and $\widehat{C}_{3}\widehat{\Lambda}_{3}\widehat{C}_{3}^{-1}=\widehat{M}_{3}$. Impose $\widetilde{\Lambda}_{3}=\widehat{\Lambda}_{3}=\Lambda$. With $M_{1},M_{2},M_{3},M_{4}$ defined above, one checks
$\begin{split}M_{1}M_{2}M_{3}M_{4}&=\widetilde{M}_{1}\widetilde{M}_{2}\widetilde{C}_{3}\mathfrak{T}^{-1}\widehat{C}_{3}^{-1}\widehat{M}_{1}\widehat{M}_{2}\widehat{C}_{3}\mathfrak{T}\widetilde{C}_{3}^{-1}\\ &=\widetilde{M}_{1}\widetilde{M}_{2}\widetilde{C}_{3}\mathfrak{T}^{-1}\widehat{C}_{3}^{-1}\widehat{M}_{3}^{-1}\widehat{C}_{3}\mathfrak{T}\widetilde{C}_{3}^{-1}\\ &=\widetilde{M}_{1}\widetilde{M}_{2}\widetilde{C}_{3}\mathfrak{T}^{-1}\widehat{\Lambda}_{3}^{-1}\mathfrak{T}\widetilde{C}_{3}^{-1}\\ &=\widetilde{M}_{1}\widetilde{M}_{2}\widetilde{C}_{3}\widehat{\Lambda}_{3}\widetilde{C}_{3}^{-1}\\ &=\widetilde{M}_{1}\widetilde{M}_{2}\widetilde{C}_{3}\widetilde{\Lambda}_{3}\widetilde{C}_{3}^{-1}\\ &=\widetilde{M}_{1}\widetilde{M}_{2}\widetilde{M}_{3}=\mathbb{I}_{2}\;.\end{split}$ (55)
Hence, the set of generators $\{M_{1},M_{2},M_{3},M_{4}\}$ defines a representation
$\rho:\pi_{1}\big(\mathcal{S}^{2}[\boldsymbol{\ell}],p\big)\rightarrow SL(2,\mathbb{R})\;,$
where $\mathcal{S}^{2}[\boldsymbol{\ell}]$ is a four-holed sphere with boundary lengths $\boldsymbol{\ell}=(\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widehat{\ell}_{1},\widehat{\ell}_{2})$. Since all the matrices $M_{1},M_{2},M_{3},M_{4}$ are hyperbolic and their axis in $\mathbb{H}$ are mutually disjoint, the subgroup of $SL(2,\mathbb{R})$ they generate is discrete and acts freely on $\mathbb{H}$ [21]. Therefore, the representation of $\pi_{1}(\mathcal{S}^{2}[\boldsymbol{\ell}],p)$ defined by these matrices is Fuchsian. ∎
Remark 11.4.
Lemma 11.3 is essentially Klein’s combination theorem [18] presented in a slightly different manner.
Denote by $\mathrm{Ax}(\widetilde{M}_{3})$ the axis of $\widetilde{M}_{3}$. Let $\gamma$ be the separating contour on $\mathcal{S}^{2}[\boldsymbol{\ell}]$ given by $\gamma=\mathrm{Ax}(\widetilde{M}_{3})/\langle\widetilde{M}_{3}\rangle$. It is a classical fact (see, e.g., [21, 18]) that the real Fenchel–Nielsen twist parameter $\tau_{\gamma}$ can be expressed as the hyperbolic distance along the axis of $\widetilde{M}_{3}$ between two distinguished points (see Fig. 11):
  • the intersection point $\widetilde{p}_{FN}$ of the axis of $\widetilde{M}_{3}$ with the unique geodesic orthogonal to both the axis of $\widetilde{M}_{3}$ and the axis of $M_{2}$,
  • the intersection point $\widehat{p}_{FN}^{\prime}$ of the axis of $\widetilde{M}_{3}$ with the unique geodesic orthogonal to both the axis of $\widetilde{M}_{3}$ and the axis of $M_{4}$.
Figure 11. Axis of $M_{2}$, $M_{4}$ on $\mathbb{H}$ and Fenchel-Nielsen twist $\tau_{\gamma}$.
Proposition 11.5.
For two trinions $\widetilde{\mathcal{T}}$ and $\widehat{\mathcal{T}}$ glued along the axis of $\widetilde{M}_{3}$, the real Fenchel–Nielsen twist $\tau_{\gamma}$ and the log-toric variable $\beta_{\gamma}$ are related by
$\tau_{\gamma}=\beta_{\gamma}+\ln{\Delta(\widehat{\lambda}_{1},\widehat{\lambda}_{2};\lambda)}+\ln{\Delta(\widetilde{\lambda}_{1},\widetilde{\lambda}_{2};\lambda)}\;,$ (56)
where the rational function $\Delta(x,y;z)$ is given by:
$\Delta(x,y;z)=\left[\frac{(1+xyz)(y+xz)(x+yz)}{y^{2}(z^{2}-1)^{2}(z+xy)}\right]^{1/2}\;.$
Proof.
Consider for $\widetilde{\mathcal{T}}$ the unique geodesic on $\mathbb{H}$ orthogonal to both the axes of $\widetilde{M}_{3}$ and $\widetilde{M}_{2}$. Its intersection with the axis of $\widetilde{M}_{3}$ is (see Fig. 11)
$\widetilde{p}_{FN}=\widetilde{f}_{3}^{(+)}+i\sqrt{\left(\widetilde{f}_{3}^{(+)}-\widetilde{f}_{2}^{(+)}\right)\left(\widetilde{f}_{3}^{(+)}-\widetilde{f}_{2}^{(-)}\right)}=\widetilde{f}_{3}^{(+)}+i\Delta(\widetilde{\lambda}_{1},\widetilde{\lambda}_{2};\widetilde{\lambda}_{3})\;.$
For $\widehat{\mathcal{T}}$, before gluing, the geodesic mutually orthogonal to the axis of $\widehat{M}_{2}$ and the axis of $\widehat{M}_{3}$ intersects this same axis at:
$\widehat{p}_{FN}=\widehat{f}_{3}^{(+)}+i\sqrt{\left(\widehat{f}_{3}^{(+)}-\widehat{f}_{2}^{(+)}\right)\left(\widehat{f}_{3}^{(+)}-\widehat{f}_{2}^{(-)}\right)}=\widehat{f}_{3}^{(+)}+i\Delta(\widehat{\lambda}_{1},\widehat{\lambda}_{2};\widehat{\lambda}_{3})\;.$
Upon gluing $\widetilde{\mathcal{T}}$ and $\widehat{\mathcal{T}}$ along the axis of $\widetilde{M}_{3}$, the generator $\widehat{M}_{2}$ is conjugated:
$\widehat{M}_{2}\rightarrow M_{4}=\widetilde{C}_{3}\mathfrak{T}^{-1}\widehat{C}_{3}^{-1}\widehat{M}_{2}\widehat{C}_{3}\mathfrak{T}\widetilde{C}_{3}^{-1}\;.$
This conjugation maps the fixed points as
$\widehat{f}_{3}^{(+)}\mapsto\widetilde{f}_{3}^{(+)}\;,\hskip 19.91684pt\widehat{f}_{2}^{(+)}\mapsto f_{4}^{(+)}=\widetilde{f}_{3}^{(+)}+b_{\gamma}\left(\widehat{f}_{3}^{(+)}-\widehat{f}_{2}^{(+)}\right)^{-1}\;,$
$\widehat{f}_{2}^{(-)}\mapsto f_{4}^{(-)}=\widetilde{f}_{3}^{(+)}+b_{\gamma}\left(\widehat{f}_{3}^{(+)}-\widehat{f}_{2}^{(-)}\right)^{-1}\;.$
Hence,
$\widehat{p}_{FN}\mapsto\widehat{p}_{FN}^{\prime}=\widetilde{f}_{3}^{(+)}+\frac{ib_{\gamma}}{\sqrt{\left(\widehat{f}_{3}^{(+)}-\widehat{f}_{2}^{(+)}\right)\left(\widehat{f}_{3}^{(+)}-\widehat{f}_{2}^{(-)}\right)}}=\widetilde{f}_{3}^{(+)}+ib_{\gamma}\Delta(\widehat{\lambda}_{1},\widehat{\lambda}_{2};\widehat{\lambda}_{3})^{-1}\;.$
By definition, the Fenchel–Nielsen twist is the following hyperbolic distance along the axis of $\widetilde{M}_{3}$:
$\tau_{\gamma}=\ell(\widetilde{p}_{FN},\widehat{p}_{FN}^{\prime})=\ln{\left[b_{\gamma}\Delta(\widehat{\lambda}_{1},\widehat{\lambda}_{2};\widehat{\lambda}_{3})\Delta(\widetilde{\lambda}_{1},\widetilde{\lambda}_{2};\widetilde{\lambda}_{3})\right]}=\beta_{\gamma}+\ln{\Delta(\widehat{\lambda}_{1},\widehat{\lambda}_{2};\widehat{\lambda}_{3})}+\ln{\Delta(\widetilde{\lambda}_{1},\widetilde{\lambda}_{2};\widetilde{\lambda}_{3})}\;.$
Imposing the constraint $\widetilde{\lambda}_{3}=\lambda$ and $\widehat{\lambda}_{3}=\lambda$ we get formula (56). ∎
Remark 11.6.
There is some arbitrariness in the choice of geodesics. We could also have taken the geodesic mutually orthogonal to the axis of $\widetilde{M}_{1}$ and $\widetilde{M}_{3}$ and the geodesic mutually orthogonal to the axis of $M_{3}$ and $\widetilde{M}_{3}$. The twist magnitude defined in this way would just differ from the one defined above by half a Dehn twist.
11.4. Closing an handle : the non-separating case
In the previous subsection, we have only considered the case of gluing two different trinions. Let us now discuss the case where we close a handle, i.e., we transform a trinion $\mathcal{T}$ into a one-holed torus $\mathcal{T}^{1}[\ell_{2}]$ with boundary length $\ell_{2}=\ln{\lambda_{2}}$, by identifying the geodesic boundaries corresponding to $M_{1}$ and $M_{3}$.
For monodromies (48), if we want to keep the diagonalizing matrices $C_{1}$ and $C_{3}$ lower and upper triangular respectively, with constant coefficients on the diagonal (this ensures that the coordinates $\{\beta_{\gamma},\ell_{\gamma},\ell_{2}\}$ are log-canonical for $\Omega$ in $\mathcal{M}_{1,1}[\ell_{2}]$, by (14)), we impose the constraint $\Lambda_{1}=\Lambda_{3}^{-1}=\Lambda$. This condition is satisfied without permuting the eigenvalues. The corresponding $b_{\gamma}$-jump matrix $\mathfrak{B}$ must then be diagonal:
$\mathfrak{B}=\begin{pmatrix}\sqrt{b_{\gamma}}&0\\ 0&1/\sqrt{b_{\gamma}}\end{pmatrix}\;$.
The following lemma gives a Fuchsian representation of the one-holed torus $\mathcal{T}_{1,1}[\ell_{2}]$ under such a convention.
Lemma 11.7.
Let $\{M_{1},M_{2},M_{3}\}$ be monodromies as in (48), such that $\Lambda_{1}=\Lambda_{3}^{-1}=\Lambda$. Define
$A=M_{3}\;,\hskip 19.91684ptB=C_{3}\mathfrak{B}^{-1}C_{1}^{-1},$
where $C_{1}$ (resp. $C_{3}$) diagonalizes $M_{1}$ (resp. $M_{3}$) and $\mathfrak{B}$ is the diagonal $b$-jump matrix. Then the pair $\{A,B\}$ defines a Fuchsian representation
$\rho:\pi_{1}(\mathcal{T}^{1}[\ell_{2}],p)\rightarrow SL(2,\mathbb{R})\;,$
where $\mathcal{T}^{1}[\ell_{2}]$ is a one-holed torus with boundary length $\ell_{2}=\ln\lambda_{2}$.
Proof.
Set $A=M_{3}$ and $B=C_{3}\mathfrak{B}^{-1}C_{1}^{-1}$ as above, with the constraint $\Lambda_{1}=\Lambda_{3}^{-1}=\Lambda$. Then one computes the commutator:
$\begin{split}AB^{-1}A^{-1}B&=M_{3}C_{1}\mathfrak{B}^{-1}C_{3}^{-1}M_{3}^{-1}C_{3}\mathfrak{B}C_{1}^{-1}\\ &=M_{3}C_{1}\mathfrak{B}^{-1}\Lambda_{3}^{-1}\mathfrak{B}C_{1}^{-1}\\ &=M_{3}C_{1}\Lambda_{1}C_{1}^{-1}\\ &=M_{3}M_{1}\\ &=M_{2}^{-1}.\end{split}$ (57)
Hence, the matrices A,B{A,B} satisfy the relation of the fundamental group of a one-holed torus
$AB^{-1}A^{-1}B=M_{2}^{-1}\;,$
so they define a representation of $\pi^{1}(\mathcal{T}^{1}[\ell_{2}],p)\rightarrow SL(2,\mathbb{R})$. Again, since all the matrices $A$, $B$, $M_{2}$ are hyperbolic and their axes in $\mathbb{H}$ are disjoint, the subgroup of $SL(2,\mathbb{R})$ they generate is discrete and acts freely on $\mathbb{H}$. Hence the corresponding representation is Fuchsian. ∎
Let $\mathrm{Ax}(M_{3})$ be the axis of $M_{3}$, and $\gamma$ be the separating contour on $\mathcal{T}$ given by $\gamma=\mathrm{Ax}(M_{3})/\langle M_{3}\rangle$. We have the following relation between the real Fenchel-Nielsen twist $\tau_{\gamma}$ and the log-toric variable $\beta_{\gamma}$ when $\mathcal{T}$ is transformed into a one-holed torus $\mathcal{T}^{1}[\ell_{2}]$.
Proposition 11.8.
Let $\mathcal{T}$ be a trinion with monodromies (48). Transforming $\mathcal{T}$ into a one-holed torus $\mathcal{T}^{1}[\ell_{2}]$ by gluing along the axis of $M_{3}$, the real Fenchel-Nielsen twist $\tau_{\gamma}$ is related to the log-toric variable $\beta_{\gamma}$ by
$\tau_{\gamma}=\beta_{\gamma}+\ln{\Delta^{2}(\lambda,\lambda_{2};\lambda)}\;,$ (58)
where
$\Delta^{2}(\lambda,\lambda_{2};\lambda)=\frac{(\lambda^{2}+\lambda_{2})(\lambda_{2}\lambda^{2}+1)}{(\lambda^{2}-1)^{2}\lambda_{2}^{2}}\;.$ (59)
Proof.
The real Fenchel-Nielsen twist $\tau_{\gamma}$ is defined as the hyperbolic distance between:
  • the intersection point $p_{1}$ of the axis of $M_{3}$ with the unique geodesic mutually orthogonal to the axes of $M_{2}^{\prime}=BM_{2}B^{-1}$ and $M_{3}$, and
  • the intersection point $p_{2}$ of the axis of $M_{3}$ with the unique geodesic mutually orthogonal to the axes of $M_{2}$ and $M_{3}$.
The fixed points of the relevant hyperbolic generators are:
$\mathrm{Fix}(M_{2})$ $=\left\{f_{2}^{(+)},f_{2}^{(-)}\right\}=\left\{1/\lambda_{2},1\right\}\;,$
$\mathrm{Fix}(M_{2}^{\prime})$ $=\left\{f_{2}^{(+)\prime},f_{2}^{(-)\prime}\right\}$
$=\left\{\frac{b_{\gamma}^{2}\lambda^{2}(\lambda^{2}\lambda_{2}+1)(\lambda_{2}+1)+\lambda_{2}(1-\lambda^{2})^{2}}{b_{\gamma}^{2}\lambda_{2}(\lambda^{2}-1)(\lambda^{2}\lambda_{2}+1)},\frac{b_{\gamma}^{2}\lambda^{2}(\lambda^{2}+\lambda_{2})(\lambda_{2}+1)+\lambda_{2}(1-\lambda^{2})^{2}}{b_{\gamma}^{2}(\lambda^{2}-1)(\lambda^{2}+\lambda_{2})}\right\}\;,$
$\mathrm{Fix}(M_{3})$ $=\left\{f_{3}^{(+)},f_{3}^{(-)}\right\}=\left\{\frac{\lambda^{2}(\lambda_{2}+1)}{\lambda_{2}(\lambda^{2}-1)},\infty\right\}\;.$
Assuming $b_{\gamma}>0$, one checks that
$0<f_{2}^{(+)}<1<f_{3}^{(+)}<f_{2}^{(+)\prime}<f_{2}^{(-)\prime}\;.$
Hence, the coordinates of the intersection points are
$p_{1}=i\sqrt{\left(f_{2}^{(-)\prime}-f_{3}^{(+)}\right)\left(f_{2}^{(+)\prime}-f_{3}^{(+)}\right)}=i\sqrt{\frac{(\lambda^{2}+\lambda_{2})(\lambda_{2}\lambda^{2}+1)}{\lambda_{2}^{2}(\lambda^{2}-1)^{2}}}\;,$
$p_{2}=i\sqrt{\left(f_{2}^{(-)}-f_{3}^{(+)}\right)\left(f_{2}^{(+)}-f_{3}^{(+)}\right)}=i\sqrt{\frac{(\lambda^{2}-1)^{2}}{b_{\gamma}^{2}(\lambda^{2}+\lambda_{2})(\lambda_{2}\lambda^{2}+1)}}\;.$
Therefore, the Fenchel-Nielsen twist is
$\tau_{\gamma}=\ell(p_{2},p_{1})=\beta_{\gamma}+\ln{\frac{(\lambda^{2}+\lambda_{2})(\lambda_{2}\lambda^{2}+1)}{(\lambda^{2}-1)^{2}\lambda_{2}^{2}}}=\beta_{\gamma}+\ln{\Delta^{2}(\lambda,\lambda_{2};\lambda)}\;.$
11.5. Generating function on $\mathcal{M}_{0,4}[\mathbf{\ell}]$ and $\mathcal{M}_{1,1}[\ell_{2}]$
Consider first the moduli space of four-holed spheres with fixed boundary lengths $\boldsymbol{\ell}=(\widetilde{\ell}_{1},\widetilde{\ell}_{2},\widehat{\ell}_{1},\widehat{\ell}_{2})$. We denote this space by
$\mathcal{M}_{0,4}[\boldsymbol{\ell}]\;.$
The Weil–Petersson form $\omega_{WP}$ on $\mathcal{M}_{0,4}[\boldsymbol{\ell}]$ can be expressed in log-canonical coordinates $\{\ell_{\gamma}=\ln\lambda,\,\beta_{\gamma}=\ln b_{\gamma}\}$ as
$\omega_{WP}=\mathrm{d}\beta_{\gamma}\wedge\mathrm{d}\ell_{\gamma}=\mathrm{d}\!\left(\beta_{\gamma}\,\mathrm{d}\ell_{\gamma}\right)\;.$ (60)
In Fenchel–Nielsen coordinates $\{\ell_{\gamma},\tau_{\gamma}\}$ its expression is [20]:
$\omega_{WP}=\mathrm{d}\tau_{\gamma}\wedge\mathrm{d}\ell_{\gamma}=\mathrm{d}\!\left(\tau_{\gamma}\,\mathrm{d}\ell_{\gamma}\right)\;.$ (61)
Subtracting (60) from (61), we obtain
$\mathrm{d}\!\left((\tau_{\gamma}-\beta_{\gamma})\,\mathrm{d}\ell_{\gamma}\right)=0\;,$
which implies the existence of a function $g_{0}(\lambda_{0})$ such that
$\frac{\tau_{\gamma}-\beta_{\gamma}}{\lambda}=\frac{\mathrm{d}g_{0}}{\mathrm{d}\lambda}\;.$
Using Proposition 11.5, $g_{0}(\lambda)$ is given by the following antiderivative (integration constants are irrelevant)
$\begin{split}g_{0}(\lambda)&=\int\frac{\ln{\Delta(\widehat{\lambda}_{1},\widehat{\lambda}_{2};\lambda)}+\ln{\Delta(\widetilde{\lambda}_{1},\widetilde{\lambda}_{2};\lambda)}}{\lambda}\mathrm{d}\lambda\\ &=H(\widehat{\lambda}_{1},\widehat{\lambda}_{2};\lambda)+H(\widetilde{\lambda}_{1},\widetilde{\lambda}_{2};\lambda)-\mathrm{Li}_{2}(1-\lambda^{2})-\ln(\lambda^{2})\ln(\lambda^{2}-1)\;,\end{split}$ (62)
where we have introduced the auxiliary function
$H(x,y;z)=-\frac{1}{2}\Bigl[\mathrm{Li}_{2}(-xyz)+\mathrm{Li}_{2}\Bigl(-\frac{xz}{y}\Bigr)+\mathrm{Li}_{2}\Bigl(-\frac{yz}{x}\Bigr)-\mathrm{Li}_{2}\Bigl(-\frac{z}{xy}\Bigr)\Bigr]-\ln(y)\ln(z)\;,$ (63)
and $\mathrm{Li}_{2}(x)=-\int_{0}^{x}\frac{\ln(1-t)}{t}\,\mathrm{d}t$ denotes Euler’s dilogarithm. Note that all arguments of the dilogarithms in the expressions above are in $(0,1)$ or negative since all the lambda’s are greater than one, ensuring that the standard branch of Euler’s dilogarithm $\mathrm{Li}_{2}(x)=-\int_{0}^{x}\frac{\ln(1-t)}{t}\,\mathrm{d}t$ is well-defined and real-valued.
Now consider the moduli space $\mathcal{M}_{1,1}[\ell_{2}]$ of one-holed tori with fixed boundary length $\ell_{2}$. The Weil–Petersson form $\omega_{WP}$ in log-canonical coordinates $\{\ell_{\gamma},\,\beta_{\gamma}\}$ is
$\omega_{WP}=\mathrm{d}\beta_{\gamma}\wedge\mathrm{d}\ell_{\gamma}\;,$
while in Fenchel–Nielsen coordinates $\{\ell,\tau\}$ it takes the form [19]:
$\omega_{WP}=\mathrm{d}\tau_{\gamma}\wedge\mathrm{d}\ell_{\gamma}\;.$
By the same reasoning as above, there exists a generating function $g_{1}(\lambda)$ satisfying
$\frac{\tau_{\gamma}-\beta_{\gamma}}{\lambda}=\frac{\mathrm{d}g_{1}}{\mathrm{d}\lambda}\;.$
According to Proposition 11.8, $g_{1}(\lambda)$ is given by
$\begin{split}g_{1}(\lambda)&=\int\frac{\ln{\Delta^{2}(\lambda,\lambda_{2};\lambda)}}{\lambda}\mathrm{d}\lambda\\ &=-\mathrm{Li}_{2}(-\lambda_{2}\lambda^{2})-\mathrm{Li}_{2}(-\frac{\lambda^{2}}{\lambda_{2}})-\mathrm{Li}_{2}(1-\lambda^{2})-\ln{\lambda^{2}}\ln{(\lambda^{2}-1)}-2\ln{\lambda_{2}}\ln{\lambda}\;.\end{split}$ (64)
Remark 11.9.
The generating functions $g_{0}$ and $g_{1}$ are equal under the identification
$\lambda_{0}=\widehat{\lambda}_{1}=\widetilde{\lambda}_{1}=\lambda\;,\qquad\widehat{\lambda}_{2}=\widetilde{\lambda}_{2}=\lambda_{2}\;.$
Therefore, the two expressions (62) and (64) are consistent.
11.6. Generating function on $\mathcal{M}_{g}$
Consider the moduli space $\mathcal{M}_{g}$ obtained by a complete trinion decomposition, together with its associated trinion graph $\Gamma_{trin}$. Recall that on an open dense subset of $\mathcal{M}_{g}$ the form $\Omega$ must coincide with the Weil-Petersson form $\omega_{WP}$. One can compare the expression (51) for $\Omega$ in log-canonical coordinates with the one of $\omega_{WP}$ in Fenchel-Nielsen coordinates given by [20]:
$\begin{split}\omega_{WP}=\sum_{j=1}^{3g-3}\mathrm{d}\tau_{\gamma_{j}}\wedge\mathrm{d}\ell_{\gamma_{j}}=\frac{1}{2}\sum_{v\in V(\Gamma_{trin})}\left(\sum_{e\perp v}{\mathrm{d}}\tau_{e}\wedge{\mathrm{d}}\ell_{e}\right)\;.\end{split}$ (65)
In order to describe explicitly the symplectomorphism between the two sets of coordinates, one must distinguish three possible configurations of edges adjacent to given vertex $v^{(j)}$ of $\Gamma_{trin}$. Those three different configurations are shown in Figure 12. Let $e_{n},e_{r},e_{m}$ be the three edges incident to $v^{(j)}$, with counter-clockwise ordering $(e_{n},e_{r},e_{m})$. When the vertex $v^{(j)}$ is in configuration (I) and (II), the edges $e_{n},e_{r},e_{m}$ are called standard edges. When the vertex $v^{(j)}$ is in configuration (III), the edge $e_{m}$ is a standard edge and $e_{n}$ is called a loop edge.
Figure 12. The three possible configurations of a vertex $v^{(j)}$ of the trinion graph $\Gamma_{trin}$.
Proposition 11.10.
Fixing the trinion graph $\Gamma_{trin}$, the Goldman form $\Omega$ (51) coincides with the Weil-Petersson form $\omega_{WP}$ (65) on an open dense subset of $\mathcal{M}_{g}$.
Proof.
The proof is by inspection of the contribution $\omega_{WP}^{v^{(j)}}$ of a vertex $v^{(j)}$ (for every possible configurations (I), (II), (III), shown in Figure 12) to the Weil-Petersson symplectic $\omega_{WP}$.
Consider first a vertex $v^{(j)}$ in configuration (I), linked to three other vertices $v^{(i)},v^{(k)},v^{(l)}$. Let $e_{m}=[v^{(j)},v^{(i)}]$ be the standard edge corresponding to the gluing of the trinion $\mathcal{T}^{(j)}$ with $\mathcal{T}^{(i)}$.
According to Proposition 11.5, we have
$\tau_{e_{m}}=\beta_{e_{m}}+\ln{\Delta(\lambda_{e_{n}},\lambda_{e_{r}};\lambda_{e_{m}})}+\ln{\Delta(\lambda_{e_{q}},\lambda_{e_{p}};\lambda_{e_{m}})}\;,$ (66)
where the two edges $\{e_{n},e_{r}\}$ with ordering $(e_{n},e_{r},e_{m})$ are incident to $v^{(j)}$, and the two edges $\{e_{p},e_{q}\}$ with ordering $(e_{q},e_{p},e_{m})$ are incident to $v^{(i)}$ (see Figure 12, (I)).
Similarly, for the standard edges $e_{n}=[v^{(j)},v^{(k)}]$ and $e_{r}=[v^{(j)},v^{(l)}]$:
$\tau_{e_{n}}$ $=\beta_{e_{n}}+\ln\Delta(\lambda_{e_{r}},\lambda_{e_{m}};\lambda_{e_{n}})+\ln\Delta(\lambda_{e_{s}},\lambda_{e_{t}};\lambda_{e_{n}})\;,$ (67)
$\tau_{e_{r}}$ $=\beta_{e_{r}}+\ln\Delta(\lambda_{e_{m}},\lambda_{e_{n}};\lambda_{e_{r}})+\ln\Delta(\lambda_{e_{v}},\lambda_{e_{w}};\lambda_{e_{r}})\;.$ (68)
Only the first two terms in (66)–(68) correspond to the contribution of the vertex $v^{(j)}$. Inserting these expressions into (65) and keeping only the terms depending on $\lambda_{e_{m}},\lambda_{e_{n}},\lambda_{e_{r}}$, we get
$\begin{split}\omega_{WP}^{v^{(j)}}&={\mathrm{d}}\tau_{e_{m}}\wedge{\mathrm{d}}\ell_{e_{m}}+{\mathrm{d}}\tau_{e_{n}}\wedge{\mathrm{d}}\ell_{e_{n}}+{\mathrm{d}}\tau_{e_{r}}\wedge{\mathrm{d}}\ell_{e_{r}}\\ &={\mathrm{d}}\beta_{e_{m}}\wedge{\mathrm{d}}\ell_{e_{m}}+{\mathrm{d}}\beta_{e_{n}}\wedge{\mathrm{d}}\ell_{e_{n}}+{\mathrm{d}}\beta_{e_{r}}\wedge{\mathrm{d}}\ell_{e_{r}}+{\mathrm{d}}\ell_{e_{n}}\wedge{\mathrm{d}}\ell_{e_{r}}+\\ &+{\mathrm{d}}\ell_{e_{r}}\wedge{\mathrm{d}}\ell_{e_{m}}+{\mathrm{d}}\ell_{e_{m}}\wedge{\mathrm{d}}\ell_{e_{n}}\\ &=\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v^{(j)}\\ e<\widetilde{e}\end{subarray}}\left(\mathrm{d}\ell_{e}\wedge{\mathrm{d}}\ell_{\widetilde{e}}+\mathrm{d}\beta_{e}\wedge\mathrm{d}\ell_{e}\right)\;.\end{split}$
This calculation remains valid for a vertex $v^{(j)}$ in configuration (II), linked to only two other vertices $v^{(i)}$ and $v^{(k)}$ (Figure 12, (II)).
Let $v^{(j)}$ be linked to another vertex $v^{(i)}$ by a standard edge $e_{m}=[v^{(j)},v^{(i)}]$ and to itself by a loop edge $e_{n}=[v^{(j)},v^{(j)}]$ (Figure 12, (III)). Then using (56) and (58)
$\tau_{e_{m}}$ $=\beta_{e_{m}}+\ln\Delta(\lambda_{e_{n}},\lambda_{e_{n}};\lambda_{e_{m}})+\ln\Delta(\lambda_{e_{p}},\lambda_{e_{q}};\lambda_{e_{m}})\;,$ (69)
$\tau_{e_{n}}$ $=\beta_{e_{n}}+\ln\Delta^{2}(\lambda_{e_{n}},\lambda_{e_{m}};\lambda_{e_{n}})\;.$ (70)
Inserting (69)-(70) in (65) one can check the following
$\begin{split}\omega_{WP}^{v^{(j)}}&={\mathrm{d}}\tau_{e_{m}}\wedge{\mathrm{d}}\ell_{e_{m}}+{\mathrm{d}}\tau_{e_{n}}\wedge{\mathrm{d}}\ell_{e_{n}}\\ &={\mathrm{d}}\beta_{e_{m}}\wedge{\mathrm{d}}\ell_{e_{m}}+{\mathrm{d}}\beta_{e_{n}}\wedge{\mathrm{d}}\ell_{e_{n}}\\ &=\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v^{(j)}\\ e<\widetilde{e}\end{subarray}}\left(\mathrm{d}\ell_{e}\wedge{\mathrm{d}}\ell_{\widetilde{e}}+\mathrm{d}\beta_{e}\wedge\mathrm{d}\ell_{e}\right),\end{split}$ (71)
since the first terms in the last sum cancel each other.
Therefore, summing over all the vertices of $\Gamma_{trin}$, taking all three configurations into account, forms (65) and (51) indeed coincide.
Consider a symplectic potential (or Liouville form) $\theta_{FN}$ for $\omega_{WP}$ in Fenchel-Nielsen coordinates such that ${\mathrm{d}}\theta_{FN}=\omega_{WP}$. A possible choice (see equation (65)) is:
$\theta_{FN}=\frac{1}{2}\sum_{v\in V(\Gamma_{trin})}\sum_{e\perp v}\tau_{e}\mathrm{d}\ell_{e}\;.$
Similarly, for the same trinion graph, a symplectic potential in log-canonical coordinates (cf. equation (51)) can be written as
$\theta_{LC}=\theta_{LC}^{(1)}+\theta_{LC}^{(2)}\;,$
$\theta_{LC}^{(1)}=\frac{1}{2}\sum_{v\in V(\Gamma_{trin})}\sum_{e\perp v}\beta_{e}\mathrm{d}\ell_{e}\;,\hskip 19.91684pt\theta_{LC}^{(2)}=-\frac{1}{2}\sum_{v\in V(\Gamma_{trin})}\sum_{\begin{subarray}{c}e,\widetilde{e}\perp v\\ e<\widetilde{e}\end{subarray}}\ell_{\widetilde{e}}\mathrm{d}\ell_{e}\;.$
Hence, there is a generating function $G(\Gamma_{trin})$ on $\mathcal{M}_{g}$ depending only on lengths $\ell(e)$ such that:
$\theta_{FN}-\theta_{LC}=\sum_{v\in V(\Gamma_{trin})}\sum_{e\perp v}\frac{\partial G(\Gamma_{trin})}{\partial\ell_{e}}{\mathrm{d}}\ell_{e}\;.$ (72)
The generating function $G(\Gamma_{trin})$ can be expressed as follows.
Theorem 11.11.
Given the edges configurations shown in Figure 12, the generating function $G(\Gamma_{trin})$ can be written as:
$G(\Gamma_{trin})=\frac{1}{2}\sum_{e_{m}=[v^{(j)},v^{(i)}]}\left[g^{(j)}_{0}(\lambda_{e_{n}},\lambda_{e_{r}};\lambda_{e_{m}})+g^{(i)}_{0}(\lambda_{e_{p}},\lambda_{e_{q}};\lambda_{e_{m}})\right]+\frac{1}{2}\sum_{e_{n}=[v^{(j)},v^{(j)}]}g_{1}^{(j)}(\lambda_{e_{m}};\lambda_{e_{n}})\;,$ (73)
where standard edges contributions are given by
$g_{0}^{(j)}(\lambda_{e_{n}},\lambda_{e_{r}};\lambda_{e_{m}})=-\frac{1}{2}\Big[\Phi(\lambda_{e_{n}},\lambda_{e_{r}},\lambda_{e_{m}})+\mathrm{Li}_{2}(1-\lambda_{e_{m}}^{2})+\ln(\lambda_{e_{m}}^{2})\ln(\lambda_{e_{m}}^{2}-1)\Big]\;,$
$g_{0}^{(i)}(\lambda_{e_{p}},\lambda_{e_{q}};\lambda_{e_{m}})=-\frac{1}{2}\Big[\Phi(\lambda_{e_{p}},\lambda_{e_{q}},\lambda_{e_{m}})+\mathrm{Li}_{2}(1-\lambda_{e_{m}}^{2})+\ln(\lambda_{e_{m}}^{2})\ln(\lambda_{e_{m}}^{2}-1)\Big]\;,$
and loop edges contributions by
$g_{1}^{(j)}(\lambda_{e_{m}};\lambda_{e_{n}})=-\Phi(\lambda_{e_{n}},\lambda_{e_{m}},\lambda_{e_{n}})-\mathrm{Li}_{2}(1-\lambda_{e_{n}}^{2})-\ln(\lambda_{e_{n}}^{2})\ln(\lambda_{e_{n}}^{2}-1)-\ln{\lambda_{e_{m}}}\ln{\lambda_{e_{n}}}+\frac{\ln^{2}{\lambda_{e_{n}}}}{2}\;.$
Here the auxiliary function $\Phi$ is defined as
$\Phi(x,y,z)=\mathrm{Li}_{2}(-xyz)+\mathrm{Li}_{2}\Big(-\frac{xz}{y}\Big)+\mathrm{Li}_{2}\Big(-\frac{yz}{x}\Big)-\mathrm{Li}_{2}\Big(-\frac{z}{xy}\Big)\;.$
Proof.
We inspect the contribution of each edge to $G(\Gamma_{trin})$.
Let $v^{(j)}$ be in configuration (I) (see Figure 12). Consider the standard edge $e_{m}=[v^{(j)},v^{(i)}]$. Recall the ordering of edges $(e_{n},e_{r},e_{m})$ at $v^{(j)}$ and $(e_{q},e_{p},e_{m})$ at $v^{(i)}$. We want to write the term in (72) corresponding to the edge $e_{m}$. This term can be splitted into two. Fixing $e=e_{m}$ in the difference $\theta_{FN}-\theta_{LC}^{(1)}$, and using Proposition 11.5, we obtain
$(\tau_{e_{m}}-\beta_{e_{m}}){\mathrm{d}}\ln{\lambda_{e_{m}}}=\frac{\ln\Delta(\lambda_{e_{n}},\lambda_{e_{r}};\lambda_{e_{m}})+\ln\Delta(\lambda_{e_{q}},\lambda_{e_{p}};\lambda_{e_{m}})}{\lambda_{e_{m}}}\mathrm{d}\lambda_{e_{m}}\;.$ (74)
Integrating expression (74) we get
$\begin{split}&\int\frac{\ln\Delta(\lambda_{e_{n}},\lambda_{e_{r}};\lambda_{e_{m}})+\ln\Delta(\lambda_{e_{p}},\lambda_{e_{q}};\lambda_{e_{m}})}{\lambda_{e_{m}}}\,\mathrm{d}\lambda_{e_{m}}=-\frac{1}{2}\left[\Phi(\lambda_{e_{n}},\lambda_{e_{r}},\lambda_{e_{m}})+\Phi(\lambda_{e_{p}},\lambda_{e_{q}},\lambda_{e_{m}})\right]-\\ &-\mathrm{Li}_{2}(1-\lambda_{e_{m}}^{2})-\ln(\lambda_{e_{m}}^{2})\ln(\lambda_{e_{m}}^{2}-1)-\ln{\lambda_{e_{n}}}\ln{\lambda_{e_{m}}}-\ln{\lambda_{e_{q}}}\ln{\lambda_{e_{m}}}\;.\end{split}$ (75)
Now, fixing $e=e_{m}$ in $\theta_{LC}^{(2)}$ we have
$\sum_{\begin{subarray}{c}\widetilde{e}\perp v^{(j)}\\ e_{m}<\widetilde{e}\end{subarray}}\ell_{\widetilde{e}}\,{\mathrm{d}}\ell_{e_{m}}+\sum_{\begin{subarray}{c}\widetilde{e}\perp v^{(i)}\\ e_{m}<\widetilde{e}\end{subarray}}\ell_{\widetilde{e}}\,{\mathrm{d}}\ell_{e_{m}}=(\ln{\lambda_{e_{n}}}+\ln{\lambda_{e_{q}}}){\mathrm{d}}\ln{\lambda_{e_{m}}}\;.$ (76)
Integrating expression (76) and summing with (75) we recover
$g_{0}^{(j)}(\lambda_{e_{n}},\lambda_{e_{r}};\lambda_{e_{m}})+g_{0}^{(i)}(\lambda_{e_{p}},\lambda_{e_{q}};\lambda_{e_{m}})\;.$
This computation remains unmodified for $v^{(j)}$ in configuration (II).
Let $v^{(j)}$ be in configuration (III). Setting $\lambda_{e_{r}}=\lambda_{e_{n}}$ in the preceding computation, we obtain that the standard edge $e_{m}$ contributes to the generating function by
$g_{0}^{(j)}(\lambda_{e_{n}},\lambda_{e_{n}};\lambda_{e_{m}})+g_{0}^{(i)}(\lambda_{e_{p}},\lambda_{e_{q}};\lambda_{e_{m}})\;.$
Let us finally consider the term of (72) corresponding to the loop edge $e_{n}$ itself. Due to Proposition 11.8, we have:
$\left(\tau_{e_{n}}-\beta_{e_{n}}\right)\frac{\mathrm{d}\lambda_{e_{n}}}{\lambda_{e_{n}}}=\frac{\ln\Delta^{2}(\lambda_{e_{n}},\lambda_{e_{m}};\lambda_{e_{n}})}{\lambda_{e_{n}}}\mathrm{d}\lambda_{e_{n}}\;,$
which, after integration with respect to $\lambda_{e_{n}}$, gives
$\int\frac{\ln\Delta^{2}(\lambda_{e_{n}},\lambda_{e_{m}};\lambda_{e_{n}})}{\lambda_{e_{n}}}\,\mathrm{d}\lambda_{e_{n}}=-\Phi(\lambda_{e_{n}},\lambda_{e_{m}},\lambda_{e_{n}})-\mathrm{Li}_{2}(1-\lambda_{e_{n}}^{2})-\ln(\lambda_{e_{n}}^{2})\ln(\lambda_{e_{n}}^{2}-1)-2\ln{\lambda_{e_{m}}}\ln{\lambda_{e_{n}}}\;.$ (77)
Fixing $e=e_{n}$ in $\theta_{FG}^{(2)}$ in this situation we obtain:
$\sum_{\begin{subarray}{c}\widetilde{e}\perp v^{(j)}\\ e_{m}<\widetilde{e}\end{subarray}}\ell_{\widetilde{e}}\,{\mathrm{d}}\ell_{e_{m}}=(\ln{\lambda_{e_{n}}}+\ln{\lambda_{e_{m}}}){\mathrm{d}}\ln{\lambda_{e_{n}}}\;.$ (78)
Integrating (78) and summing with (77) we get $g_{1}(\lambda_{e_{m}};\lambda_{e_{n}})$.
Summing over all edges in $\Gamma_{\rm trin}$ we recover expression (73) for $G(\Gamma_{\rm trin})$. ∎
12. Higher $SL(N)$ groups
The generalization of our construction to an arbitrary $SL(N)$ character variety (in both complex and real cases) of $\pi_{1}({\mathcal{C}})$ for compact genus $g$ Riemann surfaces is pretty straightforward if one uses the system of higher rank Fock-Goncharov coordinates [8]. Technically, in the higher $SL(N)$ case the jumps matrices on edges $S_{e}$ become anti-diagonal; each such matrix contains N1N-1 coordinates while matrices $A$ also contain some of the coordinates (one coordinate for $SL(3)$ case, three for $SL(4)$ etc) [3]. The logarithms of eigenvalues of the monodromy matrices around the separating contours and the logarithms of the ratios of the corresponding toric variables form an additional set of N1N-1 conjugated pairs of coordinates for the Goldman symplectic form. The technical details for an arbitrary $SL(N)$ case are very similar to [3].
13. Conclusion and future research directions
In this paper, we have constructed explicitly new systems of log-canonical coordinates on $SL(2,\mathbb{C})$ character varieties of compact Riemann surfaces as combinations of shear type coordinates and length-twist type coordinates. For complete trinion decomposition characterized by a chosen trinion graph $\Gamma_{trin}$, the relationship to Fenchel-Nielsen coordinates on $\mathcal{M}_{g}$ has been established through a generating function $\Gamma_{trin}$. These new systems of log-canonical coordinates as well as generating functions may find several potential applications.
First, consider a one-holed torus $\mathcal{T}^{1}[\ell_{2}]$ as in Section 11.4, together with some choice of $SL(2,\mathbb{R})$ monodromies $\{A,B\}$ satisfying $AB^{-1}A^{-1}B=M_{2}$. Consider the Hamiltonian of the two-body Ruijsenaars system given by (cf. [17] [9], [15] equation (42)):
$H(p,q)=2\cosh{q}\sqrt{\frac{4\sinh^{2}{p}-m-2}{4\sinh^{2}{p}-4}}\;,$
according to the authors of [15], this Hamiltonian coincides with $\mathrm{tr}A$ on $\mathcal{M}_{1,1}[\ell_{2}]$ (the parameter $m$ is chosen such that $\mathrm{tr}M_{2}=m$). The action-angle coordinates in their work satisfy $\{p,q\}_{G}=1$, so they must coincide with Fenchel-Nielsen coordinates (maybe with a different convention for the twist $\tau_{\gamma}$). Thus, the generating function $g_{1}(\lambda)$ given by expression (64) should be related to the Hamilton-Jacobi action generating the canonical transformation $(\ell_{\gamma}=\ln{\lambda},\beta_{\gamma})\mapsto(q,p)$. This correspondence may be generalized to $\mathcal{M}_{g}$, for which the generating function $G(\Gamma_{trin})$ could probably be interpreted as a Hamilton-Jacobi action for some $SL(2,\mathbb{R})$ Hitchin-type integrable system.
Then we have also remarked that for $\mathcal{V}_{2}$, choosing the trinion graph to be the “theta graph” (see Figure 15, left, in Appendix), the generating function $G(\widetilde{\Gamma}_{trin})$ given by (82) closely resembles the ’Neumann-Zagier potential’ (cf equation (19) in [22]). The possible connection between the generating function $G(\Gamma_{trin})$ and classical limits of quantum invariants on trivalent graphs is unexpected and remains at this stage conjectural. The question also remains open whether the combination of dilogarithms appearing in $G(\Gamma_{trin})$ has an interpretation in terms of hyperbolic volumes.
Finally, there exist some analogues of the Fenchel-Nielsen coordinates for $SL(3,{\mathbb{C}})$ character varieties that have been studied very recently [23]. It would be interesting to relate this construction to our framework applied to such a case.
14. Appendix
In this appendix we explicitly compute the forms of Theorem 7.1 and Theorem 8.1 on $\mathcal{M}_{2}$ (with one, two and three contours) and on $\mathcal{M}_{3}$ (with one, two, three and four contours). For complete trinion decompositions some explicit expressions of the generating function of Theorem 11.11 are also given.
14.1. $\mathcal{M}_{2}$
14.1.1. One contour
For $\mathcal{M}_{2}$ the idea is to select a separating contour $\gamma$ which cuts a double-torus into two one-holed tori $\widetilde{{\mathcal{C}}}$ and $\widehat{{\mathcal{C}}}$, endowed with graphs $\widetilde{\Gamma}$ and $\widehat{\Gamma}$ as shown in Figure 13 (the internal edges and the contour $\gamma$ are not represented), whose associated shear type coordinates are $\{\widetilde{\zeta}_{e_{1}},\widetilde{\zeta}_{e_{2}},\widetilde{\zeta}_{e_{3}}\}$ and $\{\widehat{\zeta}_{e_{1}},\widehat{\zeta}_{e_{2}},\widehat{\zeta}_{e_{3}}\}$, respectively.
Figure 13. Single contour splitting for $\mathcal{M}_{2}$ and associated graphs.
Proposition 14.1.
In log-canonical coordinates $\{\widetilde{\zeta}_{e_{2}},\widetilde{\zeta}_{e_{3}},\widehat{\zeta}_{e_{2}},\widehat{\zeta}_{e_{3}},\ell_{\gamma},\beta_{\gamma}\}$ the Weil-Petersson form on $\mathcal{M}_{2}$ can be written :
$\omega_{WP}=2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{3}}+2\mathrm{d}\widehat{\zeta}_{e_{2}}\wedge\mathrm{d}\widehat{\zeta}_{e_{3}}+\mathrm{d}\ell_{\gamma}\wedge\left(\mathrm{d}\widetilde{\zeta}_{e_{2}}+\mathrm{d}\widetilde{\zeta}_{e_{3}}\right)+\mathrm{d}\ell_{\gamma}\wedge\left(\mathrm{d}\widehat{\zeta}_{e_{2}}+\mathrm{d}\widehat{\zeta}_{e_{3}}\right)+\mathrm{d}\beta_{\gamma}\wedge\mathrm{d}\ell_{\gamma}\;.$ (79)
Or, equivalently, the Poisson structure is given by the following Poisson tensor:
$\mathbb{P}_{ij}=\begin{bmatrix}0&-2&0&0&-1&0\\ 2&0&0&0&-1&0\\ 0&0&0&-2&-1&0\\ 0&0&2&0&-1&0\\ 1&1&1&1&0&-1\\ 0&0&0&0&1&0\\ \end{bmatrix}\;$.
Proof.
Using equation (40) for the vertex $\widetilde{v}$ of $\widetilde{\Gamma}$ in Figure 13, together with the twist-length type contribution $2\mathrm{d}\beta_{\widetilde{\gamma}}\wedge\mathrm{d}\ell_{\widetilde{\gamma}}$ coming from $\widetilde{q}$, without any gluing constraint, we get:
$\Omega(\widetilde{\Gamma})=2\mathrm{d}\widetilde{\zeta}_{e_{1}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{2}}+2\mathrm{d}\widetilde{\zeta}_{e_{1}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{3}}+2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{3}}+2\mathrm{d}\beta_{\widetilde{\gamma}}\wedge\mathrm{d}\ell_{\widetilde{\gamma}}\;$.
Similarly, using (41) for the vertex $\widehat{v}$ of $\widehat{\Gamma}$, together with the twist-length type contribution $2\mathrm{d}\beta_{\widehat{\gamma}}\wedge\mathrm{d}\ell_{\widehat{\gamma}}$ coming from $\widehat{q}$, we obtain:
$\Omega(\widehat{\Gamma})=2\mathrm{d}\widehat{\zeta}_{e_{1}}\wedge\mathrm{d}\widehat{\zeta}_{e_{2}}+2\mathrm{d}\widehat{\zeta}_{e_{1}}\wedge\mathrm{d}\widehat{\zeta}_{e_{3}}+2\mathrm{d}\widehat{\zeta}_{e_{2}}\wedge\mathrm{d}\widehat{\zeta}_{e_{3}}+2\mathrm{d}\beta_{\widehat{\gamma}}\wedge\mathrm{d}\ell_{\widehat{\gamma}}\;$.
Therefore we have:
$\left\{\begin{array}[]{ll}\widetilde{c}_{12}=\widetilde{c}_{13}=\widetilde{c}_{23}=2\\ \widehat{c}_{12}=\widehat{c}_{13}=\widehat{c}_{23}=2\end{array}\right$.
Hence we get ; $\widetilde{c}_{23}-\widetilde{c}_{13}+\widetilde{c}_{12}=2$, $\widehat{c}_{23}-\widehat{c}_{13}+\widehat{c}_{12}=2$, so applying Theorem 7.1 equation (79) follows. ∎
14.1.2. Two contours
Here one of the contour, $\gamma_{2}$, is non-separating. The resulting pieces $\widetilde{\mathcal{C}}$ and $\widehat{\mathcal{C}}$ by cutting along $\gamma_{1}$ and $\gamma_{2}$ are a one-holed torus and a three-holed sphere (i.e. a trinion) respectively, see Figure 14. We can associate to $\widetilde{\mathcal{C}}$, endowed with $\widetilde{\Gamma}$, shear-type coordinates $\{\widetilde{\zeta}_{e_{1}},\widetilde{\zeta}_{e_{2}},\widetilde{\zeta}_{e_{3}}\}$ whereas $\widehat{\mathcal{C}}$, with $\widehat{\Gamma}$, comes with shear-type coordinates $\{\widehat{\zeta}_{e_{1}},\widehat{\zeta}_{e_{2}},\widehat{\zeta}_{e_{3}}\}$, but all of them can be replaced by length-twist coordinates $\{\ell_{\gamma_{1}},\ell_{\gamma_{2}},\beta_{\gamma_{1}},\beta_{\gamma_{2}}\}$ due to the constraints.
Proposition 14.2.
In log-canonical coordinates $\{\widetilde{\zeta}_{e_{2}},\widetilde{\zeta}_{e_{3}},\ell_{\gamma_{1}},\ell_{\gamma_{2}},\beta_{\gamma_{1}},\beta_{\gamma_{2}}\}$ the Weil-Petersson form on $\mathcal{M}_{2}$ can be written :
$\omega_{WP}=2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{3}}+\mathrm{d}\ell_{\gamma_{1}}\wedge\left(\mathrm{d}\widetilde{\zeta}_{e_{2}}+\mathrm{d}\widetilde{\zeta}_{e_{3}}\right)+\mathrm{d}\beta_{\gamma_{1}}\wedge\mathrm{d}\ell_{\gamma_{1}}+\mathrm{d}\beta_{\gamma_{2}}\wedge\mathrm{d}\ell_{\gamma_{2}}\;.$
Figure 14. Two contour splitting for $\mathcal{M}_{2}$ and associated graphs.
Proof.
As in the previous example we still have (the contribution from $q_{1}^{+}$ is now given by $2\mathrm{d}\beta_{\widetilde{\gamma}_{1}}\wedge\mathrm{d}\ell_{\widetilde{\gamma}_{1}}$):
$\Omega(\widetilde{\Gamma})=2\mathrm{d}\widetilde{\zeta}_{e_{1}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{2}}+2\mathrm{d}\widetilde{\zeta}_{e_{1}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{3}}+2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{3}}+2\mathrm{d}\beta_{\widetilde{\gamma}_{1}}\wedge\mathrm{d}\ell_{\widetilde{\gamma}_{1}}\;,$ (80)
whereas equation (41) applied to the vertices $v_{1}^{-}$, $v_{2}^{-}$, $v_{2}^{+}$ of $\widehat{\Gamma}$ (together with twist-length type contributions coming from $q_{1}^{-}$, $q_{2}^{-}$, $q_{2}^{+}$) gives:
$\Omega(\widehat{\Gamma})=\mathrm{d}\widehat{\zeta}_{e_{3}}\wedge\mathrm{d}\widehat{\zeta}_{e_{1}}+\mathrm{d}\widehat{\zeta}_{e_{1}}\wedge\mathrm{d}\widehat{\zeta}_{e_{2}}+\mathrm{d}\widehat{\zeta}_{e_{2}}\wedge\mathrm{d}\widehat{\zeta}_{e_{3}}+2\mathrm{d}\beta_{\widehat{\gamma}_{1}}\wedge\mathrm{d}\ell_{\widehat{\gamma}_{1}}+2\mathrm{d}\beta_{\widehat{\gamma}_{2}}\wedge\mathrm{d}\ell_{\widehat{\gamma}_{2}}+2\mathrm{d}\beta_{\widehat{\gamma}_{3}}\wedge\mathrm{d}\ell_{\widehat{\gamma}_{3}}\;.$ (81)
The constraints for $\gamma_{1}$ reads:
$2(\widetilde{\zeta}_{e_{1}}+\widetilde{\zeta}_{e_{2}}+\widetilde{\zeta}_{e_{3}})$ $=\ell_{\widetilde{\gamma}_{1}}\;,$
$\widehat{\zeta}_{e_{1}}+\widehat{\zeta}_{e_{3}}$ $=\ell_{\widehat{\gamma}_{1}}\;,$
$\ell_{\widetilde{\gamma}_{1}}=\ell_{\widehat{\gamma}_{1}}$ $=\ell_{\gamma_{1}}\;,$
and the constraints for the non-separating contour $\gamma_{2}$ are given by:
$\widehat{\zeta}_{e_{1}}+\widehat{\zeta}_{e_{2}}$ $=\ell_{\widehat{\gamma}_{2}}\;,$
$\widehat{\zeta}_{e_{3}}+\widehat{\zeta}_{e_{2}}$ $=\ell_{\widehat{\gamma}_{3}}\;,$
$\ell_{\widehat{\gamma}_{2}}=\ell_{\widehat{\gamma}_{3}}$ $=\ell_{\gamma_{2}}\;.$
Hence:
$\widetilde{\zeta}_{e_{1}}$ $=\frac{1}{2}\ell_{\gamma_{1}}-\widetilde{\zeta}_{e_{2}}-\widetilde{\zeta}_{e_{3}}\;,\hskip 19.91684pt\widehat{\zeta}_{e_{2}}=\ell_{\gamma_{2}}-\frac{1}{2}\ell_{\gamma_{1}}\;,$
$\widehat{\zeta}_{e_{1}}$ $=\frac{1}{2}\ell_{\gamma_{1}}\;,\hskip 19.91684pt\widehat{\zeta}_{e_{3}}=\frac{1}{2}\ell_{\gamma_{1}}\;.$
Inserting those constraints in (80) and (81) we obtain:
[…]
14.1.3. Three contours
In this situation, we get complete decompositions into trinions. The two possible trinion graphs $\widetilde{\Gamma}_{trin}$ and $\widehat{\Gamma}_{trin}$ are shown in Figure 15. The log-canonical coordinates associated to $\widetilde{\Gamma}_{trin}$ (resp. $\widehat{\Gamma}_{trin}$) are denoted $\{\ell_{\widetilde{e}_{i}},\beta_{\widetilde{e}_{i}}\}_{i=1}^{3}$ (resp. $\{\ell_{\widehat{e}_{i}},\beta_{\widehat{e}_{i}}\}_{i=1}^{3}$).
Proposition 14.3.
The Weil-Petersson form $\omega_{WP}$ on $\mathcal{M}_{2}$ coincides with the following forms in coordinates $\{\ell_{\widetilde{e}_{i}},\beta_{\widetilde{e}_{i}}\}_{i=1}^{3}$, $\{\ell_{\widehat{e}_{i}},\beta_{\widehat{e}_{i}}\}_{i=1}^{3}$:
  • $\Omega(\widetilde{\Gamma}_{trin})=\mathrm{d}\beta_{\widetilde{e}_{1}}\wedge\mathrm{d}\ell_{\widetilde{e}_{1}}+\mathrm{d}\beta_{\widetilde{e}_{2}}\wedge\mathrm{d}\ell_{\widetilde{e}_{2}}+\mathrm{d}\beta_{\widetilde{e}_{3}}\wedge\mathrm{d}\ell_{\widetilde{e}_{3}}$,
  • $\Omega(\widehat{\Gamma}_{trin})=\mathrm{d}\beta_{\widehat{e}_{1}}\wedge\mathrm{d}\ell_{\widehat{e}_{1}}+\mathrm{d}\beta_{\widehat{e}_{2}}\wedge\mathrm{d}\ell_{\widehat{e}_{2}}+\mathrm{d}\beta_{\widehat{e}_{3}}\wedge\mathrm{d}\ell_{\widehat{e}_{3}}$.
Figure 15. The two trinion graphs $\widetilde{\Gamma}_{trin}$ and $\widehat{\Gamma}_{trin}$.
Proof.
The proof is by direct application of Theorem 9.2 to $\widetilde{\Gamma}_{trin}$ and $\widehat{\Gamma}_{trin}$.
Using formula (51) for $\widetilde{\Gamma}_{trin}$ one can get:
$\Omega(\widetilde{\Gamma}_{trin})=\frac{1}{2}\Big[{\mathrm{d}}\ell_{\widetilde{e}_{3}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{2}}+{\mathrm{d}}\ell_{\widetilde{e}_{1}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{3}}+{\mathrm{d}}\ell_{\widetilde{e}_{2}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{1}}\Big]+\frac{1}{2}\Big[{\mathrm{d}}\ell_{\widetilde{e}_{1}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{2}}+$
$+{\mathrm{d}}\ell_{\widetilde{e}_{2}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{3}}+{\mathrm{d}}\ell_{\widetilde{e}_{3}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{1}}\Big]+{\mathrm{d}}\beta_{\widetilde{e}_{1}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{1}}+{\mathrm{d}}\beta_{\widetilde{e}_{2}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{2}}+{\mathrm{d}}\beta_{\widetilde{e}_{3}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{3}}$
$={\mathrm{d}}\beta_{\widetilde{e}_{1}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{1}}+{\mathrm{d}}\beta_{\widetilde{e}_{2}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{2}}+{\mathrm{d}}\beta_{\widetilde{e}_{3}}\wedge{\mathrm{d}}\ell_{\widetilde{e}_{3}}\;.$
Similarly, for $\widehat{\Gamma}_{trin}$:
$\Omega(\widehat{\Gamma}_{trin})=\frac{1}{2}\Big[{\mathrm{d}}\ell_{\widehat{e}_{2}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{1}}+{\mathrm{d}}\ell_{\widehat{e}_{1}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{1}}+{\mathrm{d}}\ell_{\widehat{e}_{1}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{2}}\Big]+\frac{1}{2}\Big[{\mathrm{d}}\ell_{\widehat{e}_{2}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{3}}+$
${\mathrm{d}}\ell_{\widehat{e}_{3}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{3}}+{\mathrm{d}}\ell_{\widehat{e}_{3}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{2}}\Big]+{\mathrm{d}}\beta_{\widehat{e}_{1}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{1}}+{\mathrm{d}}\beta_{\widehat{e}_{2}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{2}}+{\mathrm{d}}\beta_{\widehat{e}_{3}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{3}}$
$={\mathrm{d}}\beta_{\widehat{e}_{1}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{1}}+{\mathrm{d}}\beta_{\widehat{e}_{2}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{2}}+{\mathrm{d}}\beta_{\widehat{e}_{3}}\wedge{\mathrm{d}}\ell_{\widehat{e}_{3}}\;.$
Proposition 14.4.
For the graph $\widetilde{\Gamma}_{trin}$, the generating function $G(\widetilde{\Gamma}_{trin})$ is given by
$\begin{split}&G(\widetilde{\Gamma}_{trin})=-3\left[\mathrm{Li}_{2}\left(\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{2}}\widetilde{\lambda}_{e_{3}}\right)+\mathrm{Li}_{2}\left(-\frac{\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{2}}}{\widetilde{\lambda}_{e_{3}}}\right)+\mathrm{Li}_{2}\left(-\frac{\widetilde{\lambda}_{e_{3}}\widetilde{\lambda}_{e_{1}}}{\widetilde{\lambda}_{e_{2}}}\right)+\mathrm{Li}_{2}\left(-\frac{\widetilde{\lambda}_{e_{2}}\widetilde{\lambda}_{e_{3}}}{\widetilde{\lambda}_{e_{1}}}\right)\right]-\\ &-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{1}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{1}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{1}}^{2}-1\right)}-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{2}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{2}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{2}}^{2}-1\right)}-\\ &-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{3}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{3}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{3}}^{2}-1\right)}+\frac{3}{2}\ln^{2}{(\widetilde{\lambda}_{e_{1}})}+\frac{3}{2}\ln^{2}{(\widetilde{\lambda}_{e_{2}})}+\frac{3}{2}\ln^{2}{(\widetilde{\lambda}_{e_{3}})}-\\ &-2\ln{(\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{2}}\widetilde{\lambda}_{e_{3}})}\;,\end{split}$ (82)
and for the graph $\widehat{\Gamma}_{trin}$ by
$\begin{split}&G(\widehat{\Gamma}_{trin})=\frac{3}{2}\left[\mathrm{Li}_{2}(-\widehat{\lambda}_{e_{1}}^{2}\widehat{\lambda}_{e_{2}})+\mathrm{Li}_{2}\left(-\frac{\widehat{\lambda}_{e_{1}}^{2}}{\widehat{\lambda}_{e_{2}}}\right)+\mathrm{Li}_{2}(-\widehat{\lambda}_{e_{3}}^{2}\widehat{\lambda}_{e_{2}})+\mathrm{Li}_{2}\left(-\frac{\widehat{\lambda}_{e_{3}}^{2}}{\widehat{\lambda}_{e_{2}}}\right)\right]+\frac{5}{2}\mathrm{Li}_{2}\left(-\widehat{\lambda}_{e_{2}}\right)-\\ &-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{1}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{1}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{1}}^{2}-1\right)}-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{2}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{2}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{2}}^{2}-1\right)}-\\ &-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{3}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{3}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{3}}^{2}-1\right)}+2\ln^{2}{\widehat{\lambda}_{e_{2}}}-4\ln{\widehat{\lambda}_{e_{2}}}+\frac{3}{2}\ln^{2}{(\widetilde{\lambda}_{e_{1}})}+\frac{3}{2}\ln^{2}{(\widetilde{\lambda}_{e_{3}})}\;.\end{split}$ (83)
Proof.
According to Theorem 11.11, for the graph $\widetilde{G}_{trin}$ which contains only standard edges, we have:
$\begin{split}G(\widetilde{\Gamma}_{trin})&=g_{0}^{(1)}(\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{2}};\widetilde{\lambda}_{e_{1}})+g_{0}^{(2)}(\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{3}};\widetilde{\lambda}_{e_{1}})+g_{0}^{(1)}(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{3}};\widetilde{\lambda}_{e_{2}})+g_{0}^{(2)}(\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{1}};\widetilde{\lambda}_{e_{2}})+\\ &+g_{0}^{(1)}(\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{1}};\widetilde{\lambda}_{e_{3}})+g_{0}^{(2)}(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{2}};\widetilde{\lambda}_{e_{3}})\\ &=-\frac{1}{2}\Big[\Phi(\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{1}})+\Phi(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{2}})+\Phi(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{2}})+\Phi(\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{1}};\widetilde{\lambda}_{e_{2}})+\\ &+\Phi(\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{3}})+\Phi(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{3}})\Big]-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{1}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{1}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{1}}^{2}-1\right)}-\\ &-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{2}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{2}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{2}}^{2}-1\right)}-\mathrm{Li}_{2}\left(1-\widetilde{\lambda}_{e_{3}}^{2}\right)-\ln{\left(\widetilde{\lambda}_{e_{3}}^{2}\right)}\ln{\left(\widetilde{\lambda}_{e_{3}}^{2}-1\right)}\;.\end{split}$ (84)
We recall
$\Phi(x,y,z)=\mathrm{Li}_{2}(-xyz)+\mathrm{Li}_{2}\Big(-\frac{xz}{y}\Big)+\mathrm{Li}_{2}\Big(-\frac{yz}{x}\Big)-\mathrm{Li}_{2}\Big(-\frac{z}{xy}\Big)\;.$
Using the identity
$\mathrm{Li}_{2}(z)+\mathrm{Li}_{2}\left(\frac{1}{z}\right)=-\frac{\pi^{2}}{6}-\frac{\ln^{2}{(-z)}}{2}\;,$
one can rewrite the auxilliary function $\Phi(x,y,z)$ as (the constant $\frac{\pi^{2}}{6}$ is irrelevant):
$\Phi(x,y,z)=\mathrm{Li}_{2}(-xyz)+\mathrm{Li}_{2}\Big(-\frac{xz}{y}\Big)+\mathrm{Li}_{2}\Big(-\frac{yz}{x}\Big)+\mathrm{Li}_{2}\Big(-\frac{xy}{z}\Big)+\frac{1}{2}\ln^{2}{\left(\frac{xy}{z}\right)}\;.$
Therefore, the first term in (84)
$-\frac{1}{2}\Big[\Phi(\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{1}})+\Phi(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{2}})+\Phi(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{2}})+\Phi(\widetilde{\lambda}_{e_{3}},\widetilde{\lambda}_{e_{1}};\widetilde{\lambda}_{e_{2}})+$
$+\Phi(\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{3}})+\Phi(\widetilde{\lambda}_{e_{1}},\widetilde{\lambda}_{e_{2}},\widetilde{\lambda}_{e_{3}})\Big]$
can be rewritten
$-3\left[\mathrm{Li}_{2}\left(\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{2}}\widetilde{\lambda}_{e_{3}}\right)+\mathrm{Li}_{2}\left(-\frac{\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{2}}}{\widetilde{\lambda}_{e_{3}}}\right)+\mathrm{Li}_{2}\left(-\frac{\widetilde{\lambda}_{e_{3}}\widetilde{\lambda}_{e_{1}}}{\widetilde{\lambda}_{e_{2}}}\right)+\mathrm{Li}_{2}\left(-\frac{\widetilde{\lambda}_{e_{2}}\widetilde{\lambda}_{e_{3}}}{\widetilde{\lambda}_{e_{1}}}\right)\right]+$
$+\frac{1}{2}\left[\ln^{2}{\left(\frac{\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{2}}}{\widetilde{\lambda}_{e_{3}}}\right)}+\ln^{2}{\left(\frac{\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{2}}}{\widetilde{\lambda}_{e_{2}}}\right)}+\ln^{2}{\left(\frac{\widetilde{\lambda}_{e_{2}}\widetilde{\lambda}_{e_{3}}}{\widetilde{\lambda}_{e_{1}}}\right)}+\ln^{2}{\left(\frac{\widetilde{\lambda}_{e_{1}}\widetilde{\lambda}_{e_{3}}}{\widetilde{\lambda}_{e_{2}}}\right)}\right]\;.$
Simplifying the second sum of four logarithmic terms and summing with the remaining terms in (84) we obtain equation (82).
Now for the graph $\widehat{\Gamma}_{trin}$ that contains two loop edges and one standard edge, we have, according to Theorem 11.11
$\begin{split}G(\widehat{\Gamma}_{trin})&=g_{1}^{(1)}(\widehat{\lambda}_{e_{2}};\widehat{\lambda}_{e_{1}})+g_{0}^{(1)}(\widehat{\lambda}_{e_{1}},\widehat{\lambda}_{e_{1}};\widehat{\lambda}_{e_{2}})+g_{0}^{(2)}(\widehat{\lambda}_{e_{3}},\widehat{\lambda}_{e_{3}};\widehat{\lambda}_{e_{2}})+g_{1}^{(2)}(\widehat{\lambda}_{e_{2}};\widehat{\lambda}_{e_{3}})\\ &=-\left[\Phi(\widehat{\lambda}_{e_{1}},\widehat{\lambda}_{e_{2}},\widehat{\lambda}_{e_{1}})+\frac{1}{2}\Phi(\widehat{\lambda}_{e_{1}},\widehat{\lambda}_{e_{1}},\widehat{\lambda}_{e_{2}})+\Phi(\widehat{\lambda}_{e_{3}},\widehat{\lambda}_{e_{3}},\widehat{\lambda}_{e_{2}})+\Phi(\widehat{\lambda}_{e_{3}},\widehat{\lambda}_{e_{2}},\widehat{\lambda}_{e_{3}})\right]-\\ &-\mathrm{Li}_{2}\left(1-\widehat{\lambda}_{e_{1}}^{2}\right)-\ln{\left(\widehat{\lambda}_{e_{1}}^{2}\right)}\ln{\left(\widehat{\lambda}_{e_{1}}^{2}-1\right)}-\mathrm{Li}_{2}\left(1-\widehat{\lambda}_{e_{2}}^{2}\right)-\ln{\left(\widehat{\lambda}_{e_{2}}^{2}\right)}\ln{\left(\widehat{\lambda}_{e_{2}}^{2}-1\right)}-\\ &-\mathrm{Li}_{2}\left(1-\widehat{\lambda}_{e_{3}}^{2}\right)-\ln{\left(\widehat{\lambda}_{e_{3}}^{2}\right)}\ln{\left(\widehat{\lambda}_{e_{3}}^{2}-1\right)}-\ln{\widehat{\lambda}_{e_{1}}}\ln{\widehat{\lambda}_{e_{2}}}-\ln{\widehat{\lambda}_{e_{3}}}\ln{\widehat{\lambda}_{e_{2}}}+\frac{1}{2}\ln^{2}{\widehat{\lambda}_{e_{1}}}+\frac{1}{2}\ln^{2}{\widehat{\lambda}_{e_{3}}}\;.\end{split}$ (85)
Using again (14.1.3), one can rewrite
$\left[\Phi(\widehat{\lambda}_{e_{1}},\widehat{\lambda}_{e_{2}},\widehat{\lambda}_{e_{1}})+\frac{1}{2}\Phi(\widehat{\lambda}_{e_{1}},\widehat{\lambda}_{e_{1}},\widehat{\lambda}_{e_{2}})+\Phi(\widehat{\lambda}_{e_{3}},\widehat{\lambda}_{e_{3}},\widehat{\lambda}_{e_{2}})+\Phi(\widehat{\lambda}_{e_{3}},\widehat{\lambda}_{e_{2}},\widehat{\lambda}_{e_{3}})\right]$
as
$\begin{split}&\frac{3}{2}\left[\mathrm{Li}_{2}(-\widehat{\lambda}_{e_{1}}^{2}\widehat{\lambda}_{e_{2}})+\mathrm{Li}_{2}\left(-\frac{\widehat{\lambda}_{e_{1}}^{2}}{\widehat{\lambda}_{e_{2}}}\right)+\mathrm{Li}_{2}(-\widehat{\lambda}_{e_{3}}^{2}\widehat{\lambda}_{e_{2}})+\mathrm{Li}_{2}\left(-\frac{\widehat{\lambda}_{e_{3}}^{2}}{\widehat{\lambda}_{e_{2}}}\right)\right]+\frac{5}{2}\mathrm{Li}_{2}\left(-\widehat{\lambda}_{e_{2}}\right)+2\ln^{2}{\widehat{\lambda}_{e_{2}}}+\\ &+\ln^{2}{\widehat{\lambda}_{e_{1}}}+\ln^{2}{\widehat{\lambda}_{e_{3}}}-4\ln{\widehat{\lambda}_{e_{2}}}\;.\end{split}$ (86)
Summing this last equation with the other terms of (85) we recover (83). ∎
14.2. $\mathcal{M}_{3}$
14.2.1. One contour
In this case $\widetilde{\mathcal{C}}$ is a one-holed double torus and $\widehat{\mathcal{C}}$ is a one-holed torus, the associated graphs $\widetilde{\Gamma}$ and $\widehat{\Gamma}$ are shown in Figure 16. We associate to $\widetilde{\mathcal{C}}$, endowed with $\widetilde{\Gamma}$, shear-type coordinates $\{\widetilde{\zeta}_{e_{1}},\widetilde{\zeta}_{e_{2}},\widetilde{\zeta}_{e_{3}},\widetilde{\zeta}_{e_{4}},\widetilde{\zeta}_{e_{5}},\widetilde{\zeta}_{e_{6}},\widetilde{\zeta}_{e_{7}},\widetilde{\zeta}_{e_{8}},\widetilde{\zeta}_{e_{9}}\}$ and to $\widehat{\mathcal{C}}$, with $\widehat{\Gamma}$, shear-type coordinates $\{\widehat{\zeta}_{e_{1}},\widehat{\zeta}_{e_{2}},\widehat{\zeta}_{e_{3}}\}$.
Figure 16. Single contour splitting for $\mathcal{M}_{3}$ and associated graphs.
Proposition 14.5.
In log-canonical coordinates $\{\widetilde{\zeta}_{e_{2}},\widetilde{\zeta}_{e_{3}},\widetilde{\zeta}_{e_{4}},\widetilde{\zeta}_{e_{5}},\widetilde{\zeta}_{e_{6}},\widetilde{\zeta}_{e_{7}},\widetilde{\zeta}_{e_{8}},\widetilde{\zeta}_{e_{9}},\widehat{\zeta}_{e_{2}},\widehat{\zeta}_{e_{3}},\ell_{\gamma},\beta_{\gamma}\}$ the Weil-Petersson form on $\mathcal{M}_{2}$ can be written :
$\begin{split}&\omega_{WP}=2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{3}}+\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{4}}+2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{5}}+2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{6}}++2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{8}}+2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{9}}+\\ &+2\mathrm{d}\widetilde{\zeta}_{e_{2}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{4}}-3\mathrm{d}\widetilde{\zeta}_{e_{3}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{5}}++2\mathrm{d}\widetilde{\zeta}_{e_{3}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{6}}-2\mathrm{d}\widetilde{\zeta}_{e_{3}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{8}}+2\mathrm{d}\widetilde{\zeta}_{e_{3}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{9}}-2\mathrm{d}\widetilde{\zeta}_{e_{4}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{5}}\\ &-2\mathrm{d}\widetilde{\zeta}_{e_{4}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{7}}-2\mathrm{d}\widetilde{\zeta}_{e_{4}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{8}}+2\mathrm{d}\widetilde{\zeta}_{e_{4}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{9}}+2\mathrm{d}\widetilde{\zeta}_{e_{5}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{9}}-2\mathrm{d}\widetilde{\zeta}_{e_{6}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{8}}-2\mathrm{d}\widetilde{\zeta}_{e_{7}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{8}}+\\ &+2\mathrm{d}\widetilde{\zeta}_{e_{7}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{9}}+4\mathrm{d}\widetilde{\zeta}_{e_{8}}\wedge\mathrm{d}\widetilde{\zeta}_{e_{9}}+2\mathrm{d}\widehat{\zeta}_{e_{2}}\wedge\mathrm{d}\widehat{\zeta}_{e_{3}}+\mathrm{d}\ell_{\gamma}\wedge\left(\mathrm{d}\widehat{\zeta}_{e_{2}}+\mathrm{d}\widehat{\zeta}_{e_{3}}\right)+\mathrm{d}\ell_{\gamma}\wedge\Big(\mathrm{d}\widetilde{\zeta}_{e_{2}}+2\mathrm{d}\widetilde{\zeta}_{e_{3}}\\ &+2\mathrm{d}\widetilde{\zeta}_{e_{4}}+\mathrm{d}\widetilde{\zeta}_{e_{5}}+\mathrm{d}\widetilde{\zeta}_{e_{6}}+2\mathrm{d}\widetilde{\zeta}_{e_{7}}+2\mathrm{d}\widetilde{\zeta}_{e_{8}}+2\mathrm{d}\widetilde{\zeta}_{e_{9}}\Big)+\mathrm{d}\beta_{\gamma}\wedge\mathrm{d}\ell_{\gamma}\;.\end{split}$ (87)
Proof.
The proof is by direct application of expression (40) and Theorem 7.1. For sake of brevity we only write the non-zero coefficients $\widetilde{c}_{ij}$ for $\Omega(\Gamma_{1})$:
[…]
Whereas for $\Omega(\Gamma_{2})$ the same non-zero coefficients are given by $\widehat{c}_{12}=\widehat{c}_{13}=\widehat{c}_{23}=2$. We let to the reader the straightforward but long computation of each $\widetilde{c}_{ij}-\widetilde{c}_{1j}+\widetilde{c}_{1i}$ in order to check (87). ∎
14.2.2. Two contours
Choose two separating contours $\gamma_{1}$ and $\gamma_{2}$. As depicted in Figure 17, bottom, we obtain by cutting along these contours a one-holed torus ${\mathcal{C}}^{(1)}$, a two-holed torus ${\mathcal{C}}^{(2)}$ and another one-holed torus ${\mathcal{C}}^{(3)}$, with respective graphs and shear-type variables: $\Gamma^{(1)}$ and $\{\zeta_{e_{1}}^{(1)},\zeta_{e_{2}}^{(1)},\zeta_{e_{3}}^{(1)}\}$ , $\Gamma^{(2)}$ and $\{\zeta^{(2)}_{e_{1}},\zeta^{(2)}_{e_{2}},\zeta^{(2)}_{e_{3}},\zeta^{(2)}_{e_{4}},\zeta^{(2)}_{e_{5}},\zeta^{(2)}_{e_{6}}\}$, $\Gamma^{(3)}$ and $\{\zeta^{(3)}_{e_{1}},\zeta^{(3)}_{e_{2}},\zeta^{(3)}_{e_{3}}\}$.
Figure 17. Two contour splitting for $\mathcal{M}_{3,0}$ and associated graphs.
Proposition 14.6.
The Weil-Petersson symplectic form on $\mathcal{M}_{3}$ in log-canonical coordinates
$\{\zeta^{(1)}_{e_{2}},\zeta^{(1)}_{e_{3}},\zeta^{(2)}_{e_{2}},\zeta^{(2)}_{e_{3}},\zeta^{(2)}_{e_{4}},\zeta^{(2)}_{e_{6}},\zeta^{(3)}_{e_{2}},\zeta^{(3)}_{e_{3}},\ell_{\gamma_{1}},\beta_{\gamma_{1}},\ell_{\gamma_{2}},\beta_{\gamma_{2}}\}$ is given by:
$\begin{split}&\omega_{WP}=2\mathrm{d}\zeta^{(1)}_{e_{2}}\wedge\mathrm{d}\zeta^{(1)}_{e_{3}}-2\mathrm{d}\zeta^{(2)}_{e_{2}}\wedge\mathrm{d}\zeta^{(2)}_{e_{4}}+2\mathrm{d}\zeta^{(2)}_{e_{2}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}+2\mathrm{d}\zeta^{(2)}_{e_{3}}\wedge\mathrm{d}\zeta^{(2)}_{e_{4}}+2\mathrm{d}\zeta^{(2)}_{e_{3}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}+\\ &+2\mathrm{d}\zeta^{(2)}_{e_{4}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}+2\mathrm{d}\zeta^{(3)}_{e_{2}}\wedge\mathrm{d}\zeta^{(3)}_{e_{3}}+\mathrm{d}\ell_{\gamma_{1}}\wedge\left(\mathrm{d}\zeta^{(1)}_{e_{2}}+\mathrm{d}\zeta^{(1)}_{e_{3}}\right)+\mathrm{d}\ell_{\gamma_{1}}\wedge\left(\mathrm{d}\zeta^{(2)}_{e_{2}}+\mathrm{d}\zeta^{(2)}_{e_{3}}-\mathrm{d}\zeta^{(2)}_{e_{6}}\right)+\\ &+\mathrm{d}\ell_{\gamma_{2}}\wedge\left(\mathrm{d}\zeta^{(2)}_{e_{3}}-\mathrm{d}\zeta^{(2)}_{e_{2}}-\mathrm{d}\zeta^{(2)}_{e_{4}}+\mathrm{d}\zeta^{(2)}_{e_{6}}\right)+\mathrm{d}\ell_{\gamma_{2}}\wedge\left(\mathrm{d}\zeta^{(3)}_{e_{2}}+\mathrm{d}\zeta^{(3)}_{e_{3}}\right)+\mathrm{d}\beta_{\gamma_{1}}\wedge\mathrm{d}\ell_{\gamma_{1}}+\mathrm{d}\beta_{\gamma_{2}}\wedge\mathrm{d}\ell_{\gamma_{2}}\end{split}$ (88)
Proof.
In this situation, by equation (40) (plus the twist-length contributions) we have:
$\begin{split}\Omega(\Gamma^{(1)})&=2\,\mathrm{d}\zeta^{(1)}_{e_{1}}\wedge\mathrm{d}\zeta^{(1)}_{e_{2}}+2\,\mathrm{d}\zeta^{(1)}_{e_{1}}\wedge\mathrm{d}\zeta^{(1)}_{e_{3}}+2\,\mathrm{d}\zeta^{(1)}_{e_{2}}\wedge\mathrm{d}\zeta^{(1)}_{e_{3}}+2\mathrm{d}\beta_{\gamma_{1}^{(1)}}\wedge\mathrm{d}\ell_{\gamma_{1}^{(1)}}\;,\\[6.45831pt] \Omega(\Gamma^{(2)})&=\mathrm{d}\zeta^{(2)}_{e_{1}}\wedge\mathrm{d}\zeta^{(2)}_{e_{2}}-\mathrm{d}\zeta^{(2)}_{e_{1}}\wedge\mathrm{d}\zeta^{(2)}_{e_{4}}+\mathrm{d}\zeta^{(2)}_{e_{1}}\wedge\mathrm{d}\zeta^{(2)}_{e_{5}}-\mathrm{d}\zeta^{(2)}_{e_{1}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}\\ &\quad{}+\mathrm{d}\zeta^{(2)}_{e_{2}}\wedge\mathrm{d}\zeta^{(2)}_{e_{3}}-\mathrm{d}\zeta^{(2)}_{e_{2}}\wedge\mathrm{d}\zeta^{(2)}_{e_{4}}-\mathrm{d}\zeta^{(2)}_{e_{2}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}+\mathrm{d}\zeta^{(2)}_{e_{3}}\wedge\mathrm{d}\zeta^{(2)}_{e_{4}}\\ &\quad{}-\mathrm{d}\zeta^{(2)}_{e_{3}}\wedge\mathrm{d}\zeta^{(2)}_{e_{5}}+\mathrm{d}\zeta^{(2)}_{e_{3}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}+\mathrm{d}\zeta^{(2)}_{e_{4}}\wedge\mathrm{d}\zeta^{(2)}_{e_{5}}+\mathrm{d}\zeta^{(2)}_{e_{4}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}\\ &\quad{}+\mathrm{d}\zeta^{(2)}_{e_{5}}\wedge\mathrm{d}\zeta^{(2)}_{e_{6}}+2\mathrm{d}\beta_{\gamma_{1}^{(2)}}\wedge\mathrm{d}\ell_{\gamma_{1}^{(2)}}+2\mathrm{d}\beta_{\gamma_{2}^{(2)}}\wedge\mathrm{d}\ell_{\gamma_{2}^{(2)}}\;,\\[6.45831pt] \Omega(\Gamma^{(3)})&=2\,\mathrm{d}\zeta^{(3)}_{e_{1}}\wedge\mathrm{d}\zeta^{(3)}_{e_{2}}+2\,\mathrm{d}\zeta^{(3)}_{e_{1}}\wedge\mathrm{d}\zeta^{(3)}_{e_{3}}+2\,\mathrm{d}\zeta^{(3)}_{e_{2}}\wedge\mathrm{d}\zeta^{(3)}_{e_{3}}+2\mathrm{d}\beta_{\gamma_{1}^{(3)}}\wedge\mathrm{d}\ell_{\gamma_{1}^{(3)}}\;.\end{split}$ (89)
The constraints for the first contour $\gamma_{1}$ read:
$2(\zeta^{(1)}_{e_{1}}+\zeta^{(1)}_{e_{2}}+\zeta^{(1)}_{e_{3}})$ $=\ell_{\gamma_{1}^{(1)}}\;,$
$2\zeta^{(2)}_{e_{1}}+2\zeta^{(2)}_{e_{2}}+\zeta^{(2)}_{e_{3}}+\zeta^{(2)}_{e_{4}}+\zeta^{(2)}_{e_{5}}+\zeta^{(2)}_{e_{6}}$ $=\ell_{\gamma_{1}^{(2)}}\;,$
$\ell_{\gamma_{1}^{(1)}}=\ell_{\gamma_{1}^{(2)}}$ $=\ell_{\gamma_{1}}\;.$
Whereas for the second contour $\gamma_{2}$ we have:
$\zeta^{(2)}_{e_{3}}+\zeta^{(2)}_{e_{4}}+\zeta^{(2)}_{e_{5}}+\zeta^{(2)}_{e_{6}}$ $=\ell_{\gamma_{2}^{(2)}}\;,$
$2(\zeta^{(3)}_{e_{1}}+\zeta^{(3)}_{e_{2}}+\zeta^{(3)}_{e_{3}})$ $=\ell_{\gamma_{1}^{(3)}}\;,$
$\ell_{\gamma_{2}^{(2)}}=\ell_{\gamma_{1}^{(3)}}$ $=\ell_{\gamma_{2}}\;.$
Therefore one can eliminate the variables $\zeta^{(1)}_{e_{1}}$, $\zeta^{(2)}_{e_{1}}$, $\zeta^{(2)}_{e_{5}}$ and $\zeta^{(3)}_{e_{1}}$ according to:
$\zeta^{(1)}_{e_{1}}$ $=\frac{1}{2}\ell_{\gamma_{1}}-\zeta^{(1)}_{e_{2}}-\zeta^{(1)}_{e_{3}}\;,\hskip 19.91684pt\zeta^{(3)}_{e_{1}}=\frac{1}{2}\ell_{\gamma_{2}}-\zeta^{(3)}_{e_{2}}-\zeta^{(3)}_{e_{3}}\;$,
$\zeta^{(2)}_{e_{1}}$ $=\frac{1}{2}(\ell_{\gamma_{1}}-\ell_{\gamma_{2}})-\zeta^{(2)}_{e_{2}}\;,\hskip 19.91684pt\zeta^{(2)}_{e_{5}}=\ell_{\gamma_{2}}-\zeta^{(2)}_{e_{3}}-\zeta^{(2)}_{e_{4}}-\zeta^{(2)}_{e_{6}}\;$.
Inserting these expressions into (89), summing $\Omega(\Gamma^{(1)})$ with $\Omega(\Gamma^{(2)})$ and $\Omega(\Gamma^{(3)})$ while defining $\beta_{\gamma_{1}}=2(\beta_{\gamma_{1}^{(1)}}+\beta_{\gamma_{1}^{(2)}})$ and $\beta_{\gamma_{2}}=2(\beta_{\gamma_{2}^{(2)}}+\beta_{\gamma_{1}^{(3)}})$ one can obtain (88). ∎
14.2.3. Three contours
Let us think about the 3-torus as a sphere with three handles as shown on Figure 18. Cutting these handles along three separating contours $\gamma_{1},\gamma_{2},\gamma_{3}$ we get four pieces: ${\mathcal{C}}^{(3)}$ a trinion and three one-holed tori ${\mathcal{C}}^{(1)}$, ${\mathcal{C}}^{(2)}$ and ${\mathcal{C}}^{(4)}$. Denote the associated graphs by $\Gamma^{(1)},\Gamma^{(2)},\Gamma^{(3)},\Gamma^{(4)}$.
Figure 18. Three contour splitting for $\mathcal{M}_{3,0}$ and associated graphs.
Proposition 14.7.
The Weil-Petersson form on $\mathcal{M}_{3}$ in log-canonical coordinates
$\{\zeta^{(1)}_{e_{2}},\zeta^{(1)}_{e_{3}},\zeta^{(2)}_{e_{2}},\zeta^{(2)}_{e_{3}},\zeta^{(4)}_{e_{2}},\zeta^{(4)}_{e_{3}},\ell_{\gamma_{1}},\ell_{\gamma_{2}},\ell_{\gamma_{3}},\beta_{\gamma_{1}},\beta_{\gamma_{2}},\beta_{\gamma_{3}}\}$ is given by:
$\begin{split}&\omega_{WP}=2\mathrm{d}\zeta^{(1)}_{e_{2}}\wedge\mathrm{d}\zeta^{(1)}_{e_{3}}+2\mathrm{d}\zeta^{(2)}_{e_{2}}\wedge\mathrm{d}\zeta^{(2)}_{e_{3}}+2\mathrm{d}\zeta^{(4)}_{e_{2}}\wedge\mathrm{d}\zeta^{(4)}_{e_{3}}+\mathrm{d}\ell_{\gamma_{1}}\wedge\left(\mathrm{d}\zeta^{(1)}_{e_{2}}+\mathrm{d}\zeta^{(1)}_{e_{3}}\right)+\\ &+\mathrm{d}\ell_{\gamma_{2}}\wedge\left(\mathrm{d}\zeta^{(2)}_{e_{2}}+\mathrm{d}\zeta^{(2)}_{e_{3}}\right)+\mathrm{d}\ell_{\gamma_{3}}\wedge\left(\mathrm{d}\zeta^{(4)}_{e_{2}}+\mathrm{d}\zeta^{(4)}_{e_{3}}\right)+\mathrm{d}\ell_{\gamma_{2}}\wedge\mathrm{d}\ell_{\gamma_{1}}+\mathrm{d}\ell_{\gamma_{3}}\wedge\mathrm{d}\ell_{\gamma_{2}}+\mathrm{d}\ell_{\gamma_{1}}\wedge\mathrm{d}\ell_{\gamma_{3}}+\\ &+\mathrm{d}\beta_{\gamma_{1}}\wedge\mathrm{d}\ell_{\gamma_{1}}+\mathrm{d}\beta_{\gamma_{2}}\wedge\mathrm{d}\ell_{\gamma_{2}}+\mathrm{d}\beta_{\gamma_{3}}\wedge\mathrm{d}\ell_{\gamma_{3}}\;.\end{split}$ (90)
Proof.
Applying equation (40), and taking into account the twist-length contributions, let us write:
$\begin{split}\Omega(\Gamma^{(1)})&=2\mathrm{d}\zeta^{(1)}_{e_{1}}\wedge\mathrm{d}\zeta^{(1)}_{e_{2}}+2\mathrm{d}\zeta^{(1)}_{e_{1}}\wedge\mathrm{d}\zeta^{(1)}_{e_{3}}+2\mathrm{d}\zeta^{(1)}_{e_{2}}\wedge\mathrm{d}\zeta^{(1)}_{e_{3}}+2\mathrm{d}\beta_{\gamma_{1}^{(1)}}\wedge\mathrm{d}\ell_{\gamma_{1}^{(1)}}\;,\\[6.45831pt] \Omega(\Gamma^{(2)})&=2\mathrm{d}\zeta^{(2)}_{e_{1}}\wedge\mathrm{d}\zeta^{(2)}_{e_{2}}+2\mathrm{d}\zeta^{(2)}_{e_{1}}\wedge\mathrm{d}\zeta^{(2)}_{e_{3}}+2\mathrm{d}\zeta^{(2)}_{e_{2}}\wedge\mathrm{d}\zeta^{(2)}_{e_{3}}+2\mathrm{d}\beta_{\gamma_{1}^{(2)}}\wedge\mathrm{d}\ell_{\gamma_{1}^{(2)}}\;,\\[6.45831pt] \Omega(\Gamma^{(3)})&=\mathrm{d}\zeta^{(3)}_{e_{2}}\wedge\mathrm{d}\zeta^{(3)}_{e_{1}}+\mathrm{d}\zeta^{(3)}_{e_{3}}\wedge\mathrm{d}\zeta^{(3)}_{e_{2}}+\mathrm{d}\zeta^{(3)}_{e_{1}}\wedge\mathrm{d}\zeta^{(3)}_{e_{3}}+2\mathrm{d}\beta_{\gamma_{1}^{(3)}}\wedge\mathrm{d}\ell_{\gamma_{1}^{(3)}}+2\mathrm{d}\beta_{\gamma_{2}^{(3)}}\wedge\mathrm{d}\ell_{\gamma_{2}^{(3)}}+\\ &+2\mathrm{d}\beta_{\gamma_{3}^{(3)}}\wedge\mathrm{d}\ell_{\gamma_{3}^{(3)}}\;,\\[6.45831pt] \Omega(\Gamma^{(4)})&=2\mathrm{d}\zeta^{(4)}_{e_{1}}\wedge\mathrm{d}\zeta^{(4)}_{e_{2}}+2\mathrm{d}\zeta^{(4)}_{e_{1}}\wedge\mathrm{d}\zeta^{(4)}_{e_{3}}+2\mathrm{d}\zeta^{(4)}_{e_{2}}\wedge\mathrm{d}\zeta^{(4)}_{e_{3}}+2\mathrm{d}\beta_{\gamma_{1}^{(4)}}\wedge\mathrm{d}\ell_{\gamma_{1}^{(4)}}\;.\end{split}$ (91)
The constraints for $\gamma_{1}$ reads:
$2(\zeta^{(1)}_{e_{1}}+\zeta^{(1)}_{e_{2}}+\zeta^{(1)}_{e_{3}})$ $=\ell_{\gamma_{1}^{(1)}}\;,$
$\zeta^{(3)}_{e_{1}}+\zeta^{(3)}_{e_{3}}$ $=\ell_{\gamma_{1}^{(3)}}\;,$
$\ell_{\gamma_{1}^{(1)}}=\ell_{\gamma_{1}^{(3)}}$ $=\ell_{\gamma_{1}}\;.$
Those for $\gamma_{2}$ are given by:
$2(\zeta^{(2)}_{e_{1}}+\zeta^{(2)}_{e_{2}}+\zeta^{(2)}_{e_{3}})$ $=\ell_{\gamma_{1}^{(2)}}\;,$
$\zeta^{(3)}_{e_{1}}+\zeta^{(3)}_{e_{2}}$ $=\ell_{\gamma_{2}^{(3)}}\;,$
$\ell_{\gamma_{1}^{(2)}}=\ell_{\gamma_{2}^{(3)}}$ $=\ell_{\gamma_{2}}\;.$
And for $\gamma_{3}$ we have:
$2(\zeta^{(4)}_{e_{1}}+\zeta^{(4)}_{e_{2}}+\zeta^{(4)}_{e_{3}})$ $=\ell_{\gamma_{1}^{(4)}}\;,$
$\zeta^{(3)}_{e_{2}}+\zeta^{(3)}_{e_{3}}$ $=\ell_{\gamma_{3}^{(3)}}\;,$
$\ell_{\gamma_{1}^{(4)}}=\ell_{\gamma_{3}^{(3)}}$ $=\ell_{\gamma_{3}}\;.$
From these constraints we get:
$\zeta^{(1)}_{e_{1}}$ $=\frac{1}{2}\ell_{\gamma_{1}}-\zeta^{(1)}_{e_{2}}-\zeta^{(1)}_{e_{3}}\;,\hskip 19.91684pt\zeta^{(2)}_{e_{1}}=\frac{1}{2}\ell_{\gamma_{2}}-\zeta^{(2)}_{e_{2}}-\zeta^{(2)}_{e_{3}}\;,\hskip 19.91684pt\zeta^{(4)}_{e_{1}}=\frac{1}{2}\ell_{\gamma_{3}}-\zeta^{(4)}_{e_{2}}-\zeta^{(4)}_{e_{3}}\;$,
$\zeta^{(3)}_{e_{1}}$ $=\frac{1}{2}\left(\ell_{\gamma_{1}}+\ell_{\gamma_{2}}-\ell_{\gamma_{3}}\right)\;,\hskip 19.91684pt\zeta^{(3)}_{e_{2}}=\frac{1}{2}\left(\ell_{\gamma_{2}}-\ell_{\gamma_{1}}+\ell_{\gamma_{3}}\right)\;,\hskip 19.91684pt\zeta^{(3)}_{e_{3}}=\frac{1}{2}\left(\ell_{\gamma_{1}}-\ell_{\gamma_{2}}+\ell_{\gamma_{3}}\right)\;.$
Inserting these expressions into (91) while setting $\beta_{\gamma_{1}}=2(\beta_{\gamma_{1}^{(1)}}+\beta_{\gamma_{1}^{(3)}})$, $\beta_{\gamma_{2}}=2(\beta_{\gamma_{1}^{(2)}}+\beta_{\gamma_{2}^{(3)}})$ and $\beta_{\gamma_{3}}=2(\beta_{\gamma_{1}^{(4)}}+\beta_{\gamma_{3}^{(3)}})$ we get (90). ∎
14.2.4. Four contours
With four contours we recover complete trinion decompositions. As an example, take the trinion graph to be the “tetrahedron graph” $\Gamma_{tetra}$ shown in Figure 19.
Figure 19. Tetrahedron graph $\Gamma_{tetra}$.
Proposition 14.8.
The Weil-Petersson form $\omega_{WP}$ on $\mathcal{M}_{3}$ coincides with the following form in coordinates $\{\ell_{e_{j}},\beta_{e_{j}}\}_{j=1}^{6}$
$\begin{split}&\omega_{WP}={\mathrm{d}}\ell_{e_{2}}\wedge{\mathrm{d}}\ell_{e_{1}}+{\mathrm{d}}\ell_{e_{1}}\wedge{\mathrm{d}}\ell_{e_{3}}+{\mathrm{d}}\ell_{e_{3}}\wedge{\mathrm{d}}\ell_{e_{2}}+{\mathrm{d}}\ell_{e_{4}}\wedge{\mathrm{d}}\ell_{e_{2}}+{\mathrm{d}}\ell_{e_{2}}\wedge{\mathrm{d}}\ell_{e_{5}}+\\ &+{\mathrm{d}}\ell_{e_{5}}\wedge{\mathrm{d}}\ell_{e_{4}}+{\mathrm{d}}\ell_{e_{1}}\wedge{\mathrm{d}}\ell_{e_{4}}+{\mathrm{d}}\ell_{e_{4}}\wedge{\mathrm{d}}\ell_{e_{6}}+{\mathrm{d}}\ell_{e_{6}}\wedge{\mathrm{d}}\ell_{e_{1}}+{\mathrm{d}}\ell_{e_{3}}\wedge{\mathrm{d}}\ell_{e_{6}}+\\ &+{\mathrm{d}}\ell_{e_{6}}\wedge{\mathrm{d}}\ell_{e_{5}}+{\mathrm{d}}\ell_{e_{5}}\wedge{\mathrm{d}}\ell_{e_{3}}+{\mathrm{d}}\beta_{e_{1}}\wedge{\mathrm{d}}\ell_{e_{1}}+{\mathrm{d}}\beta_{e_{2}}\wedge{\mathrm{d}}\ell_{e_{2}}+{\mathrm{d}}\beta_{e_{3}}\wedge{\mathrm{d}}\ell_{e_{3}}+\\ &+{\mathrm{d}}\beta_{e_{4}}\wedge{\mathrm{d}}\ell_{e_{4}}+{\mathrm{d}}\beta_{e_{5}}\wedge{\mathrm{d}}\ell_{e_{5}}+{\mathrm{d}}\beta_{e_{6}}\wedge{\mathrm{d}}\ell_{e_{6}}\;.\end{split}$ (92)
Proof.
The proof is by direct computation using Theorem 9.2, equation (51), applied to $\Gamma_{tetra}$. The form $\Omega(\Gamma_{tetra})$ coincides with $\omega_{WP}$ on $\mathcal{M}_{3}$ by Proposition 11.10. ∎
Proposition 14.9.
For the tetrahedron graph $\Gamma_{tetra}$, the generating function $G(\Gamma_{tetra})$ takes the following expression
$\begin{split}&G(\Gamma_{tetra})=-\frac{3}{2}\left[\mathrm{Li}_{2}\left(\lambda_{e_{1}}\lambda_{e_{2}}\lambda_{e_{3}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{1}}\lambda_{e_{2}}}{\lambda_{e_{3}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{3}}\lambda_{e_{1}}}{\lambda_{e_{2}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{2}}\lambda_{e_{3}}}{\lambda_{e_{1}}}\right)\right]-\\ &-\mathrm{Li}_{2}\left(1-\lambda_{e_{1}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{2}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{3}}^{2}\right)+\frac{\ln^{2}{(\lambda_{e_{1}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{2}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{3}})}}{4}-\\ &-\frac{\ln{(\lambda_{e_{1}}\lambda_{e_{2}}\lambda_{e_{3}})}}{2}-\frac{3}{2}\left[\mathrm{Li}_{2}\left(\lambda_{e_{2}}\lambda_{e_{4}}\lambda_{e_{5}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{2}}\lambda_{e_{4}}}{\lambda_{e_{5}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{2}}\lambda_{e_{5}}}{\lambda_{e_{4}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{4}}\lambda_{e_{5}}}{\lambda_{e_{2}}}\right)\right]-\\ &-\mathrm{Li}_{2}\left(1-\lambda_{e_{2}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{4}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{5}}^{2}\right)+\frac{\ln^{2}{(\lambda_{e_{2}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{4}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{5}})}}{4}-\\ &-\frac{\ln{(\lambda_{e_{2}}\lambda_{e_{4}}\lambda_{e_{5}})}}{2}-\frac{3}{2}\left[\mathrm{Li}_{2}\left(\lambda_{e_{3}}\lambda_{e_{5}}\lambda_{e_{6}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{3}}\lambda_{e_{5}}}{\lambda_{e_{6}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{3}}\lambda_{e_{6}}}{\lambda_{e_{5}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{5}}\lambda_{e_{6}}}{\lambda_{e_{3}}}\right)\right]-\\ &-\mathrm{Li}_{2}\left(1-\lambda_{e_{3}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{5}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{6}}^{2}\right)+\frac{\ln^{2}{(\lambda_{e_{3}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{5}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{6}})}}{4}-\\ &-\frac{\ln{(\lambda_{e_{3}}\lambda_{e_{5}}\lambda_{e_{6}})}}{2}-\frac{3}{2}\left[\mathrm{Li}_{2}\left(\lambda_{e_{1}}\lambda_{e_{4}}\lambda_{e_{6}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{1}}\lambda_{e_{4}}}{\lambda_{e_{6}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{4}}\lambda_{e_{6}}}{\lambda_{e_{1}}}\right)+\mathrm{Li}_{2}\left(-\frac{\lambda_{e_{1}}\lambda_{e_{6}}}{\lambda_{e_{4}}}\right)\right]-\\ &-\mathrm{Li}_{2}\left(1-\lambda_{e_{1}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{4}}^{2}\right)-\mathrm{Li}_{2}\left(1-\lambda_{e_{6}}^{2}\right)+\frac{\ln^{2}{(\lambda_{e_{1}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{4}})}}{4}+\frac{\ln^{2}{(\lambda_{e_{6}})}}{4}-\\ &-\frac{\ln{(\lambda_{e_{1}}\lambda_{e_{4}}\lambda_{e_{6}})}}{2}\;.\end{split}$ (93)
Proof.
Applying Theorem 11.11 to the graph $G_{tetra}$ which has only standard edges we get
$\begin{split}G(\Gamma_{tetra})&=g_{0}^{(1)}(\lambda_{e_{3}},\lambda_{e_{2}};\lambda_{e_{1}})+g_{0}^{(3)}(\lambda_{e_{4}},\lambda_{e_{6}};\lambda_{e_{1}})+g_{0}^{(1)}(\lambda_{e_{1}},\lambda_{e_{3}};\lambda_{e_{2}})+g_{0}^{(2)}(\lambda_{e_{5}},\lambda_{e_{4}};\lambda_{e_{2}})+\\ &+g_{0}^{(1)}(\lambda_{e_{2}},\lambda_{e_{1}};\lambda_{e_{3}})+g_{0}^{(4)}(\lambda_{e_{6}},\lambda_{e_{5}};\lambda_{e_{3}})+g_{0}^{(2)}(\lambda_{e_{2}},\lambda_{e_{5}};\lambda_{e_{4}})+g_{0}^{(3)}(\lambda_{e_{6}},\lambda_{e_{1}};\lambda_{e_{4}})+\\ &+g_{0}^{(2)}(\lambda_{e_{4}},\lambda_{e_{2}};\lambda_{e_{5}})+g_{0}^{(4)}(\lambda_{e_{3}},\lambda_{e_{6}};\lambda_{e_{5}})+g_{0}^{(3)}(\lambda_{e_{1}},\lambda_{e_{4}};\lambda_{e_{6}})+g_{0}^{(4)}(\lambda_{e_{5}},\lambda_{e_{3}};\lambda_{e_{6}})\;.\end{split}$ (94)
By a direct computation very similar to the proof for the expression of $G(\widetilde{\Gamma}_{trin})$ in Proposition 14.4 we get equation (93). ∎